### v3

# Equivariant scaling asymptotics for Poisson and Szegő kernels on Grauert tube boundaries

Abstract
Let $(M,\kappa)$ be a closed and connected real-analytic Riemannian manifold, acted upon by a compact Lie group of isometries $G$. We consider the following two kinds of equivariant asymptotics along a fixed Grauer tube boundary $X^{\tau}$ of $(M,\kappa)$.
1): Given the induced unitary representation of $G$ on the eigenspaces of the Laplacian of $(M,\kappa)$, these split over the irreducible representations of $G$. On the other hand, the eigenfunctions of the Laplacian of $(M,\kappa)$ admit a simultaneous complexification to some Grauert tube. We study the asymptotic concentration along $X^{\tau}$ of the complexified eigenfunctions pertaining to a fixed isotypical component.
2): There are furthermore an induced action of $G$ as a group of CR and contact automorphisms on $X^{\tau}$, and a corresponding unitary representation on the Hardy space $H(X^{\tau})$. The action of $G$ on $X^{\tau}$ commutes with the homogeneous ‘geodesic flow’  and the representation on the Hardy space commutes with the elliptic self-adjoint Toeplitz operator induced by the generator of the goedesic flow. Hence each eigenspace of the latter also splits over the irreducible representations of $G$. We study the asymptotic concentration of the eigenfunctions in a given isotypical component.
We also give some applications of these asymptotics.
1 Introduction
Let $M$ be a compact and connected $d$-dimensional real-analytic (in the following, $\mathcal{C}^{\varpi}$) manifold. There exists an essentially unique complex manifold $(\tilde{M},J)$, the so-called Bruhat-Whitney complexification of $M$, in which $M$ embeds as totally real submanifold (i.e. $T_{x}M\cap J_{x}\left(T_{x}M\right)=\{0\}$ for any $x\in M$) [BW]. More precisely, $(\tilde{M},J)$ is uniquely determined as a germ of complex manifold along $M$; in particular, since both $(\tilde{M},J)$ and $(\tilde{M},-J)$ are complexifications of $M$, there is an anti-holomorphic involution $\sigma:\tilde{M}\rightarrow\tilde{M}$ having $M$ as its fixed locus.
As proved and discussed in [GStz-1991], [GStz-1992], [Lem], [LS], [SZ02] (see also [PW-1989] and [PW-1991]), the choice of a $\mathcal{C}^{\varpi}$ Riemannian metric $\kappa$ on $M$ singles out a unique $\mathcal{C}^{\varpi}$ exhaustion function on $\tilde{M}$ (perhaps after passing to a smaller tubular neighbourhood of $M$ in $\tilde{M}$), with the following properties:
  1. 1.
    $\rho:\tilde{M}\rightarrow[0,+\infty)$, and $M=\rho^{-1}(0)$;
  2. 2.
    $\rho=\rho\circ\sigma$;
  3. 3.
    $\rho$ is strictly plurisubharmonic, i.e. it induces a Kähler form $\Omega$ and a Riemannian metric $\hat{\kappa}$ on $\tilde{M}$ given by
    $\Omega:=\imath\partial\overline{\partial}\rho,\qquad\hat{\kappa}:=\Omega(\cdot,J(\cdot)).$ (1)
  4. 4.
[…]
  5. 5.
    on $\tilde{M}\setminus M$ the function $\sqrt{\rho}$ satisfies the complex homogeneous Monge-Ampère equation:
    $\mathrm{det}\left(\frac{\partial^{2}\sqrt{\rho}}{\partial z_{i}\partial\overline{z}_{j}}\right)=0.$
Without pretense of completeness, we refer the reader to [B], [BH], [G], [Z-2007], [Z-2012], [Z-2020], [CR1] and [CR2] for a wider foundational discussion.
Let $\tau_{0}^{2}:=\sup(\rho)$; then $\tau_{0}\in(0,+\infty]$, and for any $\tau\in(0,\tau_{0})$ the (open) Grauert tube of radius $\tau$ and its boundary are given by
$\tilde{M}^{\tau}:=\rho^{-1}\big{(}[0,\tau^{2})\big{)},\qquad X^{\tau}:=\partial\tilde{M}^{\tau}=\left\{\sqrt{\rho}=\tau\right\}.$ (2)
Grauert tubes have been studied extensively in recent years, both for their intrinsic geometric interest and in relation to the a fundamental holomorphic extension property of the eigenfunctions of the non-negative Laplacian $\Delta$ of $(M,\kappa)$, which was first discovered by Bouted de Monvel [BdM-1978]; this foundational result was also proved and analyzed in [Leb], [S-2014], [S-2015],[Z-2020] (see furthermore the discussion in [GStz-1992]). Grauert tube techniques have also been applied to the study of nodal sets (see [CT-2016], [CT-2018], [TZ-2009], [TZ-2021]).
Let us recall Boutet de Monvel’s foundational result, as formulated in [Leb]. Denote by
$0=\mu_{1}^{2}<\mu_{2}^{2}<\cdots\uparrow+\infty\qquad\text{where}\qquad\mu_{j}\geq 0$ (3)
the distinct eigenvalues of $\Delta$; for every $j=1,2,\cdots$, let $W_{j}\subset\mathcal{C}^{\infty}(M)$ be the (finite-dimensional) eigenspace for the eigenvalue $\mu_{j}^{2}$, and let $(\varphi_{j,k})_{k=1}^{n_{j}}$ a real orthonormal basis of $W_{j}$. For $\mu\in\mathbb{R}$, let us set $\langle\mu\rangle:=\left(1+\mu^{2}\right)^{1/2}$.
Theorem 1.
(Boutet de Monvel) There exists $\tau_{1}\in(0,\tau_{0}]$ such that the following holds:
  1. 1.
    every eigenfunction $\varphi$ of $\Delta$ admits a holomorphic extension $\tilde{\varphi}$ to $\tilde{M}^{\tau_{1}}$, hence for any $\tau\in(0,\tau_{1})$ the restriction $\tilde{\varphi}^{\tau}:=\left.\tilde{\varphi}\right|_{X^{\tau}}$ to $X^{\tau}$ is a CR function;
  2. 2.
    the sequence of functions
    $\left(e^{-\tau\,\mu_{j}}\,\langle\mu_{j}\rangle^{(d-1)/4}\,\tilde{\varphi}_{j,k}^{\tau}\right)_{j\geq 1,\,k=1,\ldots,n_{j}}$
    is a Riesz basis of $H(X^{\tau})$.
If $\tau\in(0,\tau_{0})$, on $X^{\tau}$ we have the following geometric structures (see §3.2 of [P-2024], to which we conform for conventions and notation):
  1. 1.
    a contact form $\alpha^{\tau}$, given a follows: if $\jmath^{\tau}:X^{\tau}\hookrightarrow\tilde{M}$ is the inclusion,
    $\alpha^{\tau}:={\jmath^{\tau}}^{*}(\alpha),\qquad\text{where}\qquad\alpha:=\Im(\partial\rho);$ (4)
  2. 2.
    a CR structure
    $\mathcal{H}^{\tau}:=TX^{\tau}\cap J(TX^{\tau})=\ker(\alpha^{\tau})$ (5)
    (the maximal complex sub-vector bundle of the tangent bundle $TX^{\tau}$);
  3. 3.
    the closed symplectic cone sprayed by $\alpha^{\tau}$ in the complement of the zero section $X^{\tau}_{0}$ of the cotangent bundle $T^{\vee}X^{\tau}$ of $X^{\tau}$:
    $\Sigma^{\tau}:=\left\{\left(x,r\,\alpha^{\tau}_{x}\right)\,:\,x\in X^{\tau},\,r>0\right\}\subseteq T^{\vee}X^{\tau}\setminus X^{\tau}_{0};$ (6)
  4. 4.
    the smooth vector field $\upsilon_{\sqrt{\rho}}^{\tau}\in\mathfrak{X}(X^{\tau})$ given by the restriction to $X^{\tau}$ of the Hamiltonian vector field $\upsilon_{\sqrt{\rho}}$ of $\sqrt{\rho}$ with respect to $\Omega$;
  5. 5.
    the Reeb vector field $\mathcal{R}^{\tau}$ of $(X^{\tau},\alpha^{\tau})$, given by
    $\mathcal{R}^{\tau}:=-\frac{1}{\tau}\,\upsilon_{\sqrt{\rho}}^{\tau};$ (7)
  6. 6.
    the real vector sub-bundle $\mathcal{T}^{\tau}$ of the tangent bundle $TX^{\tau}$ generated $\mathcal{R}^{\tau}$, so that
    $TX^{\tau}=\mathcal{T}^{\tau}\oplus\mathcal{H}^{\tau}$ (8)
    (we shall occasionally refer to $\mathcal{T}^{\tau}$ and $\mathcal{H}^{\tau}$ as, respectively, the $\alpha^{\tau}$-vertical and the $\alpha^{\tau}$-horizontal tangent bundles of $X^{\tau}$);
  7. 7.
    a ‘Riemannian’  volume form $\mathrm{vol}^{R}_{X^{\tau}}$, given by the contraction of the gradient of $\sqrt{\rho}$ with the symplectic volume form of $(\tilde{M}\setminus M,\Omega)$, restricted to $X^{\tau}$.
In addition, on $X^{\tau}$ we have the following analytic structures:
  1. 1.
    the Hardy space $H(X^{\tau}):=\mathrm{ker}(\overline{\partial}_{b})\subseteq L^{2}(X^{\tau})$;
  2. 2.
    the corresponding Szegő projector and its distributional kernel, the so-called Szegő kernel:
    $\Pi^{\tau}:L^{2}(X^{\tau})\longrightarrow H(X^{\tau}),\qquad\Pi^{\tau}(\cdot,\cdot)\in\mathcal{D}^{\prime}(X^{\tau}\times X^{\tau})$ (9)
    (what follows is pivoted on the microlocal description of $\Pi^{\tau}$ as a Fourier integral operator in [BdM-S]);
  3. 3.
    a privileged collection of CR functions (elements of $H(X^{\tau})$), given by the restrictions $\tilde{\varphi}^{\tau}$ to $X^{\tau}$ of the holomorphic extensions $\tilde{\varphi}$ of the eigenfunctions $\varphi$ of $\Delta$ (if $\tau\in(0,\tau_{1})$);
  4. 4.
    the self-adjoint first-order elliptic Toeplitz operator
    $\mathfrak{D}^{\tau}_{\sqrt{\rho}}:=\Pi^{\tau}\circ D^{\tau}_{\sqrt{\rho}}\circ\Pi^{\tau},\qquad\text{where}\qquad D^{\tau}_{\sqrt{\rho}}:=\imath\,\upsilon^{\tau}_{\sqrt{\rho}}.$ (10)
In fact, the principal symbol of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$ (i.e., the restriction to $\Sigma^{\tau}$ in (6) of the principal symbol of $D^{\tau}_{\sqrt{\rho}}$) is strictly positive; hence the spectrum of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$ is discrete and accumulates only at $\infty$ [BdM-G].
In a series of papers, Zelditch has laid the foundations and paved the way for the asymptotic study of both the complexified eigenfunctions of the Laplacian and of the eigenfunctions of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$ (see e.g. [Z-2007], [Z-2012], [Z-2014], [Z-2017], [Z-2020]). These asymptotics are encrypted in certain ‘tempered’  complexified spectral projectors (in the case of the Laplacian) and in a smoothed spectral projector (in the case of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$), whose definitions we briefly recall.
In the case of the Laplacian, the asymptotic distribution of the complexified eigenfunctions of $\Delta$, restricted to $X^{\tau}$, is encapsulated in the asymptotics for $\lambda\rightarrow+\infty$ of the tempered complexified projector kernels $P^{\tau}_{\chi,\,\lambda}\in\mathcal{C}^{\infty}(X^{\tau}\times X^{\tau})$ given by
$P^{\tau}_{\chi,\,\lambda}(x,y):=\sum_{j\geq 1}\hat{\chi}(\lambda-\mu_{j})e^{-2\tau\mu_{j}}\sum_{k=1}^{n_{j}}\tilde{\varphi}^{\tau}_{j,k}(x)\,\overline{\tilde{\varphi}^{\tau}_{j,k}(y)},$ (11)
where $\chi\in\mathcal{C}^{\infty}_{c}(\mathbb{R})$ (the space of compactly supported smooth functions on $\mathbb{R}$) and $\hat{\chi}$ is its Fourier transform. Since the restrictions $(\tilde{\varphi}^{\tau}_{j,k})_{k}$ are not orthonormal, the inner sum may not be interpreted as a genuine projector, even in a rescaled sense.
Similarly, let $\lambda_{1}^{\tau}<\lambda_{2}^{\tau}<\cdots\uparrow+\infty$ denote the distinct eigenvalues of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$, and for each $j$ let $H(X^{\tau})_{j}\subset H(X^{\tau})$ be the eigenspace of $\lambda_{j}^{\tau}$. For $j=1,2,\ldots$, let $(\rho_{j,k}^{\tau})_{k=1}^{\ell_{j}^{\tau}}$ be an orthonormal basis of $H(X^{\tau})_{j}$. The asymptotic distribution of the $\rho_{j,k}^{\tau}$’s and of the $\lambda_{j}^{\tau}$’s is probed by smoothed projection kernels of the form $\Pi^{\tau}_{\chi,\,\lambda}\in\mathcal{C}^{\infty}(X^{\tau}\times X^{\tau})$ given by
$\Pi^{\tau}_{\chi,\,\lambda}(x,y):=\sum_{j\geq 1}\hat{\chi}\left(\lambda-\lambda_{j}^{\tau}\right)\,\sum_{k=1}^{\ell_{j}^{\tau}}\rho_{j,k}(x)\,\overline{\rho_{j,k}(y)}.$ (12)
This is the Schwartz kernel of the smoothed projector
$\Pi^{\tau}_{\chi,\,\lambda}:=\sum_{j\geq 0}\hat{\chi}\left(\lambda-\lambda_{j}^{\tau}\right)\Pi^{\tau}_{j},$ (13)
where $\Pi^{\tau}_{j}:L^{2}(X^{\tau})\rightarrow H(X^{\tau})_{j}$ is the orthogonal projector.
Remark 2.
In the following, for notational simplicity we shall omit the suffix $\tau$ from the spectral data $\lambda_{j}^{\tau}$ and $\ell_{j}^{\tau}$.
While (11) is peculiar to the Grauert tube setting, (12) is the counterpart of classically studied kernels in the spectral theory of pseudodifferential operators (see e.g. [DG] and [GrSj]); in the Toeplitz setting, in the geometric context of positive line bundles and with an emphasis on local scaling asymptotics, they have been used in [P-2009], [P-2010], [P-2011], [P1-2012], [P-2017], [P-2018], [ZZ-2018], [ZZ1-2019], [ZZ2-2019].
In two recent papers ([CR1] and [CR2]), Chang and Rabinowitz have made groundbreaking progress in pushing forward the analogy between the line bundle and the Grauert tube settings, providing local scaling asymptotics for (11) and (12) bearing a striking similarity with the local scaling asymptotics for the Fourier components of Szegő kernels in the line bundle setting ([Z-1998], [BSZ] and [SZ02]). Besides the description of $\Pi^{\tau}$ as a Fourier integral operator in [BdM-S], which underpins the whole theory, the approach of Chang and Rabinowitz in based on the following two pillars. The first is the description, due to Zelditch, of certain 1-parameter groups of Toeplitz operators as ‘dynamical Toeplitz operators’, which goes back to [Z-1997] and is pervasive in his treatment of the subject. The second is a clever use of the ‘Heisenberg local coordinates’  adapted to a hypersurface in a complex manifold introduced by Folland and Stein in [FS1] and [FS2]. The approach of Chang and Rabinowitz has been reviewed and elaborated in [P-2024], where a refinement of their asymptotic expansions is given in the near-diagonal case, where $\chi\in\mathcal{C}^{\infty}_{0}\big{(}(-\epsilon,\epsilon)\big{)}$.
The aim of this paper is to extend the results of [P-2024] to the (off-diagonal and) equivariant setting. Suppose given a smooth action $\mu:G\times M\rightarrow M$ of a $d_{G}$-dimensional compact Lie group $G$ as a group of symmetries on $(M,\kappa)$. Then there are naturally induced unitary representations of $G$ on $L^{2}(M)$ and $H(X^{\tau})$, commuting with $\Delta$ and $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$, respectively. Therefore, for every $j\geq 1$ both the eigenspaces $W_{j}\subset L^{2}(M)$ of $\Delta$ and $H(X^{\tau})_{j}\subset H(X^{\tau})$ are invariant.
Notation 3.
Let $\hat{G}$ denote the unitary dual of $G$; for every $\nu\in\hat{G}$, we shall adopt the following notation.
  1. 1.
    $V_{\nu}$ is the representation space of $\nu$, and $\dim(\nu):=\dim(V_{\nu})$;
  2. 2.
    $\Xi_{\nu}:G\rightarrow\mathbb{C}$ is the character of $\nu$.
  3. 3.
    If $G$ is unitarily represented on a Hilbert space $H$, $H_{\nu}\subseteq H$ will denote the $\nu$-th isotypical component of $H$.
By the Theorem of Peter and Weyl (see e.g. [S-95]) there are unitary equivariant Hilbert direct sum decompositions
$L^{2}(M)=\bigoplus_{\nu\in\hat{G}}L^{2}(M)_{\nu},\quad W_{j}=\bigoplus_{\nu\in\hat{G}}W_{j,\nu},$ (14)
where $W_{j,\nu}:=W_{j}\cap L^{2}(M)_{\nu}$, and similarly
$H(X^{\tau})=\bigoplus_{\nu\in\hat{G}}H(X^{\tau})_{\nu},\quad H(X^{\tau})_{j}=\bigoplus_{\nu\in\hat{G}}H(X^{\tau})_{j,\nu},$ (15)
where $H(X^{\tau})_{j,\nu}:=H(X^{\tau})_{j}\cap H(X^{\tau})_{\nu}$.
For every $(j,\nu)$ such that $W_{j,\nu}\neq(0)$, let $(\varphi_{j,\nu,k})_{k=1}^{n_{j,\nu}}$ be a real orthonormal basis of $W_{j,\nu}$. The equivariant analogue of (11) is the smooth kernel $P^{\tau}_{\chi,\nu,\lambda}(\cdot,\cdot)\in\mathcal{C}^{\infty}(X^{\tau}\times X^{\tau})$ given by
$P^{\tau}_{\chi,\nu,\lambda}(x,y):=\sum_{j\geq 1}\hat{\chi}(\lambda-\mu_{j})e^{-2\tau\mu_{j}}\sum_{k=1}^{n_{j,\nu}}\tilde{\varphi}^{\tau}_{j,\nu,k}(x)\,\overline{\tilde{\varphi}^{\tau}_{j,\nu,k}(y)}.$ (16)
Similarly, for every $(j,\nu)$ for which $H(X^{\tau})_{j,\nu}\neq(0)$, let $(\rho_{j,\nu,k})_{k=1}^{\ell_{j,\nu}}$ be an orthonormal basis of $H(X^{\tau})_{j,\nu}$. The equivariant analogue of (12) is the smooth kernel $\Pi^{\tau}_{\chi,\nu,\lambda}(\cdot,\cdot)\in\mathcal{C}^{\infty}(X^{\tau}\times X^{\tau})$ given by
$\Pi^{\tau}_{\chi,\nu,\lambda}(x,y):=\sum_{j\geq 1}\hat{\chi}\left(\lambda-\lambda_{j}\right)\,\sum_{k=1}^{\ell_{j,\nu}}\rho_{j,\nu,k}(x)\,\overline{\rho_{j,\nu,k}(y)}.$ (17)
Thus $\Pi^{\tau}_{\chi,\nu,\lambda}(\cdot,\cdot)$ is the Schwartz kernel of the smoothed projector
$\Pi^{\tau}_{\chi,\nu,\lambda}:=\sum_{j\geq 1}\hat{\chi}\left(\lambda-\lambda_{j}\right)\,\sum_{k=1}^{\ell_{j,\nu}}\,\Pi^{\tau}_{\nu,j}=P_{\nu}\circ\Pi_{\chi,\lambda},$ (18)
where $\Pi^{\tau}_{\nu,j}:L^{2}(X^{\tau})\rightarrow H(X^{\tau})_{j,\nu}$ and $P_{\nu}:L^{2}(X^{\tau})\rightarrow L^{2}(X^{\tau})_{\nu}$ are the orthogonal projectors.
We shall provide local scaling asymptotics for (16) and (17). Unlike [P-2024], we are not assuming here that $\chi$ is supported near the origin; in the special action-free case, we shall thus recover a refinement of the near-graph scaling asymptotics of Chang and Rabinowitz in [CR2].
The smooth action $\mu:G\times M\rightarrow M$ extends (perhaps after decreasing $\tau_{0}$) to an action $\tilde{\mu}:G\times\tilde{M}^{\tau_{0}}\rightarrow\tilde{M}^{\tau_{0}}$ of $G$ as a group of biholomorphisms. Since $\mu$ is isometric for $(M,\kappa)$, $\tilde{\mu}$ preserves $\rho$. Thus $G$ acts as a group of Hamiltonian automorphisms of the Kähler manifold $(\tilde{M}^{\tau_{0}},J,\Omega)$, with a moment map $\Phi:\tilde{M}^{\tau_{0}}\rightarrow\mathfrak{g}^{\vee}$ vanishing identically on $M$.
Since $\tilde{\mu}$ preserves $\rho$, it also preserves every $X^{\tau}$, with its CR and contact structure, for $\tau\in(0,\tau_{0})$; let $\tilde{\mu}^{\tau}:G\times X^{\tau}\rightarrow X^{\tau}$ be the resulting contact and CR action. Let us set
$Z:=\Phi^{-1}(0)\subseteq\tilde{M}^{\tau_{0}},\qquad Z^{\tau}:=Z\cap X^{\tau}.$ (19)
What follows depends on the following hypothesis.
Assumption 4.
We shall assume that:
  1. 1.
    $Z\setminus M\neq\emptyset$ (hence $Z^{\tau}\neq\emptyset$ for every $\tau\in(0,\tau_{0})$);
  2. 2.
    $\tilde{\mu}^{\tau}$ is locally free on $Z^{\tau}$.
Remark 5.
Let us notice the following regarding the two conditions in Assumption 4.
  1. 1.
    The second requirement is equivalent to $0\in\mathfrak{g}^{\vee}$ being a regular value of $\left.\Phi\right|_{\tilde{M}^{\tau_{0}}\setminus M}$ (see e.g. [$M$], [GStb]); therefore, it implies that $Z\setminus M$ and $Z^{\tau}$ are submanifolds of $\tilde{M}$ and $X^{\tau}$, respectively. Our scaling asymptotics rely on a direct sum decomposition of $TX^{\tau}$ along $Z^{\tau}$ that depends on this smoothness assumption (see §1.1.2).
  2. 2.
    $\tilde{\mu}^{\tau}$ is a fortiori locally free on $Z^{\tau}$ if $\mu$ itself is locally free, e.g., if $M$ is a principal $G$-bundle; for more examples, see §2.3.
  3. 3.
    Under Assumption 4, $d>d_{G}$ (see Corollary 55 below); for instance, the case of a compact Lie group acting on itself by, say, left translations is not covered by the present analysis. This is because $Z^{\tau}=\emptyset$ in this case (we shall consider this specific situation in a separate paper).
Let $\Gamma:\mathbb{R}\times\tilde{M}\rightarrow\tilde{M}$ denote the flow of $\upsilon_{\sqrt{\rho}}$ ($\Gamma$ will be referred to - with some abuse of language - as the ‘homogeneous geodesic flow’,   since it is intertwined with the latter by the imaginary time exponential map - see [GStz-1992], [LS], [GLS]). Its restriction to $X^{\tau}$ is the flow of $\upsilon_{\sqrt{\rho}}^{\tau}$, and will be denoted $\Gamma^{\tau}:\mathbb{R}\times X^{\tau}\rightarrow X^{\tau}$. Since $\tilde{\mu}^{\tau}$ and $\Gamma^{\tau}$ commute, there is a product action of $G\times\mathbb{R}$ on $X^{\tau}$; our first result is that (16) and (17) asymptotically concentrate near certain compact loci $\mathfrak{X}^{\tau}_{\chi}\subset X^{\tau}\times X^{\tau}$ determined by $\chi$, the moment map, and the orbits of the latter action.
Definition 6.
If $x\in X^{\tau}$ and $\chi\in\mathcal{C}^{\infty}_{c}(\mathbb{R})$, we set
$x^{G\times\chi}:=\left\{\tilde{\mu}^{\tau}_{g}\circ\Gamma^{\tau}_{t}(x)\,:\,g\in G,\,t\in\mathrm{supp}(\chi)\right\}.$
We then pose
$\mathfrak{X}^{\tau}_{\chi}:=\left\{(x_{1},x_{2})\in Z^{\tau}\times Z^{\tau}\,:\,x_{1}\in x_{2}^{G\times\chi}\right\}.$
We also set
$x^{\chi}:=\left\{\Gamma^{\tau}_{t}(x)\,:\,t\in\mathrm{supp}(\chi)\right\}.$
Theorem 7.
For any $C,\,\epsilon^{\prime}>0$, we have
$P^{\tau}_{\chi,\nu,\lambda}(x,y)=O\left(\lambda^{-\infty}\right)\quad\text{and}\quad\Pi^{\tau}_{\chi,\nu,\lambda}(x,y)=O\left(\lambda^{-\infty}\right),$
uniformly for
$\max\left\{\mathrm{dist}_{X^{\tau}}\left(x,y^{G\times\chi}\right),\mathrm{dist}_{X^{\tau}}\left(x,Z^{\tau}\right)\right\}\geq C\,\lambda^{\epsilon^{\prime}-1/2}.$
We are thus led to consider the asymptotics of $P^{\tau}_{\chi,\nu,\lambda}(x,y)$ and $\Pi^{\tau}_{\chi,\nu,\lambda}(x,y)$ when $(x,y)$ ranges in a shrinking neighourhood of a fixed pair $(x_{1},x_{2})\in\mathfrak{X}^{\tau}_{\chi}$.
1.1 Scaling asymptotics
The previous estimates motivate the expectation that near $\mathfrak{X}^{\tau}_{\chi}$ the two kernels satisfy scaling asymptotics on a scale of $O(\lambda^{-1/2})$ exhibing an exponential decay along directions transverse to $\mathfrak{X}^{\tau}_{\chi}$. The aim of the following statements is to substantiate this expectation.
Before giving precise statements, it is in order to premise a few general remarks. While the general approach is heuristically inspired by the line bundle setting, the arguments are significantly more involved in the present context, the primary reason being that the geodesic flow is generally not holomorphic. As the reader will appreciate, Zelditch’ method of dynamical Toeplitz operators plays an essential conceptual and technical role in dealing with this difficulty. Another key ingredient is the use of suitable sets of local coordinates (whose construction goes back to Folland and Stein in [FS1] and [FS2], and which were first used in this context by Chang and Rabinowitz), specifically adapted to the local CR geometry of $X^{\tau}$.
More precisely, as in [P-2024] the asymptotics in point will be formulated in suitable systems of local coordinates at points $x\in X^{\tau}$, called normal Heisenberg local coordinates (in the following, NHLC’s), which are a slight specialization of the Heisenberg local coordinates in [CR1] and [CR2] (see §3.3 of [P-2024]).
NHLC’s centered at $x\in X^{\tau}$ will be written, in additive notation, as $x+(\theta,\mathbf{v})$, where $(\theta,\mathbf{v})\in\mathbb{R}\times\mathbb{R}^{2d-2}$ belongs to a neighbourhood of the origin. In terms of (8), we have
$\left.\frac{\partial}{\partial\theta}\right|_{x}=\mathcal{R}^{\tau}(x)\in\mathcal{T}_{x}^{\tau},\qquad\left.\frac{\partial}{\partial\mathbf{v}}\right|_{x}\in\mathcal{H}_{x}^{\tau}\quad\forall\,\mathbf{v}\in\mathbb{R}^{2d-2}.$ (20)
Let us fix $(x_{1},x_{2})\in\mathfrak{X}^{\tau}_{\chi}$ and NHLC’s on $X^{\tau}$ at $x_{1}$ and $x_{2}$. We shall work in rescaled coordinates and set
$x_{j,\lambda}:=x_{j}+\left(\frac{\theta_{j}}{\sqrt{\lambda}},\frac{\mathbf{v}_{j}}{\sqrt{\lambda}}\right)\qquad(j=1,2).$ (21)
The non-holomorphicity, or equivalently the non-unitarity, of the geodesic flow is encapsulated in the appearance of a (generally) non-unitary matrix $B$ in the local description of $\Gamma^{\tau}$ (see (25) and §2.6.2), and in the ensuing oscillatory and Gaussian integrals computing the asymptotics.
In order to obtain tractable computations and extract more intelligible geometric information, we have in some cases restricted the type of rescaled directions. Specifically, in the near-graph action-free setting of [CR2] (Theorem 15 below) we consider arbitrary rescaled displacements; similarly, in the near-diagonal equivariant case (Theorem 28) we allow arbitrary displacements away from $G$-orbits. However, in the general equivariant case we consider more specific choices of directions (see below).
1.1.1 Scaling asymptotics in the action-free case
To fix ideas and ease of exposition, let us first consider separately the action-free case (i.e. where $G$ is trivial); we shall write $x^{\chi}$ for $x^{G\times\chi}$. In this case, $Z^{\tau}=X^{\tau}$ and by Theorem 7 we have
$P^{\tau}_{\chi,\lambda}(x,y)=O\left(\lambda^{-\infty}\right)\quad\text{and}\quad\Pi^{\tau}_{\chi,\lambda}(x,y)=O\left(\lambda^{-\infty}\right),$
uniformly for
$\mathrm{dist}_{X^{\tau}}\left(x,y^{\chi}\right)\geq C\,\lambda^{\epsilon^{\prime}-1/2}.$
We shall present near-graph scaling asymptotics for $P^{\tau}_{\chi,\lambda}(x_{1,\lambda},x_{2,\lambda})$ and $\Pi^{\tau}_{\chi,\lambda}(x_{1,\lambda},x_{2,\lambda})$ refining those in [CR2], where rescaling is according to Heisenberg type; furthermore, we shall provide an explicit description of the leading order term and an estimate on the degree of the polynomials in the rescaled variables intervening in the lower order terms of the asymptotic expansions.
The leading order term describing the exponential decay in the scaling asymptotics of $P^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ and $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ for $\lambda\rightarrow+\infty$ depends on a real-quadratic complex-valued form on the vector subspace $\mathcal{H}_{x_{1}}^{\tau}\times\mathcal{H}_{x_{2}}^{\tau}\subseteq T_{x_{1}}X^{\tau}\times T_{x_{2}}X^{\tau}$. To describe the latter form, we need a brief digression.
Definition 8.
Given a symplectic matrix $A\in\mathrm{Sp}(2d-2)$, we shall set
$A_{c}:=\mathcal{W}\,A\,\mathcal{W}^{-1}\qquad\text{where}\qquad\mathcal{W}:=\frac{1}{\sqrt{2}}\,\begin{pmatrix}I_{d-1}&\imath\,I_{d-1}\\ I_{d-1}&-\imath\,I_{d-1}\end{pmatrix}$.
Then
$A_{c}=\begin{pmatrix}P&Q\\ \overline{Q}&\overline{P}\end{pmatrix}$, (22)
where $P$ is invertible and $\|P\,\mathbf{z}\|\geq\|\mathbf{z}\|$, $\forall\,\mathbf{z}\in\mathbb{C}^{d-1}$ (§4.1 of [F]).
Definition 9.
Let us identify $\mathbb{R}^{2d-2}\cong\mathbb{C}^{d-1}$ in the usual norm-preserving manner, so that if $Z_{j}\in\mathbb{C}^{d-1}$ corresponds to $\mathbf{v}_{j}\in\mathbb{R}^{2d-2}$ then
$h_{0}(Z_{1},Z_{2})=g_{0}(\mathbf{v}_{1},\mathbf{v}_{2})-\imath\,\omega_{0}(\mathbf{v}_{1},\mathbf{v}_{2}),\qquad(\mathbf{v}_{1},\,\mathbf{v}_{2}\in\mathbb{C}^{d-1}),$
where $h_{0}$, $g_{0}$, and $\omega_{0}$ denote the standard Hermitian, Euclidean, and symplectic structures, respectively. Given $A\in\mathrm{Sp}(2d-2)$, let us define $\Psi_{A}:\mathbb{R}^{2d-2}\times\mathbb{R}^{2d-2}\rightarrow\mathbb{C}$ (or, equivalently, $\Psi_{A_{c}}:\mathbb{C}^{d-1}\times\overline{\mathbb{C}}^{d-1}\rightarrow\mathbb{C}$) as follows. With $P$ and $Q$ as in Definition 8,
$\Psi_{A}(\mathbf{v}_{1},\mathbf{v}_{2})=\Psi_{A_{c}}(Z_{1},Z_{2})$
$:=$ $\frac{1}{2}\,\left(Z_{1}^{\dagger}\,\overline{Q}\,P^{-1}\,Z_{1}+2\,\overline{Z}_{2}^{\dagger}\,P^{-1}\,Z_{1}-\overline{Z}_{2}^{\dagger}\,P^{-1}\,Q\,\overline{Z}_{2}-\|Z_{1}\|^{2}-\|Z_{2}\|^{2}\right).$
Remark 10.
$\Psi_{A}$ plays an important role in the theory of the metaplectic representation (see the discussions in §4 of [F], [D], [ZZ-2018]).
Definition 11.
If $(V,h^{\prime})$ is a Hermitian complex $k$-dimensional vector space, so that $h=g_{h}-\imath\,\omega_{h}$, where $g_{h}:=\Re(h)$ and $\omega_{h}:=-\Im(h)$ are, respectively, an Euclidean product and a symplectic bilinear form on $V$. We define $\psi_{2}^{\omega_{h}}:V\times V\rightarrow\mathbb{C}$ by
$\psi_{2}^{\omega_{h}}(v,v^{\prime}):=h(v,v^{\prime})-\frac{1}{2}\,\|v\|_{h}^{2}-\frac{1}{2}\,\|v^{\prime}\|_{h}^{2}=-\imath\,\omega_{h}(v,v^{\prime})-\frac{1}{2}\,\|v-v^{\prime}\|_{h}^{2},$
where $\|v\|_{h}:=\sqrt{h(v,v)}$.
Notation 12.
When $(V,h)=\left(\mathbb{C}^{k},h_{0}\right)$, where $h_{0}$ is the standard Hermitian product, we shall also view $\psi_{2}:=\psi_{2}^{\omega_{0}}$ as being defined on $\mathbb{R}^{2\,k}\times\mathbb{R}^{2k}$, where $\mathbb{R}^{2k}\cong\mathbb{C}^{k}$ in the standard manner. If $\mathbf{v}_{j}\in\mathbb{R}^{2\,k}$, $j=1,2$, corresponds to $Z_{j}\in\mathbb{C}^{k}$ under the previous identification, we shall equivalently write $\psi_{2}(Z_{1},Z_{2})$ as $\psi_{2}(\mathbf{v}_{1},\mathbf{v}_{2})$ to emphasize the symplectic structure.
Remark 13.
Let us identify the unitary group U(d1)U(d-1) as the maximal compact subgroup $\hat{U}(d-1):=\mathrm{Sp}(2d-2)\cap O(2d-2)$ of $\mathrm{Sp}(2d-2)$ in the standard manner. Then $A\in\hat{U}(d-1)\leqslant\mathrm{Sp}(2d-2)$ if and only if in (22) we have $Q=0$ and $P\in U(d-1)\leqslant\mathrm{GL}_{\mathbb{C}}(d-1)$. Hence if $A\in\hat{U}(d-1)$ then
$\Psi_{A}(\mathbf{v}_{1},\mathbf{v}_{2})$ $=$ $\overline{Z}_{2}^{\dagger}\,P^{-1}\,Z_{1}-\frac{1}{2}\,\|Z_{1}\|^{2}-\frac{1}{2}\,\|Z_{2}\|^{2}$
$=$ $h_{0}(Z_{1},P\,Z_{2})-\frac{1}{2}\,\|Z_{1}\|^{2}-\frac{1}{2}\,\|Z_{2}\|^{2}$
$=$ $\psi_{2}(Z_{1},P\,Z_{2})=\psi_{2}(\mathbf{v}_{1},A\,\mathbf{v}_{2})=\psi_{2}(A^{-1}\,\mathbf{v}_{1},\mathbf{v}_{2}).$
Remark 14.
Let us identify $T_{x}X^{\tau}\cong\mathbb{R}\oplus\mathbb{R}^{2d-2}$ through NHLC’s at $x$, hence $\mathcal{H}^{\tau}_{x}\cong\mathbb{R}^{2d-2}\cong\mathbb{C}^{d-1}$. Then the previous invariants may be viewed as defined on $\mathcal{H}^{\tau}_{x}$, and it is natural to expect that they have an intrinsic geometric meaning. This is so, but the symplectic and Euclidean pairings on $\mathcal{H}^{\tau}_{x}$ corresponding to $\omega_{0}=-\Im(h_{0})$ and $g_{0}=\Re(h_{0})$ are not $\Omega_{x}$ and $\hat{\kappa}_{x}$, but rather their halves $\omega_{x}:=\frac{1}{2}\,\Omega_{x}$ and $\tilde{\kappa}_{x}:=\frac{1}{2}\,\hat{\kappa}_{x}$ (see [P-2024]).
Let us return to our geometric setting and consider, in the action-free case, a pair $(x_{1},x_{2})\in\mathfrak{X}^{\tau}_{\chi}$. Hence there exists $t\in\mathrm{supp}(\chi)$ such that $x_{1}=\Gamma^{\tau}_{t}(x_{2})$. It turns out that if $\mathrm{supp}(\chi)$ is sufficiently small (shorter than $2\,\epsilon$ for some $\epsilon>0$, say), then for any $(x_{1},x_{2})\in\mathfrak{X}^{\tau}_{\chi}$ there exists a unique $t_{1}=t_{1}(x_{1},x_{2})\in\mathrm{supp}(\chi)$ such that $x_{1}=\Gamma^{\tau}_{t_{1}}(x_{2})$ (see Lemma 57 below). If furthermore NHLC’s on $X^{\tau}$ are chosen at $x_{1}$ and $x_{2}$, there is a unique $B=B_{x_{1},x_{2}}\in\mathrm{Sp}(2d-2)$ such that
$\Gamma^{\tau}_{-t_{1}}\big{(}x_{1}+(\theta,\mathbf{v})\big{)}=x_{2}+\big{(}\theta+R_{3}(\theta,\mathbf{v}),\,B\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v})\big{)},$ (23)
where $R_{k}$ (respecively, $\mathbf{R}_{k}$) denotes, here and in the following, a generic real-valued (respectively, vector-valued) function on an open neighbourhood of the origin of some Euclidean space, vanishing to $k$-th order at the origin; (23) is a special case of Lemma 67 below.
Theorem 15.
Assume that $\mathrm{supp}(\chi)$ is sufficiently small and $x_{1}\in x_{2}^{\mathrm{supp}(\chi)}$. Let $t_{1}=t_{1}(x_{1},x_{2})$ be as above, and let $B=B_{x_{1},x_{2}}$ be as in (23). Suppose $C>0$ and $\epsilon^{\prime}\in(0,1/6)$. Then, uniformly for $(\theta_{j},\mathbf{v}_{j})\in T_{x_{j}}X^{\tau}$ with $\|(\theta_{j},\mathbf{v}_{j})\|\leq C\,\lambda^{\epsilon^{\prime}}$, there are asymptotic expansions
$\Pi^{\tau}_{\chi,\lambda}(x_{1,\lambda},x_{2,\lambda})$
$\sim$ $\frac{1}{\sqrt{2\,\pi}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1}\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})+\Psi_{B^{-1}}(\mathbf{v}_{1},\mathbf{v}_{2})\right]}\cdot e^{-\imath\,\lambda\,t_{1}}$
$\cdot e^{\imath\,\theta^{\tau}(x_{1},x_{2})}\cdot\left[\chi(t_{1})+\sum_{k\geq 1}\,\lambda^{-k/2}\,F_{k}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2}\right)\right],$
where $e^{\imath\,\theta^{\tau}(x_{1},x_{2})},\,e^{\imath\,\tilde{\theta}^{\tau}(x_{1},x_{2})}\in S^{1}$ and $F_{k}(x_{1},x_{2};\cdot)$, $\widetilde{F}_{k}(x_{1},x_{2};\cdot)$ are polynomials in the rescaled variables of degree $\leq 3\,k$ and parity $k$.
When $x_{1}=x_{2}$ and $\chi\in\mathcal{C}^{\infty}_{c}\big{(}(-\epsilon,\epsilon)\big{)}$ for suitably small $\epsilon>0$ then $M=I_{2d-2}$, $t_{1}=0$ and $\theta^{\tau}(x_{1},x_{2})=0$ Thus we recover the near-diagonal scaling asymptotic in [P-2024]. Furthermore, we recover the near-graph scaling asymptotics of [CR2], with an explicit determination of the leading factor, by fixing $(\theta_{j},\mathbf{v}_{j})$ and rescaling according to Heisenberg type (that is, in the form $(\theta_{j}/\lambda,\mathbf{v}_{j}/\sqrt{\lambda})$). The analogue of Theorem 15 for $P^{\tau}_{\chi,\nu,\lambda}$ is discussed in Section 1.1.4.
1.1.2 A decomposition of $T_{x}X^{\tau}$ in the general equivariant case
In order to state our results in the general equivariant setting, we need to introduce a finer decomposition of $T_{x}X^{\tau}$, valid at $x\in Z^{\tau}$, than the one dictated by $\alpha^{\tau}$ in (8); this decomposition depends on $\mu^{\tau}$ and $\Phi$, and is the analogue of the decomposition appearing in the equivariant asymptotics in the line bundle setting (see [MZ] and [P-2008]).
Notation 16.
For any $\xi\in\mathfrak{g}$ (the Lie algebra of $G$) we shall denote by $\xi_{X^{\tau}}$ the induced vector field on $X^{\tau}$, and for any $x\in X^{\tau}$ we shall denote by $\mathfrak{g}_{X^{\tau}}(x)\subseteq T_{x}X^{\tau}$ the tangent space at $x$ to the $G$-orbit through $x$.
Remark 17.
Suppose $x\in Z^{\tau}$. Then
  1. 1.
    $\mathfrak{g}_{X^{\tau}}(x)\subseteq\mathcal{H}^{\tau}_{x}$;
  2. 2.
    since $\mathcal{H}^{\tau}_{x}=T_{x}X^{\tau}\cap J_{x}(T_{x}X^{\tau})$, we also have $J_{x}\big{(}\mathfrak{g}_{X^{\tau}}(x)\big{)}\subset\mathcal{H}^{\tau}_{x}$;
  3. 3.
    under Assumption 4, $\dim\big{(}\mathfrak{g}_{X^{\tau}}(x)\big{)}=d_{G}$.
Definition 18.
Suppose $x\in Z^{\tau}$. We set
$T_{x}^{v}X^{\tau}:=\mathfrak{g}_{X^{\tau}}(x),\quad T_{x}^{t}X^{\tau}:=J_{x}\big{(}\mathfrak{g}_{X^{\tau}}(x)\big{)},\quad T_{x}^{h}X^{\tau}:=\mathcal{H}^{\tau}_{x}\cap\left(T_{x}^{v}X^{\tau}\oplus T^{t}_{x}X^{\tau}\right)^{\perp},$
where $\perp$ denotes the Hermitian (equivalently, symplectic or Riemannian) orthocomplement of the complex subspace $T_{x}^{v}X^{\tau}\oplus T^{t}_{x}X^{\tau}\subseteq T_{x}\tilde{M}$.
We shall refer to $T_{x}^{v}X^{\tau},\,T_{x}^{t}X^{\tau},\,T_{x}^{h}X^{\tau}$ as the $\tilde{\mu}^{\tau}$-vertical, $\tilde{\mu}^{\tau}$-transverse, and $\tilde{\mu}^{\tau}$-horizontal tangent spaces at $x$, respectively (the terminology being inspired by the line bundle setting).
We shall accordingly decompose $\mathbf{v}\in\mathcal{H}^{\tau}_{x}$ as
$\mathbf{v}=\mathbf{v}^{t}+\mathbf{v}^{v}+\mathbf{v}^{h},\quad\text{where}\quad\mathbf{v}^{t}\in T_{x}^{t}X^{\tau},\,\mathbf{v}^{v}\in T^{v}_{x}X^{\tau},\,\mathbf{v}^{h}\in T_{x}^{h}X^{\tau}.$
Remark 19.
At any $x\in X^{\tau}$, we have direct sum decompositions
$\mathcal{H}^{\tau}_{x}=\left(T_{x}^{v}X^{\tau}\oplus T^{t}_{x}X^{\tau}\right)\oplus T_{x}^{h}X^{\tau},$
and
$T_{x}Z^{\tau}=\mathcal{T}^{\tau}_{x}\oplus\left(T_{x}^{v}X^{\tau}\oplus T_{x}^{h}X^{\tau}\right),\qquad T_{x}X^{\tau}=T_{x}^{t}X^{\tau}\oplus T_{x}Z^{\tau},$
so that $T_{x}^{t}X^{\tau}$ is the normal space to $Z^{\tau}$ at $x$ (in $X^{\tau}$).
Notation 20.
Given a choice of NHLC’s centered for $X^{\tau}$ at $x$, we obtain a direct sum decomposition of $\mathbb{R}^{2\,d-1}=\mathbb{R}\oplus\mathbb{C}^{d-1}\cong T_{x}X^{\tau}$, dictated by $\alpha^{\tau}$, where $\mathbb{C}^{d-1}\cong\mathcal{H}^{\tau}_{x}$. The finer decomposition of $\mathcal{H}^{\tau}_{x}$ in Definition 18, dictated by $\tilde{\mu}^{\tau}$, determines a corresponding decomposition of $\mathbb{C}^{d-1}$, which - emphasizing dimensions - we shall write in the form
$\mathbb{C}^{d-1}=\mathbb{R}^{d_{G}}_{t}\oplus\mathbb{R}^{d_{G}}_{v}\oplus\mathbb{C}^{d-1-d_{G}}_{h},\quad\text{where}\quad\mathbb{R}^{d_{G}}_{t}=J_{0}(\mathbb{R}^{d_{G}}_{v}).$
We shall shift from real to complex notation, identifying
$\mathbb{C}^{d-1}\cong\mathbb{R}^{2d-2},\quad\mathbb{C}^{d-1-d_{G}}_{h}=\mathbb{R}^{2d-2-2d_{G}}_{h}.$
Any $\mathbf{u}\in\mathbb{R}^{2d-2}$ may thus be decomposed as
$\mathbf{u}=\mathbf{u}^{t}+\mathbf{u}^{v}+\mathbf{v}^{h},\quad\text{where}\quad\mathbf{u}^{t}\in\mathbb{R}^{d_{G}}_{t},\,\mathbf{u}^{v}\in\mathbb{R}^{d_{G}}_{v},\,\mathbf{u}^{h}\in\mathbb{R}^{2d-2-2d_{G}}_{h}.$
1.1.3 Scaling asymptotics for $\Pi^{\tau}_{\chi,\nu,\lambda}$
For the sake of brevity, we shall first discuss the asymptotic expansions for $\Pi^{\tau}_{\chi,\nu,\lambda}$, and then explain the necessary changes for $P^{\tau}_{\chi,\nu,\lambda}$.
Before stating our results, we need to make some recalls and introduce some further notation.
If $x\in X^{\tau}$, we shall denote by $G_{x}\leqslant G$ the stabilizer subgroup of $x$, and by $r_{x}=|G_{x}|$ its cardinality ($r_{x}$ is always finite if $x\in Z^{\tau}$).
Assume as before that Assumption 4 holds and that $\mathrm{supp}(\chi)$ is sufficiently small. Then for any $(x_{1},x_{2})\in\mathfrak{X}_{\chi}^{\tau}$ the following holds.
  1. 1.
    There exists a unique $t_{1}=t_{1}(x_{1},x_{2})$ such that $x_{1}=\tilde{\mu}_{g}^{\tau}\circ\Gamma^{\tau}_{t_{1}}(x_{2})$ for some $g\in G$ (Lemma 57 below).
  2. 2.
    There are exactly $r_{x_{1}}$ elements $g_{l}\in G$ such that $x_{1}=\tilde{\mu}_{g_{l}}^{\tau}\circ\Gamma^{\tau}_{t_{1}}(x_{2})$ (see Corollary 58 and Remark 59 below).
  3. 3.
    The condition $(x_{1},x_{2})\in\mathfrak{X}_{\chi}^{\tau}$ is tantamount to $x_{1}^{G}\cap x_{2}^{\chi}\neq\emptyset$, and one has
    $x_{1}^{G}\cap x_{2}^{\chi}=\{x_{12}\},\quad\text{where}\quad x_{12}:=\Gamma^{\tau}_{t_{1}}(x_{2}).$ (24)
Definition 21.
Under Assumption 4, the effective volume at $x\in Z^{\tau}$, denoted $V_{eff}(x)$, is the volume of the $G$-orbit through $x$, $x^{G}\subset X^{\tau}$ for the induced Riemannian density (with respect to $\tilde{\kappa} -$ see also Remark 14).
Definition 22.
Assume as above that $\chi$ has sufficiently small support, and that $(x_{1},x_{2})\in\mathfrak{X}_{\chi}^{\tau}$. Let $t_{1}(x_{1},x_{2})\in\mathrm{supp}(\chi)$ and $x_{12}\in x_{2}^{\chi}$ be as in (24). Given any choice of NHLC’s at $x_{12}$ and $x_{2}$, let $B\in\mathrm{Sp}(2d-2)$ be defined by the analogue of (23):
$\Gamma^{\tau}_{-t_{1}}\big{(}x_{12}+(\theta,\mathbf{v})\big{)}=x_{2}+\big{(}\theta+R_{3}(\theta,\mathbf{v}),\,B\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v})\big{)}$ (25)
(see Lemma 67 below and the notational remark preceding Theorem 15). Referring to the direct sum decomposition in Notation 20, let us set
$A_{\chi}(x_{1},x_{2})$ $:=$ $\int_{\mathbb{R}^{d_{G}}_{t}}\,\mathrm{d}\mathbf{u}^{t}\,\int_{\mathbb{R}^{d_{G}}_{v}}\,\mathrm{d}\mathbf{u}^{v}\,\int_{\mathbb{R}^{2d-2-2d_{G}}_{h}}\,\mathrm{d}\mathbf{u}^{h}$ (26)
$\left[e^{-\left\|\mathbf{u}^{t}\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{u}^{h}\right\|^{2}-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{u}^{t}\right)-\frac{1}{2}\,\|B\mathbf{u}\|^{2}}\right].$
Definition 23.
Let $x_{1},\,x_{2}\in X^{\tau}$ be as in (24), $M$ as in (25), and $A_{\chi}(x_{1},x_{2})$ as in Definition 22; define $P$ by (22), with $A=B$. We set
$\mathcal{F}_{\chi}(x_{1},x_{2})$ $:=$ $\chi(t_{1})\cdot\frac{\left|\det(P)\right|}{r\cdot V_{eff}(x_{1})}\cdot\frac{A_{\chi}(x_{1},x_{2})}{\pi^{d-1}},$
$\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$ $:=$ $\dim(\nu)\cdot\overline{\Xi_{\nu}\left(g_{l}\right)},$
Remark 24.
Let us note the following:
  1. 1.
    Under the previous assumptions, $A_{\chi}(x_{1},x_{2})$ is finite and non-zero, and can in principle be computed in terms of $B$ (see Appendix A of [F]);
  2. 2.
    if $B$ is orthogonal (e.g., if $B$ is the identity matrix), by an interated Gaussian integration one obtains $A_{\chi}(x_{1},x_{2})=\pi^{d-1}$;
  3. 3.
    if $B$ is orthogonal one also has $|\det(P)|=1$ (since in this case $P$ is a unitary complex matrix), hence $\mathcal{F}_{\chi}(x_{1},x_{2})=\chi(t_{1})/\big{(}r\cdot V_{eff}(x_{1})\big{)}$.
In view of Theorem 7, we expect an exponential decay of $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ (and $P^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$) along normal displacements to $Z^{\tau}$, that is, when $\mathbf{v}_{j}=\mathbf{v}_{j}^{t}\in T_{x_{j}}^{t}X^{\tau}$, and we wish to determine the leading exponent governing. Accordingly, in the general equivariant case we restrict first to dispacements along directions in $T_{x_{j}}^{t}X^{\tau}$ (Theorem 25, case 1); in this case, the leading exponent is determined explicitly. In the same setting, an expansion is also obtained, although with a less explicit determination of the exponent, for more general displacements, provided we take $\mathbf{v}_{2}^{h}=0$ (Theorem 25, case 2). As discussed at the beginning of §1.1, this unequal treatment of $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ is aimed at making the computations more tractable, while obtaining managable results that are sufficient for the applications.
Theorem 25.
Suppose that $\chi\in\mathcal{C}^{\infty}_{c}(\mathbb{R})$ has sufficiently small support, and that $(x_{1},x_{2})\in\mathfrak{X}^{\tau}_{\chi}$. Let $g_{1},\ldots,g_{r_{x_{1}}}\in G$ be the distinct elements such that $x_{1}=\mu^{\tau}_{g_{l}}\circ\Gamma^{\tau}_{t_{1}}(x_{2})$, where $t_{1}\in\mathbb{R}$ is as in (24). Fix $C>0$ and $\epsilon^{\prime}\in(0,1/6)$.
  1. 1.
    Uniformly for
    $\theta_{j}\in\mathbb{R},\quad\mathbf{v}_{j}=\mathbf{v}_{j}^{t}\in T^{t}_{x_{j}}X^{\tau},\quad|\theta_{j}|,\,\|\mathbf{v}_{j}^{t}\|\leq C\,\lambda^{\epsilon^{\prime}}$
    we have for $\lambda\rightarrow+\infty$
    $\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1,\lambda},x_{2,\lambda})\sim\sum_{l=1}^{r_{x_{1}}}\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1,\lambda},x_{2,\lambda})_{l}$
    where for each $l$ there is an asymptotic expansion
    $\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1,\lambda},x_{2,\lambda})_{l}$
    $\sim$ $e^{-\imath\,\lambda\,t_{1}}\cdot\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot e^{\frac{1}{\tau}\left[\imath\,\sqrt{\lambda}\,\,(\theta_{1}-\theta_{2})-\left(\|\mathbf{v}_{1}^{t}\|^{2}+\|\mathbf{v}_{2}^{t}\|^{2}\right)\right]}$
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{k\geq 1}\,\lambda^{-k/2}\,F_{k,l,\nu}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right];$
    here $e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\in S^{1}$ and $F_{k,l,\nu}(x_{1},x_{2};\cdot)$ is a polynomial in the rescaled variables of degree $\leq 3\,k$ and parity $k$.
  2. 2.
    Under the same assumptions and with the same notation, uniformly for
    $\theta_{j}\in\mathbb{R},\quad|\theta_{j}|\leq C\,\lambda^{\epsilon^{\prime}}$
    $\mathbf{v}_{1}=\mathbf{v}_{1}^{t}+\mathbf{v}_{1}^{h}\in T^{t}_{x_{1}}X^{\tau}\oplus T^{h}_{x_{1}}X^{\tau},\quad\|\mathbf{v}_{1}\|\leq C\,\lambda^{\epsilon^{\prime}}$
    $\mathbf{v}_{2}=\mathbf{v}_{2}^{t}\in T^{t}_{x_{2}}X^{\tau},\quad\|\mathbf{v}_{2}\|\leq C\,\lambda^{\epsilon^{\prime}},$
    we have for $\lambda\rightarrow+\infty$
    $\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1,\lambda},x_{2,\lambda})\sim\sum_{l=1}^{r_{x_{1}}}\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1,\lambda},x_{2,\lambda})_{l}$
    where for each $l$ there is an asymptotic expansion
    $\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1,\lambda},x_{2,\lambda})_{l}$
    $\sim$ $e^{-\imath\,\lambda\,t_{1}}\cdot\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})-\frac{1}{2}\,\langle\mathbf{V},D_{l}\,\mathbf{V}\rangle\right]}$
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{k\geq 1}\,\lambda^{-k/2}\,\tilde{F}_{k,l,\nu}\left(x_{1},x_{2};\theta_{1},\theta_{2},\mathbf{V}\right)\right];$
    here $\tilde{F}_{k,l}(x_{1},x_{2};\cdot)$ is a polynomial in the rescaled variables of degree $\leq 3\,k$ and parity $k$,
    $\mathbf{V}^{\dagger}:=\begin{pmatrix}(\mathbf{v}_{1}^{t})^{\dagger}&(\mathbf{v}_{1}^{h})^{\dagger}&(\mathbf{v}_{2}^{t})^{\dagger}\end{pmatrix}$,
    and $D_{l}=D_{l}^{\dagger}$, $\Re(D_{l})\gg 0$ for every $l$.
Remark 26.
The unitary factor $e^{\imath\,\theta^{\tau}_{t}(x)}$ is a smooth function on $\mathbb{R}\times X^{\tau}$ and is related to the description (by Zelditch) of the unitary Toeplitz operators $e^{\imath\,t\,\mathfrak{D}^{\tau}_{\sqrt{\rho}}}$ as ‘dynamical Toeplitz operators’ (recall (10)); it is therefore an intrinsic invariant, given a priori, of the CR structure of $X^{\tau}$. Similar considerations apply to the unitary factor $e^{\imath\,\tilde{\theta}^{\tau}_{t}(x)}$ appearing in the asymptotics of $P^{\tau}_{\chi,\nu,\lambda}$ of §1.1.4, with $e^{\imath\,t\,\mathfrak{D}^{\tau}_{\sqrt{\rho}}}$ replaced by $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ in (6) below - see (162) and (164) below.
Let us dwell on the special case where
$x_{1}=x_{2}=x\in Z^{\tau}\quad\text{and}\quad\chi\in\mathcal{C}^{\infty}_{c}\big{(}(-\epsilon,\epsilon)\big{)},$
so that $t_{1}=0$ and $e^{\imath\,\theta^{\tau}_{t}(x)}=1$. Furthermore, we fix one system of NHLC’s at $x$, in terms of which in (21) we write
$x_{1,\lambda}:=x+\left(\frac{\theta_{1}}{\sqrt{\lambda}},\frac{\mathbf{v}_{1}}{\sqrt{\lambda}}\right),\quad x_{2,\lambda}:=x+\left(\frac{\theta_{2}}{\sqrt{\lambda}},\frac{\mathbf{v}_{2}}{\sqrt{\lambda}}\right).$ (27)
In (25) we then have $M=I_{2d-2}$.
Notation 27.
Suppose $G_{x}=\{\kappa_{l}\}_{l=1}^{r_{x_{1}}}$. For every $l$, $\mathrm{d}_{x}\tilde{\mu}^{\tau}_{\kappa_{l}}:T_{x}X^{\tau}\rightarrow T_{x}X^{\tau}$ satisfies
$\mathcal{R}^{\tau}(x)=\mathrm{d}_{x}\tilde{\mu}^{\tau}_{\kappa_{l}}\left(\mathcal{R}^{\tau}(x)\right)\quad\Rightarrow\quad\mathrm{d}_{x}\tilde{\mu}^{\tau}_{\kappa_{l}}\left(\left.\frac{\partial}{\partial\theta}\right|_{x}\right)=\left.\frac{\partial}{\partial\theta}\right|_{x}$
and $\mathrm{d}_{x}\tilde{\mu}^{\tau}_{\kappa_{l}}\left(\mathcal{H}^{\tau}_{x}\right)=\mathcal{H}^{\tau}_{x}$ (recall (7) and (20)). Thus the action of $\kappa_{l}\in G_{x}$ on $T_{x}X^{\tau}=\mathcal{T}^{\tau}_{x}\oplus\mathcal{H}^{\tau}_{x}\cong\mathbb{R}\times\mathbb{R}^{2d-2}$ has the form
$\mathrm{d}_{x}\tilde{\mu}^{\tau}_{\kappa_{l}}(\theta,\mathbf{u})=(\theta,\mathbf{u})=\left(\theta,\mathrm{d}_{x}\tilde{\mu}^{\tau}_{\kappa_{l}}(\mathbf{u})\right).$
In the following, we shall adopt the short-hand
$\mathbf{u}^{(l)}:=\mathrm{d}_{x}\tilde{\mu}^{\tau}_{\kappa_{l}^{-1}}(\mathbf{u})\quad\text{for}\quad\mathbf{u}\in\mathbb{R}^{2d-2}\cong\mathcal{H}^{\tau}_{x}.$
Furthermore, if $\mathbf{u}=\mathbf{u}^{t}+\mathbf{u}^{v}+\mathbf{u}^{h}$ then $(\mathbf{u}^{t})^{(l)}=(\mathbf{u}^{(l)})^{t}$, $(\mathbf{u}^{v})^{(l)}=(\mathbf{u}^{(l)})^{v}$, $(\mathbf{u}^{h})^{(l)}=(\mathbf{u}^{(l)})^{h}$.
Theorem 28.
Under the assumptions and with the notation of Theorem 25, if $\chi\in\mathcal{C}^{\infty}_{c}\big{(}(-\epsilon,\epsilon)\big{)}$ and $x_{1}=x_{2}=x\in Z^{\tau}$, then with the previous notation the following holds: uniformly for
$\theta_{j}\in\mathbb{R},\quad\mathbf{v}_{j}=\mathbf{v}_{j}^{h}+\mathbf{v}_{j}^{t}\in T^{h}_{x_{j}}X^{\tau}\oplus T^{t}_{x_{j}}X^{\tau},\quad|\theta_{j}|,\,\|\mathbf{v}_{j}^{t}\|\leq C\,\lambda^{\epsilon^{\prime}},$
we have for $\lambda\rightarrow+\infty$
$\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1,\lambda},x_{2,\lambda})\sim\sum_{l=1}^{r_{x_{1}}}\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1,\lambda},x_{2,\lambda})_{l}$
where for each $l$ there is an asymptotic expansion
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}$ $\sim$ $\frac{1}{\sqrt{2\,\pi}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\,\frac{\dim(\nu)}{r_{x}\cdot V_{eff}(x_{1})}\cdot$
$\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})-\left\|\mathbf{v}_{1}^{t}\right\|^{2}-\left\|\mathbf{v}_{2}^{t}\right\|^{2}+\psi_{2}\left((\mathbf{v}^{h}_{1})^{(l)},\mathbf{v}^{h}_{2}\right)\right]}$
$\cdot\left[\chi(0)\cdot\overline{\Xi_{\nu}\left(\kappa_{l}\right)}+\sum_{k\geq 1}\,\lambda^{-k/2}\,F_{k,l,\nu}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right];$
here $F_{k,l,\nu}(x_{1},x_{2};\cdot)$ is a polynomial in the rescaled variables of degree $\leq 3\,k$ and parity $k$.
Theorem 28 is the analogue in the Grauert tube setting of the equivariant Szegő kernel asymptotics in the line bundle setting of [P-2008].
1.1.4 Scaling asymptotics for $P^{\tau}_{\chi,\nu,\lambda}$
The scaling asymptotics of $P^{\tau}_{\chi,\nu,\lambda}$ can be studied, with some adaptations, by arguments and techniques similar to those used for $\Pi^{\tau}_{\chi,\nu,\lambda}$.
Theorem 29.
The statements of Theorems 7, 15, 25 and 28 apply with $P^{\tau}_{\chi,\nu,\lambda}$ in place of $\Pi^{\tau}_{\chi,\nu,\lambda}$, with the following changes:
  1. 1.
    the leading order terms of the asymptotic expansions in Theorems 25, 15 and 28 are multiplied by $(\lambda/\pi\,\tau)^{-(d-1)/2}$;
  2. 2.
    the real smooth function $\theta^{\tau}_{t}$ in the definition of $F_{\chi,\lambda}$ is replaced by a possibly different function real smooth function $\tilde{\theta}^{\tau}_{t}$.
In the action-free case, for example, rather than the expansion in Theorem 15 we have
$P^{\tau}_{\chi,\lambda}(x_{1,\lambda},x_{2,\lambda})$
$\sim$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\cdot\left(\frac{1}{2}\right)^{d-1}\cdot\left(\frac{\lambda}{\tau\,\pi}\right)^{\frac{d-1}{2}}\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})+\Psi_{M^{-1}}(\mathbf{v}_{1},\mathbf{v}_{2})\right]}\cdot e^{\imath\,\tilde{\theta}^{\tau}_{t_{1}}(x)}\,$
$\cdot\left[\chi(t_{1})+\sum_{k\geq 1}\,\lambda^{-k/2}\,\widetilde{F}_{k}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2}\right)\right],$
for certain polynomials $\widetilde{F}_{k}(x_{1},x_{2};\cdot)$ in the rescaled variables of degree $\leq 3\,k$ and parity $k$. In the yet more particular near-diagonal case, the previous expansion corrects the leading order factor appearing in the expansion in Theorem 7 of [P-2024] by a power of $\pi^{(d-1)/2}$; this is related to the symbolic computation in §5 below (see Lemma 81).
The original expansion in Theorem 1.9 of [CR2] follows by rescaling according to Heisenberg type (as in the Introduction of [P-2024]).
1.2 Unrescaled Asymptotics
The equivariant scaling asymptotic expansions of Theorems 25, 28, and 29 hold uniformly, say, on compact subsets of the locus in $Z^{\tau}$ with principal orbit type, but it is not a priori obvious that they hold uniformly near the locus where the cardinality of the stabilizer has a discontinuity. In the following Theorem, which rests on the previous ones, we establish a near-graph unrescaled version of the previous asymptotic expansions; these allow for some uniform estimates which, in turn, will be useful in the following applications.
We give an explicit statement and proof in the case of $\Pi^{\tau}_{\chi,\nu,\lambda}$; the extension to $P^{\tau}_{\chi,\nu,\lambda}$ can be carried out by the same arguments used for Theorem 29, and will be left to the reader.
Theorem 30.
With the previous assumptions and notation, the following holds for $\lambda\rightarrow+\infty$.
  1. 1.
    For small enough
    $(\theta_{j},\mathbf{v}_{j}^{t})\in\mathbb{R}\times\mathbb{R}^{d_{G}}_{t}\cong\mathcal{T}^{\tau}_{x_{j}}\oplus T^{t}_{x_{j}}X^{\tau},$
    there is an asymptotic expansion
    $\Pi^{\tau}_{\chi,\nu,\lambda}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1}^{t}),x_{2}+(\theta_{2},\mathbf{v}_{2}^{t})\big{)}\sim\sum_{g_{l}\in G_{x_{1}}}\Pi^{\tau}_{\chi,\nu,\lambda}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1}^{t}),x_{2}+(\theta_{2},\mathbf{v}_{2}^{t})\big{)}_{l},$
    where for each $l$
    $\Pi^{\tau}_{\chi,\nu,\lambda}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1}^{t}),x_{2}+(\theta_{2},\mathbf{v}_{2}^{t})\big{)}_{l}$
    $\sim$ $e^{-\imath\,\lambda\,t_{1}}\,\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}$
    $\cdot e^{\frac{\lambda}{\tau}\,\left[\imath\,\,(\theta_{1}-\theta_{2})-\left(\|\mathbf{v}_{1}^{t}\|^{2}+\|\mathbf{v}_{2}^{t}\|^{2}\right)+R_{3}\left(\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right]}$
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{s\geq 1}\,\lambda^{-s}\,L_{\nu,l,s}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right],$
    where $R_{3}$ vanishes to third order at the origin and
    $\Re\left[(\imath\,\,(\theta_{1}-\theta_{2})-\left(\|\mathbf{v}_{1}^{t}\|^{2}+\|\mathbf{v}_{2}^{t}\|^{2}\right)+R_{3}\left(\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right]\leq 0.$
  2. 2.
    For small enough $(\theta_{j},\mathbf{v}_{j})\in\mathbb{R}\times\mathbb{R}^{2d-2}\cong T_{x_{j}}X^{\tau}=\mathcal{T}^{\tau}_{x_{j}}\oplus\mathcal{H}^{\tau}_{x_{j}}$ with
    $\mathbf{v}_{1}=\mathbf{v}_{1}^{t}+\mathbf{v}_{2}^{h}\in T^{t}_{x_{1}}X^{\tau}\oplus T^{h}_{x_{1}}X^{\tau},\quad\mathbf{v}_{2}=\mathbf{v}_{2}^{t}\in T^{t}_{x_{2}}X^{\tau},$
    there is an asymptotic expansion
    $\Pi^{\tau}_{\chi,\nu,\lambda}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1}),x_{2}+(\theta_{2},\mathbf{v}_{2})\big{)}\sim\sum_{g_{l}\in G_{x_{1}}}\Pi^{\tau}_{\chi,\nu,\lambda}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1}),x_{2}+(\theta_{2},\mathbf{v}_{2})\big{)}_{l},$
    where for each $l$
    $\Pi^{\tau}_{\chi,\nu,\lambda}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1}),x_{2}+(\theta_{2},\mathbf{v}_{2})\big{)}_{l}$
    $\sim$ $e^{-\imath\,\lambda\,t_{1}}\,\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}$
    $\cdot e^{\frac{\lambda}{\tau}\,\left[\imath\,\,(\theta_{1}-\theta_{2})-\frac{1}{2}\,\langle\mathbf{V},D_{l}\,\mathbf{V}\rangle+R_{3}\left(\theta_{1},\theta_{2},\mathbf{V}\right)\right]}$
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{s\geq 1}\,\lambda^{-s}\,K_{\nu,l,s}\left(x_{1},x_{2};\theta_{1},\theta_{2},\mathbf{V}\right)\right],$
where $\mathbf{V}$ is as in Theorem 25, $R_{3}$ vanishes to third order at the origin, and
$\Re\left[\imath\,\,(\theta_{1}-\theta_{2})-\frac{1}{2}\,\langle\mathbf{V},D_{l}\,\mathbf{V}\rangle+R_{3}\left(\theta_{1},\theta_{2},\mathbf{V}\right)\right]\leq 0.$
1.3 Applications
We conclude this Introduction with a sample of some of the possible applications of the previous asymptotics.
1.3.1 An equivariant global Weyl law for $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$
The previous expansions may be integrated to obtain global information on the asymptotic distribution of the $\lambda_{j}$’s. Recalling (15), let us set
$\mathcal{W}_{\nu}^{\tau}(\lambda):=\sum_{\lambda_{j}\leq\lambda}\,\dim H^{\tau}(X)_{j,\nu}\qquad(\lambda\in\mathbb{R}).$ (28)
We shall make here the simplifying assumption that $G$ acts freely on $Z^{\tau}$; then the quotient $Z^{\tau}/G$ is a compact manifold; it has a uniquely determined Riemannian structure such that the projection $Z^{\tau}\rightarrow Z^{\tau}/G$ is a Riemannian submersion. Let $\mathrm{vol}(Z^{\tau}/G)$ be the corresponding volume of $Z^{\tau}/G$.
Theorem 31.
In the situation of Theorem 28, let us assume that $G$ acts freely on $Z^{\tau}$. Then as $\lambda\rightarrow+\infty$
$\mathcal{W}_{\nu}^{\tau}(\lambda)$ $=$ $\frac{1}{2^{d_{G}/2}}\cdot\frac{\tau}{d-d_{G}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
$\cdot\left[1+O\left(\lambda^{-1}\right)\right]$
1.3.2 Pointwise estimates on eigenfunctions of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$
Theorem 30 has the following following straighforward consequence.
Proposition 32.
There exists $C_{\nu},\,\lambda_{\nu}^{\tau}>0$ such that
$\Pi^{\tau}_{\chi,\nu,\lambda}(x,x)\leq C_{\nu}\,\left(\frac{\lambda}{\tau}\right)^{d-1-d_{G}/2}\qquad\forall\,x\in X^{\tau},\,\lambda\geq\lambda_{\nu}^{\tau}.$
Corollary 33.
There exist $C_{\nu}>0$ and $j_{\nu}^{\tau}>0$ such that for any $j\geq j_{\nu}^{\tau}$ and $\lambda\geq\lambda_{\nu}^{\tau}$
$\sum_{k}|\rho_{j,\nu,k}(x)|^{2}\leq C_{\nu}\,\left(\frac{\lambda_{j}}{\tau}\right)^{d-1-d_{G}/2}\qquad(x\in X^{\tau}).$
1.3.3 Operator norm estimates for $\Pi^{\tau}_{\chi,\nu,\lambda}$
In the action-free case, Chang and Rabinowitz have established in [CR2] operator norm estimates on $\Pi^{\tau}_{\chi,\lambda}:L^{p}(X^{\tau})\rightarrow L^{q}(X^{\tau})$. Namely, they proved that for $2\leq p,q\leq+\infty$ there exist constants $C^{\tau}_{p}>0$ such that
$\left\|\Pi^{\tau}_{\chi,\lambda}\right\|_{L^{p}\rightarrow L^{q}}\leq C^{\tau}_{p}\,\lambda^{(d-1)\,\left(\frac{1}{p}-\frac{1}{q}\right)}.$ (29)
This result is the analogue of similar estimates proved in the line bundle setting by Shiffman and Zelditch in [SZ03], and the proof follows the same general lines, adapted to the Grauert tube context (for antecedents in the real domain, see [So-1988] and [So-2017]). More precisely, the argument given by Chang and Rabinowitz for (29) is based on the off-graph scaling asymptotics of $\Pi^{\tau}_{\chi,\lambda}$, paired with the Shur-Young inequality [So-2017].
The goal of this section is to describe an equivariant extension of (29), involving the operator norm of $\Pi^{\tau}_{\chi,\nu,\lambda}:L^{p}(X^{\tau})\rightarrow L^{q}(X^{\tau})$.
Theorem 30 allows an adaptation to the equivariant setting of the arguments in [CR2] for the proof of (29).
Theorem 34.
Under the previous assumptions, there exists a constant $\rho^{\tau}_{p,\nu}>0$ such that
$\left\|\Pi^{\tau}_{\chi,\nu,\lambda}\right\|_{L^{p}\rightarrow L^{q}}\leq\rho_{p,\nu}^{\tau}\left(\frac{\lambda}{\tau}\right)^{\left(\frac{1}{p}-\frac{1}{q}\right)\,\left(d-1-\frac{d_{G}}{2}\right)}.$
1.3.4 Estimates on the complexifications $\widetilde{\varphi}^{\tau}_{j,\nu,k}$’s.
We discuss some estimates on equivariant complexified eigenfunctions of $\Delta$, which are equivariant analogues of estimates in [Z-2020]. The proofs are either straightforward, or minor modifications of the ones for the eigenfunctions of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$.
Let $\chi$ be as in §1.3.2; Then $P^{\tau}_{\chi,\nu,\lambda}(x,x)\geq 0$. Furthermore, as we have remarked an analogue of Theorem 30 holds for $P^{\tau}_{\chi,\nu,\lambda}$, leading to the following analogue of Proposition 32:
Proposition 35.
For some constant $C_{\nu}>0$ and every $x\in X^{\tau}$ and $\lambda\gg 0$ we have
$P^{\tau}_{\chi,\nu,\lambda}\left(x,x\right)\leq C_{\nu}\,\left(\frac{\lambda}{\tau}\right)^{(d-1-d_{G})/2}.$
With $\lambda=\mu_{j}$ we deduce the following.
Corollary 36.
If $x\in X^{\tau}$ and $j\gg 0$,
$\sum_{k}\left|\widetilde{\varphi}^{\tau}_{j,\nu,k}(x)\right|^{2}\leq C_{\nu}\,e^{2\,\tau\,\mu_{j}}\,\left(\frac{\mu_{j}}{\tau}\right)^{(d-1-d_{G})/2}.$ (30)
As in [Z-2020], [CR1], [CR2], let us consider the (equivariant) Husimi distributions
$U_{j,\nu,k}(x):=\dfrac{\left|\widetilde{\varphi}^{\tau}_{j,\nu,k}(x)\right|^{2}}{\left\|\widetilde{\varphi}^{\tau}_{j,\nu,k}\right\|_{L^{2}(X^{\tau})}^{2}}.$ (31)
By Lemma 0.2 of [Z-2020],
$\left\|\widetilde{\varphi}^{\tau}_{j,\nu,k}\right\|_{L^{2}(X^{\tau})}^{2}\sim D_{\tau}\,e^{2\,\tau\,\mu_{j}}\,\mu_{j}^{-(d-1)/2}$
for some constant $D_{\tau}>0$. We obtain the following equivariant refinement of the action-free upper bound in [Z-2020].
Corollary 37.
Under the previous assumptions, for $j\gg 0$ and every $k$
$\sup_{x\in X^{\tau}}U_{j,\nu,k}(x)\leq D_{\tau}^{\prime\prime}\,\mu_{j}^{d-1-\frac{1}{2}\,d_{G}}.$
We conclude with an $L^{2}$-analogue of the Weyl law for $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$ in Theorem 31.
Proposition 38.
Under the assumptions and with the notation of Theorem 31, the following asymptotic expansion holds for $\lambda\rightarrow+\infty$:
$\mathrm{trace}\left(P^{\tau}_{\chi,\nu,\lambda}\right)$ $=$ $\sum_{j}\hat{\chi}(\lambda-\mu_{j})\,e^{-2\,\tau\,\mu_{j}}\,\sum_{k}\left\|\widetilde{\varphi}^{\tau}_{j,\nu,k}\right\|_{L^{2}(X^{\tau})}^{2}$
$\sim$ $\frac{1}{\sqrt{2^{d+d_{G}}\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{\frac{d-1}{2}-d_{G}}\,\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
$\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k}\,F_{k,\nu}\left(x\right)\right]$
Setting $\lambda=\mu_{j}$, we obtain the following consequence.
Corollary 39.
There is a constant $C_{\nu}^{\tau}>0$ such that for all $j\gg 0$
$\sum_{k}\left\|\widetilde{\varphi}^{\tau}_{j,\nu,k}\right\|_{L^{2}(X^{\tau})}^{2}\leq C_{\nu}^{\tau}\,e^{2\,\tau\,\mu_{j}}\,\mu_{j}^{\frac{d-1}{2}-d_{G}}.$
The asymptotic expansion in Proposition 38 may be translsated into information about the asymptotics of
$\mathcal{P}^{\tau}_{\nu}(\lambda)=\sum_{\mu_{j}\leq\lambda}e^{-2\,\tau\,\mu_{j}}\,\sum_{k}\left\|\widetilde{\varphi}^{\tau}_{j,\nu,k}\right\|_{L^{2}(X^{\tau})}^{2},$ (32)
essentially by the same argument used for Theorem 31.
Theorem 40.
In the situation of Theorem 31, let us also assume that $d\geq 2\,d_{G}$. Then
$\mathcal{P}_{\nu}^{\tau}(\lambda)$ $\sim$ $\frac{1}{\sqrt{2^{d+1+d_{G}}}\,\pi}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{\frac{d-1}{2}-d_{G}}\,\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
$\cdot\left[\dfrac{\lambda}{\frac{d+1}{2}-d_{G}}+O(1)\right].$
2 Preliminaries
We shall identify the tangent and cotangent bundles of $M$, TMTM and $T^{\vee}M$, by means of $\kappa$.
2.1 The action on the Grauert tubes
In Section 1, the Grauert tubes $\tilde{M}^{\tau}$, $\tau\in(0,\tau_{0})$, have been described as strictly pseudoconvex domains in $\tilde{M}$ intrisically associated to $\kappa$. An alternative characterization of $\tilde{M}^{\tau}$ is in terms of disk bundles $T^{\tau}M\subset T^{\vee}M$ of norm $\tau$ with respect to $\kappa$, where $T^{\vee}M$ has been endowed in a neighbourhood of the zero section $M_{0}$ with a complex structure intrinsically associated to $\kappa$ ([LS], [Lem], [Sz], [GStz-1991], [GStz-1992]). This complex structure was called adapted in [LS], and will be denoted $J_{ad}$; it is uniquely determined by the condition that the Riemannian and Monge-Ampère foliations coincide [LS].
Notation 41.
Let $\gamma:\mathbb{R}\rightarrow M$ be a geodesic, with velocity curve $\dot{\gamma}:\mathbb{R}\rightarrow TM$. Let us define $\psi_{\gamma}:\mathbb{C}\rightarrow TM$ by
$\psi_{\gamma}:a+\imath\,b\mapsto N_{b}\big{(}\dot{\gamma}(a)\big{)},$
where $N_{b}$ denotes dilation by the factor $b$.
Theorem 42.
(Lempert and Szöke) Let $J_{0}$ be the complex structure on $\mathbb{C}$. Given any sufficiently small neighbourhood $T^{\prime}M\subseteq TM$ of the zero section, $J_{ad}$ is the unique complex structure on $T^{\prime}M\subseteq TM$ such that $\psi_{\gamma}:\psi_{\gamma}^{-1}(T^{\prime}M)\rightarrow T^{\prime}M$ is $(J_{0},J_{ad})$-holomorphic for every geodesic $\gamma$ on $(M,\kappa)$.
As proved in [LS] and [GLS], for some tubular neighbourhood $T^{\prime}M\subseteq T^{\vee}M$ of the zero section the imaginary time exponential map $E:T^{\prime}M\rightarrow\tilde{M}$ (essentially defined by complexifying the ordinary exponential map) is $(J_{ad},J)$-holomorphic and intertwines the square norm function with $\rho$ and the canonical symplectic structure $\Omega_{can}$ on $T^{\vee}M$ with $\Omega$ (we conform to the notation and conventions in [P-2024], thus $\Omega_{can}=\mathrm{d}\mathbf{q}\wedge\mathrm{d}\mathbf{p}$ in local $(\mathbf{q},\mathbf{p})$ coordinates). Hence $E$ restricts to an isomorphism of Kähler manifolds
$E^{\tau}:(T^{\tau}M,\Omega_{can},J)\cong(\tilde{M}^{\tau},\Omega,J).$
Notation 43.
Given $f:M\rightarrow M$ smooth, $\mathrm{d}f:TM\rightarrow TM$ will denote its differential (tangent map).
Proposition 44.
If $f:M\rightarrow M$ is a Riemannian isometry of $(M,\kappa)$, then $\mathrm{d}f$ restricts to a biholomorphism of $(T^{\tau}M,J_{\mathrm{ad}})$ into itself, for any $\tau\in(0,\tau_{0})$.
Proof.
The claim is that
$\mathrm{d}(\mathrm{d}f)\circ J_{\mathrm{ad}}=J_{\mathrm{ad}}\circ\mathrm{d}(\mathrm{d}f)$ (33)
on $T(T^{\tau}M)$. Let $(m,v)\in TM\setminus M_{0}$, and let $\gamma:\mathbb{R}\longrightarrow M$ be the unique geodesic with $\dot{\gamma}(0)=(m,v)$. Since $f$ is an isometry, the geodesic corresponding to $\mathrm{d}f(m,v)$ is $f\circ\gamma$, and its velocity is $\mathrm{d}f(\dot{\gamma})$. Hence,
$\mathrm{d}f\circ\psi_{\gamma}(a+\imath\,b)=\mathrm{d}f\circ N_{b}\left(\dot{\gamma}(a)\right)=N_{b}\circ\mathrm{d}f\left(\dot{\gamma}(a)\right)=N_{b}\circ\dot{\overbrace{{f\circ\gamma}}}(a)=\psi_{f\circ\gamma}(a+\imath\,b).$
Thus, by Theorem 42 $\mathrm{d}f\circ\psi_{\gamma}=\psi_{f\circ\gamma}$ is $(J_{0},J_{\mathrm{ad}})$-holomorphic for every $\gamma$ (on the preimage in $\mathbb{C}$ of $T^{\tau}M$). In other worlds,
$J_{\mathrm{ad}}\circ\mathrm{d}(\mathrm{d}f\circ\psi_{\gamma})=\mathrm{d}(\mathrm{d}f\circ\psi_{\gamma})\circ J_{0}.$
By the chain rule,
$\left(\mathrm{d}(\mathrm{d}f)^{-1}\circ J_{\mathrm{ad}}\circ\mathrm{d}(\mathrm{d}f)\right)\circ\mathrm{d}\psi_{\gamma}=\mathrm{d}\psi_{\gamma}\circ J_{0}.$
By the uniqueness part in Theorem 42,
$\mathrm{d}(\mathrm{d}f)^{-1}\circ J_{\mathrm{ad}}\circ\mathrm{d}(\mathrm{d}f)=J_{0},$
i.e. (33) holds. ∎
Corollary 45.
Any isometry of $(M,\kappa)$ is real-analytic.
Therefore, if $f$ is an isometry of $(M,\kappa)$, it uniquely extends to a holomorphic map $\tilde{f}:\tilde{M}^{\tau}\rightarrow\tilde{M}$ for any sufficiently small $\tau>0$.
Corollary 46.
For any $\tau\in(0,\tau_{0})$ the following holds:
  1. 1.
    $\tilde{f}$ is defined on $\tilde{M}^{\tau}$, where it is the biholomorphism
    $\tilde{f}=E^{\tau}\circ\mathrm{d}f\circ(E^{\tau})^{-1}:\tilde{M}^{\tau}\rightarrow\tilde{M}^{\tau};$
  2. 2.
    $\tilde{f}\circ\rho=\rho$, $\tilde{f}^{*}(\Omega)=\Omega$;
  3. 3.
    $\tilde{f}$ restricts to a CR and contact automorphism $\tilde{f}^{\tau}:X^{\tau}\rightarrow X^{\tau}$.
Applying the above to a smooth action by isometries, we conclude the following.
Proposition 47.
Let $(M,\kappa)$ be a compact $\mathcal{C}^{\varpi}$ Riemannian manifold, $G$ a compact Lie group, and $\mu:G\times M\rightarrow M$ a smooth action as a group of isometries of $(M,\kappa)$. Then:
  1. 1.
    for every $g\in G$, $\mu_{g}:M\rightarrow M$ is real-analytic;
  2. 2.
    for $\tau\in(0,\tau_{0})$, $\mu$ extends to a $G$-action $\tilde{\mu}^{\tau}:G\times\tilde{M}^{\tau}\rightarrow\tilde{M}^{\tau}$ as a group of automorphisms of the Kähler manifold $(\tilde{M}^{\tau},\Omega,J)$;
  3. 3.
    $E^{\tau}$ intertwines $\tilde{\mu}^{\tau}$ on $\tilde{M}^{\tau}$ with the (co)tangent action on $T^{\tau}M$;
  4. 4.
    $(\tilde{\mu}^{\tau}_{g})^{*}(\alpha)=\alpha,\,\forall\,g\in G$;
  5. 5.
    $\tilde{\mu}$ restricts to an action $\mu^{\tau}:G\times X^{\tau}\rightarrow X^{\tau}$ as a group of automorphisms of the contact CR manifold $X^{\tau}$.
Proof.
The only thing left to prove is that $\tilde{\mu}_{g}^{*}(\alpha^{\tau})=\alpha^{\tau}$ for all $g\in G$. Since $\tilde{\mu}_{g}$ is holomorphic, preserves $\rho$ and $\alpha=\Im(\partial\rho)$, the thesis follows by restriction to $X^{\tau}$. ∎
The (co)tangent lift of $\mu$ preserves the tautological 1-form $\lambda_{\mathrm{\mathrm{can}}}$ on $T^{\vee}M$ (locally given by $\mathbf{p}\,\mathrm{d}\mathbf{q}$), and is therefore Hamiltonian with respect to $\Omega_{\mathrm{\mathrm{can}}}$, with moment map
$\Phi^{\xi}_{TM}=\langle\Phi_{TM},\xi\rangle=\lambda_{\mathrm{\mathrm{can}}}(\xi_{TM}).$
On the other hand, for every $\tau\in(0,\tau_{0})$
$\alpha=-{E^{\tau}}^{*}(\lambda_{can})\quad\text{on}\quad M^{\tau}$
(see the discussion in the introduction of [P-2024]).
Given $\xi\in\mathfrak{g}$ and $\tau\in(0,\tau_{0})$, we shall denote by $\xi_{\tilde{M}^{\tau_{0}}}$ and $\xi_{X^{\tau}}$ the induced vector fields on $\tilde{M}^{\tau_{0}}$ and $X^{\tau}$, respectively. Hence $\xi_{\tilde{M}^{\tau_{0}}}$ is tangent to $X^{\tau}$ and $\xi_{X^{\tau}}$ is the restriction of $\xi_{\tilde{M}^{\tau_{0}}}$ to $X^{\tau}$.
By the previous discussion (see also Lemma 22 of [P-2024]), we conclude the following.
Corollary 48.
With the preceding assumptions and notation, for every $\tau\in(0,\tau_{0})$ the following holds.
  1. 1.
    $\tilde{\mu}^{\tau}$ is Hamiltonian for $\Omega$, with moment map $\Phi_{\tilde{M}^{\tau}}:\tilde{M}^{\tau}\rightarrow\mathfrak{g}^{\vee}$ given by
    $\varphi^{\xi}:=\langle\Phi_{\tilde{M}^{\tau}},\xi\rangle=-\alpha(\xi_{\tilde{M}^{\tau}})\quad(\xi\in\mathfrak{g}),$
  2. 2.
    For any $\xi\in\mathfrak{g}$, we have $\xi_{X^{\tau}}=\xi_{X^{\tau}}^{\sharp}-\varphi^{\xi}\,\mathcal{R}^{\tau}$, where $\xi_{X^{\tau}}^{\sharp}$ is a smooth section of $\mathcal{H}^{\tau}$ and we write $\varphi^{\xi}$ for $\left.\varphi^{\xi}\right|_{X^{\tau}}$.
Notation 49.
We stress the notation used above: $\tilde{\mu}^{\tau}$ is the action of $G$ on $\tilde{M}^{\tau}$ as a group of Kähler automorphisms; $\mu^{\tau}$ is the restricted action of $G$ on $X^{\tau}$ by contact CR automorphisms. Furthermore, in the following we shall generally simplify notation and write $\Phi$ for $\Phi_{\tilde{M}^{\tau}}$.
Since $\rho$ is $G$-invariant, $\xi_{\tilde{M}^{\tau_{0}}},\upsilon_{\sqrt{\rho}}]=0$, hence $\Phi$ is constant along the geodesic flow. Recalling (7) and (19), we conclude the following.
Corollary 50.
$\mathcal{R}^{\tau}$ is tangent to $Z^{\tau}$.
2.2 $\Pi^{\tau}$ in NHLC’s on $X^{\tau}$
As mentioned in §1, our computations will be carried out in so-called normal Heisenberg local coordinates on $X^{\tau}$ (recall the discussion surrounding (20) and (21)); these are induced (by projection and restriction) by suitable holomorphic local coordinates on $\tilde{M}$, in which the defining equation of $X^{\tau}$ has a canonical approximate form, and which are also called NHLC’s (on $\tilde{M}$).
NHLC’s were defined in §3.3 of [P-2024], and are a slight specialization of the normal coordinates introduced in [FS1] and [FS2], and first put to use (under the name of Heisenberg local coordinates) in the present setting in [CR1] and [CR2]. In the line bundle context, Heisenberg local coordinates have been thoroughly used in the study of scaling asymptotics since their appearance in [SZ02].
In NHLC’s on $\tilde{M}$ centered at $x\in X^{\tau}$, $\alpha$, $\Omega$ and $\hat{\kappa}$ (and hence also $\omega=\frac{1}{2}\Omega$ and $\tilde{\kappa}=\frac{1}{2}\hat{\kappa}$) have a simple local approximate expression, for which we refer to Proposition 34 of [P-2024]. Another key point in our discussion is that the phase and the leading order term of the symbol of the Szegő kernel $\Pi^{\tau}$ can be computed fairly explicitly in NHLC’s on $X^{\tau}$.
As proved in [BdM-S], $\Pi^{\tau}$ in (9) is a Fourier integral operator with complex phase. Its wave front is the anti-diagonal $(\Sigma^{\tau})^{\sharp}$ of $\Sigma^{\tau}$ in (6):
$(\Sigma^{\tau})^{\sharp}:=\left\{\left(x,r\alpha^{\tau}_{x},x,-r\alpha^{\tau}_{x}\right)\ \big{|}\ x\in X^{\tau},r>0\right\}\subseteq T^{\lor}X^{\tau}\setminus(0)\times T^{\lor}X^{\tau}\setminus(0).$
Up to a smoothing kernel, the distributional kernel $\Pi^{\tau}(\cdot,\cdot)$ has the form:
$\Pi^{\tau}(x,y)\sim\int_{0}^{+\infty}e^{\imath u\psi^{\tau}(x,y)}s^{\tau}(x,y,u)\mathrm{d}u,$ (34)
where
  1. 1.
    $s^{\tau}$ is a classical symbol of the form:
    $s^{\tau}(x,y,u)\sim\sum_{j\geq 0}u^{d-1-j}s_{j}^{\tau}(x,y);$
  2. 2.
    $\psi^{\tau}$ is a phase of positive type and is determined by:
    $\psi^{\tau}:=-\imath\tilde{\phi}^{\tau}\big{|}_{X^{\tau}\times X^{\tau}}$
    where $\tilde{\phi}$ is the holomorphic extension to $\tilde{M}\times\overline{\tilde{M}}$ of the defining function $\phi^{\tau}=\rho-\tau^{2}$ of $X^{\tau}$ in $\tilde{M}$ (see the discussions in [CR1], [CR2], and §3.3.2 of [P-2024] - and of course [BdM-S]).
For the following, see Proposition 48 and Theorem 51 of [P-2024].
Theorem 51.
[…]
$\imath\psi^{\tau}(x^{\prime},x^{\prime\prime})=\imath(\theta-\eta)-\frac{1}{4\tau^{2}}(\theta-\eta)^{2}+\psi^{\omega_{x}}_{2}(z^{\prime},z^{\prime\prime})+R_{3}(\theta,z^{\prime},\overline{z^{\prime}},\eta,z^{\prime\prime},\overline{z^{\prime\prime}}).$
Moreover, in the same chart, the principal term of the symbol satisfies
$s^{\tau}_{0}(x,x)=\frac{\tau}{(2\pi)^{d}}.$
2.3 The submanifold $Z^{\tau}$
Our analysis rests on the assumption that $G$ acts locally freely on $Z^{\tau}$ in (19). Thus Assumption 4 is certainly satisfied if $\mu$ itself is locally free and $Z^{\tau}\neq\emptyset$, for instance if $M$ is the total space of a principal $G$-bundle over a non-trivial base, $p:M\rightarrow N$ with $\dim(N)>0$. In this case, under the map $E$ in §2.1, $Z$ is the image in $\tilde{M}$ of the orthocomplement of the vertical tangent bundle $\mathrm{Ver}(p)\subseteq TM$ (that is, of the horizontal tangent bundle $\mathrm{Hor}(p)\subseteq TM$) of $M$ as a principal $G$-bundle (more precisely, $Z=E\big{(}\mathrm{Hor}(p)\big{)}\cap T^{\prime}M$). Then the quotient $Z/G$ can be identified with a tubular neighbourhood of $N$ in $TN$, and $Z^{\tau}/G$ with the bundle of tangent spheres of radius $\tau$ over $N$.
The case of a principal $G$-bundle is clearly not the only circumstance where the hypothesis is statisfied; for example, if $G$ acts in a Hamiltonian manner on a compact real-analytic symplectic manifold $(R,\omega_{R})$, and $0\in\mathfrak{g}^{\vee}$ is a regular value of the moment map $\Phi_{R}:R\rightarrow\mathfrak{g}^{\vee}$, then $G$ acts locally freely on $M:=\Phi_{R}^{-1}(0)$, but in most cases the action is not free.
Furthermore, it can happen that Assumption 4 is satisfied even if $\mu$ is not locally free. Here we give two more sufficient conditions, whose proof is left to the reader.
Lemma 52.
With the previous assumptions and notation, $\mu^{\tau}$ is locally free on $Z^{\tau}$ under the following two sets of circumstances.
  1. 1.
    $\mu$ is locally free away from a finite set of fixed points.
  2. 2.
    $G$ is Abelian, and if $T_{1},\ldots,T_{r}\leqslant G$ are the subtori that appear as stabilizers of poits in $M$ then the connected components of the fixed locus of each $T_{i}$ in $M$ are single $G$-orbits of $\mu$.
We record a statement in the opposite direction, whose proof is also left to the reader.
Lemma 53.
Suppose that there exists a submanifold $N\subseteq M$ of positive dimension such that $\mu_{g}(n)=n$ for every $g\in G$ and $n\in N$. Then $\mu^{\tau}$ is not locally free on $Z^{\tau}$.
If $G$ acts locally freely on $Z\setminus M$, the latter is a submanifold of $\tilde{M}$ of codimension $d_{G}$ (Remark 5). Given that $Z$ is obviously transverse to $X^{\tau}$, we conclude the following.
Lemma 54.
Under Assumption 4, $Z^{\tau}$ is a submanifold of $X^{\tau}$ of dimension $2d-1-d_{G}$.
Since $G$ acts locally freely on $Z^{\tau}$, $d_{G}\leq\mathrm{dim}(Z^{\tau})$.
Corollary 55.
$d_{G}\leq d-1$.
2.4 Intersections of $G$-orbits and $\mathbb{R}$-orbits
The orbits of the the geodesic flow $t\mapsto\Gamma^{\tau}_{t}$ on $X^{\tau}$ are ‘vertical’  (i.e., tangent to $\mathcal{T}^{\tau}$ in (8)); in view of (48), along $Z^{\tau}$ the orbits of the $G$-action $g\mapsto\mu^{\tau}_{g}$ are ‘horizontal’  (i.e., tangent to $\mathcal{H}^{\tau}$). Since under Assumption 4 both actions are locally free on $Z^{\tau}$ and commute, the product action of $G\times\mathbb{R}$ is also locally free on $Z^{\tau}$.
Points $x_{1},\,x_{2}\in Z^{\tau}$ belong to the same $G\times\mathbb{R}$ orbit, that is, $x_{1}\in x_{2}^{G\times\mathbb{R}}$, if and only if the $\mathbb{R}$-orbit of either one intersects the $G$-orbit of the other, i.e. $x_{1}^{G}\cap x_{2}^{\mathbb{R}}\neq\emptyset$, and then the intersection might be infinite. However, given a compact subset $K\subset\mathbb{R}$ by the above there are at most finitely many $t\in K$ for which $\Gamma^{\tau}_{t}(x_{2})\in x_{1}^{G}$. A uniform statement can be given in terms of the size of $K$, but first we need to lay down some useful consequences of the local freeness of the action of $G\times\mathbb{R}$ on the compact manifold $Z^{\tau}$; the proof is left to the reader.
Lemma 56.
Under Assumption 4, the following holds.
  1. 1.
    There exists R,r>0R,\,r>0 such that if $(g,t)$ belongs to an $R$-neighbourhood of $(e_{G},0)$ then
    $\mathrm{dist}_{X^{\tau}}\left(y,\mu^{\tau}_{g}\circ\Gamma^{\tau}_{t}(y)\right)\geq r\,\mathrm{dist}_{G\times\mathbb{R}}\big{(}(g,t),(e_{G},0)\big{)}.$
  2. 2.
    There exist $C>0$ such that
    $\mathrm{dist}_{X^{\tau}}\left(x,\Gamma^{\tau}_{t}(x)\right)<C\,|t|,\quad\forall\,(x,t)\in Z^{\tau}\times\mathbb{R}.$
  3. 3.
    There exists $D>0$ such that for any $y\in Z^{\tau}$ and $\delta>0$ sufficiently small
    $\mathrm{dist}_{G}(g,G_{y})\geq\delta\quad\Rightarrow\quad\mathrm{dist}_{X^{\tau}}\left(y,\mu^{\tau}_{g}(y)\right)\geq\frac{\delta}{D}.$
Lemma 57.
Assume that $\epsilon>0$ is sufficiently small and $\chi\in\mathrm{C}^{\infty}_{c}\big{(}(t_{0}-\epsilon,t_{0}+\epsilon)\big{)}$ for some $t_{0}\in\mathbb{R}$. Then for any $x\in Z^{\tau}$ and $x^{\prime}\in x^{G\times\chi}$ there exists a unique $t\in\mathrm{supp}(\chi)$ such that $x^{\prime}=\mu^{\tau}_{g}\circ\Gamma^{\tau}_{t}(x)$ for some $g\in G$.
Proof of Lemma 57.
If the Lemma is false, for any $j=1,2,\ldots$ there exist $x_{j},\,x_{j}^{\prime}\in Z^{\tau}$, $t_{j},\,t_{j}^{\prime}\in\mathbb{R}$, and $g_{j},\,g_{j}^{\prime}\in G$ such that such that
$0<|t_{j}-t_{j}^{\prime}|<\frac{1}{j}\quad\text{and}\quad x_{j}^{\prime}=\mu^{\tau}_{g_{j}^{\prime}}\circ\Gamma^{\tau}_{t_{j}^{\prime}}(x_{j})=\mu^{\tau}_{g_{j}}\circ\Gamma^{\tau}_{t_{j}}(x_{j}).$
Hence
$x_{j}=\mu^{\tau}_{g_{j}^{-1}\,g_{j}^{\prime}}\circ\Gamma^{\tau}_{t_{j}^{\prime}-t_{j}}(x_{j}).$ (35)
By statement 2 of Lemma 56,
$\mathrm{dist}_{X^{\tau}}\left(x,\Gamma^{\tau}_{t_{j}^{\prime}-t_{j}}(x)\right)<C/j\quad\forall\,x\in X^{\tau}.$
By the triangle inequality, we conclude that
$\mathrm{dist}_{X^{\tau}}\left(x_{j},\mu^{\tau}_{g_{j}^{-1}\,g_{j}^{\prime}}(x_{j})\right)$
$\leq$ $\mathrm{dist}_{X^{\tau}}\left(x_{j},\mu^{\tau}_{g_{j}^{-1}\,g_{j}^{\prime}}\circ\Gamma^{\tau}_{t_{j}^{\prime}-t_{j}}(x_{j})\right)+\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g_{j}^{-1}\,g_{j}^{\prime}}\circ\Gamma^{\tau}_{t_{j}^{\prime}-t_{j}}(x_{j}),\mu^{\tau}_{g_{j}^{-1}\,g_{j}^{\prime}}(x_{j})\right)$
$\leq$ $0+C/j=C/j.$
Hence, by statement 3 of Lemma 56,
$\mathrm{dist}_{G}\left(g_{j}^{-1}\,g_{j}^{\prime},G_{x_{j}}\right)\leq(C\cdot D)/j.$
We may thus find $\sigma_{j}\in G_{x_{j}}$ and $\delta_{j}\in G$ such that
$g_{j}^{-1}\,g_{j}^{\prime}=\delta_{j}\,\sigma_{j},\quad\mathrm{dist}_{G}\left(\delta_{j},e_{G}\right)\leq D/j.$
Then (35) implies
$x_{j}=\mu^{\tau}_{\delta_{j}\,\sigma_{j}}\circ\Gamma^{\tau}_{t_{j}^{\prime}-t_{j}}(x_{j})=\mu^{\tau}_{\delta_{j}}\circ\Gamma^{\tau}_{t_{j}^{\prime}-t_{j}}(x_{j}),$ (37)
and $(\delta_{j},t_{j}^{\prime}-t_{j})$ belongs to a neighbourhood of $(0,e_{G})$ of radius $O\left(1/j\right)$. By statement 1 of Lemma 56 this is absurd, unless $\delta_{j}=e_{G}$ and $t_{j}=t_{j}^{\prime}$.
Corollary 58.
If $\epsilon>0$ is sufficiently small and $|\mathrm{supp}(\chi)|<2\,\epsilon$, then for any $x_{1},\,x_{2}\in Z^{\tau}$ we have the following alternative; either $x_{1}^{G}\cap x_{2}^{\chi}=\emptyset$, or else $x_{1}^{G}\cap x_{2}^{\chi}=\{x_{12}\}$ for a unique $x_{12}\in Z^{\tau}$.
Remark 59.
In the assumptions of Corollary 58, suppose that $x_{1}^{G}\cap x_{2}^{\mathrm{supp}(\chi)}\neq\emptyset$; equivalently,
$\Sigma_{\chi}(x_{1},x_{2}):=\left\{(g,t)\in G\times\mathrm{supp}(\chi)\,:\,x_{1}=\mu^{\tau}_{g}\circ\Gamma^{\tau}_{t}(x_{2})\right\}\neq\emptyset.$ (38)
Let $t_{1}=t_{1}(x_{1},x_{2})\in\mathrm{supp}(\chi)$ be the unique element whose existence is asserted in Lemma 57, and choose $h_{1}\in G$ such that $(h_{1},t_{1})\in\Sigma_{\chi}(x_{1},x_{2})$. Let $G_{x_{1}}:=\left\{\kappa_{l}\,:\,l=1,\ldots,r_{x_{1}}\right\}\leqslant G$ denote the stabilizer subgroup of $x_{1}$. Then
$\Sigma_{\chi}(x_{1},x_{2})=\left\{(\kappa_{l}\,h_{1},t_{1})\,:\,l=1,\ldots,r_{x_{1}}\right\}.$ (39)
2.5 The action in Heisenberg local coordinates
As is the Introduction (recall the discussion surrounding (21)), NHLC’s on $X^{\tau}$ centered at $x$ will be denoted in additive notation: $y(\theta,\mathbf{v})=x+(\theta,\mathbf{v})$. Given $x\in X^{\tau}$, we can find an open neighbourhood $U^{\tau}\subseteq X^{\tau}$ of $x$ and a smoothly varying family of systems of NHLC’s $y_{x^{\prime}}(\theta,\mathbf{v})=x^{\prime}+(\theta,\mathbf{v})$ centered at $x^{\prime}\in U^{\tau}$. More explicitly, for some $\delta>0$ we have a smooth map $\Upsilon:U^{\tau}\times(-\delta,\delta)\times B_{2d-2}(\mathbf{0},\delta)\rightarrow X^{\tau}$ such that for every $x^{\prime}\in U^{\tau}$ the partial function $\Upsilon(x^{\prime},\cdot,\cdot)$ is a system of NHLC’s centered at $x^{\prime}$. Hence
$y_{x^{\prime}}(\theta,\mathbf{v})=\Upsilon(x^{\prime},\theta,\mathbf{u})=x^{\prime}+(\theta,\mathbf{u}).$
We refer to the notation and conventions in [P-2024] (see also [CR1] and [CR2]). In this section, we shall adapt some arguments from [P2-2012] to express the action of $G\times\mathbb{R}$ on $X^{\tau}$ in terms of NHLC’s.
Let us first consider the geodesic flow. The following is a consequence of (105) of [P-2024].
Lemma 60.
Suppose $x\in X^{\tau}$, and let us choose Heisenberg local coordinates on $X^{\tau}$ centered at $x$. If $y=x+(\theta,\mathbf{u})$, then
$\Gamma^{\tau}_{t}(y)=x+\Big{(}\theta-\tau\,t+R_{3}(\tau,t,\theta,\mathbf{u}),\mathbf{u}+\mathbf{R}_{2}(\tau\,t,\theta,\mathbf{u})\Big{)}.$
The convention for the meaning of $R_{k}$ and $\mathbf{R}_{k}$ is as explained in the discussion of (23).
To proceed, we need to relate systems of NHLC’s centered at nearby points. We shall adapt an argument in [P2-2012] to the present setting.
Lemma 61.
For $(\theta_{,}\mathbf{u}_{j})\sim(0,\mathbf{0})$,
$x+(\theta_{2},\mathbf{u}_{2})$ $=$ $\big{(}x+(\theta_{1},\mathbf{u}_{1})\big{)}$
$+\Big{(}\theta_{2}-\theta_{1}+\omega(\mathbf{u}_{1},\mathbf{u}_{2})+R_{3}(\theta_{j},\mathbf{u}_{j}),\mathbf{u}_{2}-\mathbf{u}_{1}+\mathbf{R}_{2}(\theta_{j},\mathbf{u}_{j})\Big{)}.$
Proof.
We can write
$x+(\theta_{2},\mathbf{u}_{2})=\big{(}x+(\theta_{1},\mathbf{u}_{1})\big{)}+\big{(}\beta(\theta_{j},\mathbf{u}_{j}),\mathbf{B}(\theta_{j},\mathbf{u}_{j})\big{)}$
for certain smooth functions $\beta$ (real-valued) and $\mathbf{B}$ (vector-valued), vanishing at the origin. Let us expand $\beta$ and $\mathbf{B}$:
$\beta=\beta_{1}+\beta_{2}+R_{3},\quad\mathbf{B}=\mathbf{B}_{1}+\mathbf{R}_{2},$
where $\beta_{1}$ and $\mathbf{B}_{1}$ are linear, and $\beta_{2}$ is homogeneous of degree $2$. It is easily seen that
$\beta_{1}(\theta_{j},\mathbf{u}_{j})=\theta_{2}-\theta_{1},\quad\mathbf{B}_{1}(\theta_{j},\mathbf{u}_{j})=\mathbf{u}_{2}-\mathbf{u}_{1};$ (40)
let us then determine $\beta_{2}$.
Let us choose $\chi\in\mathcal{C}^{\infty}_{0}\big{(}[-\epsilon,\epsilon]\big{)}$ for some sufficiently small $\epsilon>0$. Keeping $\theta_{j},\mathbf{u}_{j}$ fixed, for $\lambda>0$ let $x_{j,\lambda}$ be as in (27) and apply Theorem 1 in [P-2024].
On the one hand, for $\lambda\rightarrow+\infty$ we have
$\Pi_{\chi,\lambda}(x_{1,\lambda},x_{2,\lambda})$
$=$ $\frac{\chi(0)}{\sqrt{2\,\pi}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1}\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})+\psi_{2}(\mathbf{u}_{1},\mathbf{u}_{2})\right]}\cdot\left[1+O\left(\lambda^{-1/2}\right)\right].$
On the other hand, by (40) we have
$x_{2,\lambda}$ $=$ $x_{1,\lambda}+\left(\frac{1}{\sqrt{\lambda}}\,\left(\theta_{2}-\theta_{1}+\frac{1}{\sqrt{\lambda}}\,\beta_{2}(\theta_{j},\mathbf{u}_{j})+\sqrt{\lambda}\,\beta_{3}\left(\frac{\theta_{j}}{\sqrt{\lambda}},\,\frac{\mathbf{u}_{j}}{\sqrt{\lambda}}\right)\right),\right$. (42)
$\left.\frac{1}{\sqrt{\lambda}}\,(\mathbf{u}_{2}-\mathbf{u}_{1})+\mathbf{R}_{3}\left(\frac{\theta_{j}}{\sqrt{\lambda}},\frac{\mathbf{u}_{j}}{\sqrt{\lambda}}\right)\right)$
Hence, again by Theorem 1 in [P-2024],
$\Pi_{\chi,\lambda}(x_{1,\lambda},x_{2,\lambda})$
$=$ $\frac{\chi(0)}{\sqrt{2\,\pi}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1}\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})-\imath\,\beta_{2}(\theta_{j},\mathbf{u}_{j})-\frac{1}{2}\,\|\mathbf{u}_{1}-\mathbf{u}_{2}\|^{2}\right]}\cdot\left[1+O\left(\lambda^{-1/2}\right)\right].$
Thus
$\imath\,\beta_{2}(\theta_{j},\mathbf{u}_{j})-\frac{1}{2}\,\|\mathbf{u}_{1}-\mathbf{u}_{2}\|^{2}=\psi_{2}(\mathbf{u}_{1},\mathbf{u}_{2})=-\imath\,\omega_{x}(\mathbf{u}_{1},\mathbf{u}_{2})-\frac{1}{2}\,\|\mathbf{u}_{1}-\mathbf{u}_{2}\|^{2}.$
The claim follows. ∎
Corollary 62.
Under the same assumptions,
$\Big{(}x+(\theta,\mathbf{u})\Big{)}+\big{(}\beta,\mathbf{B}\big{)}$
$=$ $x+\Big{(}\theta+\beta-\omega_{x}(\mathbf{u},\mathbf{B})+R_{3},\mathbf{u}+\mathbf{B}+\mathbf{R}_{2}\Big{)},$
where $R_{3}=R_{3}(\theta,\beta,\mathbf{u},\mathbf{B})$, $\mathbf{R}_{2}=\mathbf{R}_{2}(\theta,\beta,\mathbf{u},\mathbf{B})$.
Let us now express the $G$-action near a point $x\in X^{\tau}$ in NHLC’s. We consider both the case of a fixed $g\in G_{x}$ (the stabilizer subgroup of $x$), and that of a 1-parameter subgroup of $G$ generated by a given $\xi\in\mathfrak{g}$. Let us premise a remark. If $g\in G_{x}$, then $\mathrm{d}_{x}\mu^{\tau}_{g}:T_{x}X^{\tau}\rightarrow T_{x}X^{\tau}$ preserves both the vertical and horizontal tangent vector bundles at $x$; furthermore, $\mathrm{d}_{x}\mu^{\tau}_{g}\big{(}\mathcal{R}^{\tau}(x)\big{)}=\mathcal{R}^{\tau}(x)$. If $\mathbf{v}\in\mathcal{H}^{\tau}(x)$, we shall simplify notation and write
$\mathbf{v}_{g}:=\mathrm{d}_{x}\mu^{\tau}_{g}\big{(}\mathbf{v}\big{)}.$ (44)
Lemma 63.
Suppose $x\in X^{\tau}$ and choose a system of NHLC’s centered at $x$. Then the following holds.
  1. 1.
    If $\kappa\in G_{x}$,
    $\mu^{\tau}_{\kappa^{-1}}\big{(}x+(\theta,\mathbf{v})\big{)}=x+\Big{(}\theta+R_{3}(\theta,\mathbf{v}),\mathbf{v}_{\kappa^{-1}}+\mathbf{R}_{2}(\theta,\mathbf{v})\Big{)}.$
  2. 2.
    If $\xi\sim 0\in\mathfrak{g}$ and $(\theta,\mathbf{v})\sim(0,\mathbf{0})\in\mathbb{R}\times\mathbb{R}^{2d-2}$, then
    $\mu^{\tau}_{e^{-\xi}}\big{(}x+(\theta,\mathbf{v})\big{)}=x+\Big{(}\theta+\varphi^{\xi}(x)+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}\big{)}+R_{3},\mathbf{v}-\xi^{\sharp}_{X^{\tau}}(x)+\mathbf{R}_{2}\Big{)},$
    where $R_{3}=R_{3}(\theta,\mathbf{v},\xi)$, $\mathbf{R}_{2}=\mathbf{R}_{2}(\theta,\mathbf{v},\xi)$.
Remark 64.
If $\mathbf{v}\in\mathcal{H}^{\tau}(x)$, then $\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}\big{)}=\omega_{x}\big{(}\xi_{X^{\tau}}(x),\mathbf{v}\big{)}$.
Proof of 1.
Composing the given system of NHLC’s with $\mu^{\tau}_{\kappa^{-1}}$ yields another system of NHLC’s on $X^{\tau}$ centered at $x$. Hence, by the discussion in §3.3.5 of [P-2024], the angular coordinates of the two systems only differ to third order. ∎
Proof of 2..
With the notation of Corollary 48, since $\mu^{\tau}$ and the geodesic flow commute
$\left[\xi^{\sharp}_{X^{\tau}},\varphi^{\xi}\,\mathcal{R}^{\tau}\right]=\left[\xi^{\sharp}_{X^{\tau}},\mathcal{R}^{\tau}\right]=0.$
Let us denote by $\Upsilon_{t,V}$ the flow at time $t$ of a smooth vector field $V$ on $X^{\tau}$. Then, if $\mu^{\tau}_{t,\xi}:=\Upsilon_{t,\xi_{X^{\tau}}}$, for any $t\in\mathbb{R}$:
$\mu^{\tau}_{-t,\xi}=\Upsilon_{t,\varphi^{\xi}\,\mathcal{R}}\circ\Upsilon_{-t,\xi^{\sharp}_{X^{\tau}}}:X^{\tau}\rightarrow X^{\tau}.$
For any $x^{\prime}\in X^{\tau}$, the curve $t\mapsto\Upsilon_{-t,\xi^{\sharp}_{X^{\tau}}}(x)$ is horizontal, since $\xi^{\sharp}_{X^{\tau}}$ is tangent to $\mathcal{H}^{\tau}$ (here ‘horizontal’  is the sense of (8)). By §3.3.7 of [P-2024],
$\Upsilon_{-t,\xi^{\sharp}_{X^{\tau}}}\big{(}x+(\theta,\mathbf{v})\big{)}$
$=$ $\big{(}x+(\theta,\mathbf{v})\big{)}+\Big{(}R_{3}(t),-t\,\xi^{\sharp}_{X^{\tau}}(x)+\mathbf{R}_{2}(t)\Big{)}.$
In view of Lemma 61, this may be rewritten
$\Upsilon_{-t,\xi^{\sharp}_{X^{\tau}}}\big{(}x+(\theta,\mathbf{v})\big{)}$
$=$ $x+\Big{(}\theta+t\,\omega_{x}(\mathbf{v},\xi^{\sharp}_{X^{\tau}}(x))+R_{3},\mathbf{v}-t\,\xi^{\sharp}_{X^{\tau}}(x)+\mathbf{R}_{2}\Big{)},$
where we abridge $R_{3}=R_{3}(\theta,t,\mathbf{v})$, and similarly for $\mathbf{R}_{2}$. Then by Lemma 60
$\mu^{\tau}_{-t,\xi}\big{(}x+(\theta,\mathbf{v})\big{)}$
$=$ $\Upsilon_{t,\varphi^{\xi}\,\mathcal{R}}\Big{(}x+\Big{(}\theta+t\,\omega_{x}(\mathbf{v},\xi^{\sharp}_{X^{\tau}}(x))+R_{3},\mathbf{v}-t\,\xi^{\sharp}_{X^{\tau}}(x)+\mathbf{R}_{2}\Big{)}\Big{)}$
$=$ $x+\Big{(}\theta+t\,\omega_{x}(\mathbf{v},\xi^{\sharp}_{X^{\tau}}(x))+t\,\varphi^{\xi}\big{(}x+(\theta,\mathbf{v})\big{)}+R_{3},\mathbf{v}-t\,\xi^{\sharp}_{X^{\tau}}(x)+\mathbf{R}_{2}\Big{)}$
$=$ $x+\Big{(}\theta+t\,\omega_{x}(\mathbf{v},\xi^{\sharp}_{X^{\tau}}(x))+t\,\varphi^{\xi}(x)+2\,t\,\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp},\mathbf{v}\big{)}+R_{3},\mathbf{v}-t\,\xi^{\sharp}_{X^{\tau}}(x)+\mathbf{R}_{2}\Big{)}$
$=$ $x+\Big{(}\theta+t\,\varphi^{\xi}(x)+t\,\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp},\mathbf{v}\big{)}+R_{3},\mathbf{v}-t\,\xi^{\sharp}_{X^{\tau}}(x)+\mathbf{R}_{2}\Big{)}.$
Corollary 65.
In the situation and with the notation of Lemma 63, if in addition $x\in Z^{\tau}$ then
$\mu^{\tau}_{e^{-\xi}}\big{(}x+(\theta,\mathbf{v})\big{)}=x+\Big{(}\theta+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}\big{)}+R_{3},\mathbf{v}-\xi^{\sharp}_{X^{\tau}}(x)+\mathbf{R}_{2}\Big{)}.$
We shall need to localize our computations in $X^{\tau}$ near points $x\in Z^{\tau}$ and in $G$ near $G_{x}$. Let $G_{x}=\{\kappa_{1},\ldots,\kappa_{r}\}$, and for each $l=1,\ldots,r$ let us parametrize $G$ in the neighbourhood of $\kappa_{l}$ by setting
$g=e^{\xi}\,\kappa_{l},\quad\text{where}\quad\xi\sim 0\in\mathfrak{g}.$ (46)
Furthermore, we shall further abridge notation (44), and set
$\mathbf{v}^{(l)}:=\mathbf{v}_{k_{l}^{-1}}.$ (47)
Then, by Corollary 65,
$\mu^{\tau}_{g^{-1}}\big{(}x+(\theta,\mathbf{v})\big{)}$ $=$ $\mu^{\tau}_{\kappa_{l}^{-1}}\circ\mu^{\tau}_{e^{-\xi}}\big{(}x+(\theta,\mathbf{v})\big{)}$
$=$ $\mu^{\tau}_{\kappa_{l}^{-1}}\left(x+\Big{(}\theta+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}\big{)}+R_{3},\mathbf{v}-\xi^{\sharp}_{X^{\tau}}(x)+\mathbf{R}_{2}\Big{)}\right).$
Applying 1. of Lemma 63, we obtain (with $g$ as in (46))
$\mu^{\tau}_{g^{-1}}\big{(}x+(\theta,\mathbf{v})\big{)}$
$=$ $x+\Big{(}\theta+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}\big{)}+R_{3},\mathbf{v}^{(l)}-\xi^{\sharp}_{X^{\tau}}(x)^{(l)}+\mathbf{R}_{2}\Big{)}.$
Remark 66.
For any $\xi\in\mathfrak{g}$ and $l=1,\ldots,r$,
$\xi^{\sharp}_{X^{\tau}}(x)^{(l)}=\mathrm{Ad}_{\kappa_{l}^{-1}}(\xi)_{X^{\tau}}(x).$
2.6 $\psi^{\tau}$ near orbit intersections
Let us now assume that $x_{1},x_{2}\in Z^{\tau}$ are such that $x_{1}^{G}\cap x_{2}^{\chi}\neq\emptyset$, i.e. $\Sigma_{\chi}(x_{1},x_{2})\neq\emptyset$ (recall (38)). Thus, assuming $|\mathrm{supp}(\chi)|$ is sufficiently small, $\Sigma_{\chi}(x_{1},x_{2})$ is as in (39). By Corollary 58, $x_{1}^{G}\cap x_{2}^{\chi}=\{x_{12}\}$, where
$x_{12}:=\mu^{\tau}_{h_{1}^{-1}}(x_{1})=\Gamma^{\tau}_{t_{1}}(x_{2}).$
We shall need to expand the phase $\psi^{\tau}$ in (34) near $(x_{12},x_{12})$ and $(x_{2},x_{2})$. We assume given systems of NHLC’s on $X^{\tau}$ centered at $x_{1}$ and $x_{2}$. The choice of $h_{1}$ then uniquely determines NHLC’s on $X^{\tau}$ centered at $x_{12}$, by the condition that $\mu^{\tau}_{h_{1}^{-1}}$ be locally represented by the identity from a neighbourhood of $x_{1}$ to one of $x_{12}$.
2.6.1 $\psi^{\tau}$ near $(x_{12},x_{12})$
In the neighbourhood of $\kappa_{l}\,h_{1}$ (recall (39)), we can parameterize $G$ by setting
$g:=e^{\xi}\,\kappa_{l}\,h_{1},\quad\text{where}\quad\xi\sim 0\in\mathfrak{g}.$
Making use of (2.5), we obtain
$\mu^{\tau}_{g^{-1}}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1})\big{)}=\mu^{\tau}_{h_{1}^{-1}}\circ\mu^{\tau}_{\kappa_{l}^{-1}\,e^{-\xi}}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1})\big{)}$
$=$ $\mu^{\tau}_{h_{1}^{-1}}\left(x_{1}+\Big{(}\theta_{1}+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}_{1}\big{)}+R_{3},\mathbf{v}_{1}^{(l)}-\xi^{\sharp}_{X^{\tau}}(x)^{(l)}+\mathbf{R}_{2}\Big{)}\right)$
$=$ $x_{12}+\Big{(}\theta_{1}+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}_{1}\big{)}+R_{3},\mathbf{v}^{(l)}_{1}-\xi^{\sharp}_{X^{\tau}}(x)^{(l)}+\mathbf{R}_{2}\Big{)}.$
In the neighbourhood of $(x_{12},x_{12})$, we shall consider pairs of the form
$\left(\mu^{\tau}_{g^{-1}}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1}\big{)},x_{12}+\big{(}\theta,\mathbf{v}\big{)}\right),\quad g=e^{\xi}\,\kappa_{l}\,h_{1};$
the first entry is given by (2.6.1). By the discussion in §2.2 (see also Proposition 47 of [P-2024]), we then have:
$\imath\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}\big{(}x_{1}+(\theta_{1},\mathbf{v}_{1})\big{)},x_{12}+\big{(}\theta,\mathbf{v}\big{)}\right)$
$=$ $\imath\,\psi^{\tau}\left(x_{12}+\Big{(}\theta_{1}+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}_{1}\big{)}+R_{3},\mathbf{v}^{(l)}_{1}-\xi^{\sharp}_{X^{\tau}}(x)^{(l)}+\mathbf{R}_{2}\Big{)},x_{12}+\big{(}\theta,\mathbf{v}\big{)}\right)$
$=$ $\imath\,\Big{(}\theta_{1}-\theta+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}_{1}\big{)}\Big{)}-\frac{1}{4\,\tau^{2}}\,\Big{(}\theta_{1}-\theta+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}_{1}\big{)}\Big{)}^{2}$
$+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi^{\sharp}_{X^{\tau}}(x)^{(l)},\mathbf{v}\right)+R_{3}$
$=$ $\imath\,\Big{(}\theta_{1}-\theta+\omega_{x}\big{(}\xi_{X^{\tau}}^{\sharp}(x),\mathbf{v}_{1}\big{)}\Big{)}-\frac{1}{4\,\tau^{2}}\,(\theta_{1}-\theta)^{2}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi^{\sharp}_{X^{\tau}}(x)^{(l)},\mathbf{v}\right)+R_{3},$
where $R_{3}=R_{3}(\theta_{1},\theta,\mathbf{v}_{1},\mathbf{v},\xi)$.
2.6.2 $\psi^{\tau}$ near $(x_{2},x_{2})$
In the neighbourhood of $(x_{2},x_{2})$, we shall consider points of the form
$\left(\Gamma^{\tau}_{-t_{1}-t}\big{(}x_{12}+(\theta,\mathbf{v})\big{)},x_{2}+(\theta_{2},\mathbf{v}_{2})\right),$ (52)
where $t\sim 0$; recall that $\Gamma^{\tau}_{-t_{1}}(x_{12})=x_{2}$.
We aim to express the first entry in (52) in NHLC’s on $X^{\tau}$ centered at $x_{2}$. Since $\alpha^{\tau}$ is invariant by the geodesic flow, $\Gamma^{\tau}_{t}$ preserves the vertical and horizontal tangent bundles. In particular, for any $s\in\mathbb{R}$ and $\mathbf{v}\in\mathcal{H}^{\tau}_{x_{12}}$ with some abuse of notation we have
$\mathrm{d}_{x_{12}}\Gamma^{\tau}_{-t_{1}}\left(s\,\mathcal{R}^{\tau}(x_{12})+\mathbf{v})\right)=s\,\mathcal{R}^{\tau}(x_{2})+B\,\mathbf{v},$ (53)
where the symplectic matrix $B$ is as in (25).
Lemma 67.
Under the previous assumptions,
$\Gamma^{\tau}_{-t_{1}}\big{(}x_{12}+(\theta,\mathbf{v})\big{)}$ $=$ $x_{2}+\big{(}\theta+R_{3}(\theta,\mathbf{v}),B\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v})\big{)}.$
Before proving Lemma 67, let us give a definition.
Definition 68.
Let $\gamma:(-a,a)\rightarrow X^{\tau}$ be a smooth curve defined for some $a>0$ and set $\gamma(0)=x$. Let $k\geq 1$ be an integer. The $\gamma$ will be said to be horizontal to $k$-th order at $x$ if $\left\langle\alpha,\dot{\gamma}\right\rangle=O\left(t^{k}\right)$ for $t\sim 0$.
Lemma 69.
In the situation of Definition 68, the following conditions are equivalent:
  1. 1.
    $\gamma$ is horizontal to second order at $x$, that is, $\left\langle\alpha,\dot{\gamma}\right\rangle=O\left(t^{2}\right)$ for $t\sim 0$;
  2. 2.
    given any system of NHLC’s on $X^{\tau}$ centered at $x$, for an appropriate $\mathbf{v}\in\mathcal{H}^{\tau}_{x}$ we have
    $\gamma(t)=x+\big{(}R_{3}(t),t\,\mathbf{v}+\mathbf{R}_{2}(t)\big{)}.$
Proof.
NHLC’s $(\theta,z^{\prime})$ on $X^{\tau}$ centered at $x$ are induced, under projection and restriction, by NHLC’s $(z_{0},z_{1},\ldots,z_{d-1})$, on $\tilde{M}$ adapted to $X^{\tau}$ and centered at $x$ (see §2.2); here $z_{0}=\theta+\imath\,\eta_{0}$ (with $\theta$ and $\eta_{0}$ real). By Proposition 33 of [P-2024],
$\alpha$ $=$ $\mathrm{d}\theta_{0}+\frac{1}{2\,\imath}\,\left[\frac{1}{2\,\tau^{2}}\,\left(\overline{z}_{0}\,\mathrm{d}z_{0}-z_{0}\,\mathrm{d}\overline{z}_{0}\right)+\overline{z}^{\prime}\cdot\mathrm{d}z^{\prime}-z^{\prime}\cdot\mathrm{d}\overline{z}^{\prime}\right]$
$+R_{2}(z,\overline{z}).$
If $\gamma(t)=\big{(}\theta(t),z^{\prime}(t)\big{)}$ in local coordinates on $X^{\tau}$, then $\gamma(t)=\big{(}\theta(t)+\imath\,\eta_{0}(t),z^{\prime}(t)\big{)}$ in the corresponding NHLC’s on $\tilde{M}$. In particular, $z_{0}(t)=\theta(t)+\imath\,\eta_{0}(t)$.
By Corollary 35 of [P-2024], $\eta_{0}(t)=R_{2}(t)$. Hence
$\gamma^{*}\left(\overline{z}_{0}\,\mathrm{d}z_{0}-z_{0}\,\mathrm{d}\overline{z}_{0}\right)=2\,\imath\,\gamma^{*}(\theta\,\mathrm{d}\eta_{0}-\eta_{0}\,\mathrm{d}\theta)=R_{2}(t).$
If $z^{\prime}(t)=t\,w+\mathbf{R}_{2}(t)$, then
$\gamma^{*}\left(\overline{z}^{\prime}\cdot\mathrm{d}z^{\prime}-z^{\prime}\cdot\mathrm{d}\overline{z}^{\prime}\right)$
$=$ $\left[\left(t\,\overline{w}+\mathbf{R}_{2}(t)\right)\cdot\big{(}w+R_{1}(t)\big{)}-\left(t\,w+\mathbf{R}_{2}(t)\right)\cdot\big{(}\overline{w}+R_{1}(t)\big{)}\right]\,\mathrm{d}t$
$=$ $R_{2}(t).$
Hence, $\gamma^{*}(\alpha)=\theta^{\prime}(t)\,\mathrm{d}t+R_{2}(t)$. Thus, $\left\langle\alpha,\dot{\gamma}\right\rangle=R_{2}(t)$ if and only if $\theta^{\prime}(t)=R_{2}(t)$ if and only if $\theta(t)=R_{3}(t)$.
Proof of Lemma 67.
By Lemma 60, we have
$x_{12}+(\theta,\mathbf{v})=\Gamma^{\tau}_{-\theta/\tau}\Big{(}x_{12}+\big{(}R_{3}(\theta,\mathbf{v}),\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v})\big{)}\Big{)},$
for certain functions $R_{3}$ and $\mathbf{R}_{2}$ (with the usual conventions about vanishing orders, and with $R_{j}$, $\mathbf{R}_{j}$ allowed to vary from line to line). By (53), we can write
$\Gamma_{-t_{1}}\Big{(}x_{12}+\big{(}R_{3}(\theta,\mathbf{v}),\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v})\big{)}\Big{)}=x_{2}+\big{(}R_{2}^{(1)}(\theta,\mathbf{v}),B\mathbf{v}+\mathbf{R}^{(1)}_{2}(\theta,\mathbf{v})\big{)},$
for certain functions vanishing to the indicated orders.
By Lemma 69, the smooth curve $\gamma:(-a,a)\rightarrow X^{\tau}$, defined for some $a\geq 1$ by
$\gamma(t):=x_{12}+\big{(}R_{3}(t\,\theta,t\,\mathbf{v}),\mathbf{v}+\mathbf{R}_{2}(t\,\theta,t\,\mathbf{v})\big{)},$
is horizontal to second order at $x_{12}$. Since the geodesic flow $\Gamma^{\tau}$ preserves $\alpha^{\tau}$, $\Gamma^{\tau}_{-t_{1}}\circ\gamma$ is horizontal to second order at $x_{2}$. Hence $R_{2}^{(1)}(t\,\theta,t\,\mathbf{v})=O\left(t^{3}\right)$, and therefore $R_{2}^{(1)}(\theta,\mathbf{v})$ really vanishes to third order at $(0,\mathbf{0})$; we shall accordingly replace it by $R_{3}^{(1)}(\theta,\mathbf{v})$.
Therefore,
$\Gamma^{\tau}_{-t_{1}}\big{(}x_{12}+(\theta,\mathbf{v})\big{)}$ $=$ $\Gamma^{\tau}_{-t_{1}}\circ\Gamma^{\tau}_{-\theta/\tau}\Big{(}x_{12}+\big{(}R_{3}(\theta,\mathbf{v}),\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v})\big{)}\Big{)}$
$=$ $\Gamma^{\tau}_{-\theta/\tau}\circ\Gamma^{\tau}_{-t_{1}}\Big{(}x_{12}+\big{(}R_{3}(\theta,\mathbf{v}),\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v})\big{)}\Big{)}$
$=$ $\Gamma^{\tau}_{-\theta/\tau}\left(x_{2}+\big{(}R_{3}^{(1)}(\theta,\mathbf{v}),B\mathbf{v}+\mathbf{R}^{(1)}_{2}(\theta,\mathbf{v})\big{)}\right)$
$=$ $x_{2}+\Big{(}\theta+R_{3}^{(2)}(\theta,\mathbf{v}),B\mathbf{v}+\mathbf{R}^{(2)}_{2}(\theta,\mathbf{v})\big{)},$
where in the last equality we have made use of Lemma 60. ∎
Using one more time Lemma 60, we obtain the following upshot.
Corollary 70.
With the previous notation,
$\Gamma^{\tau}_{-t-t_{1}}\Big{(}x_{12}+(\theta,\mathbf{v})\Big{)}$
$=$ $x_{2}+\Big{(}\theta+\tau\,t+R_{3}(t,\theta,\mathbf{v}),B\mathbf{v}+\mathbf{R}_{2}(t,\theta,\mathbf{v})\Big{)}.$
Invoking again Proposition 47 of [P-2024] (see also §2.2 above), we conclude that
$\imath\,\psi^{\tau}\left(\Gamma^{\tau}_{-t_{1}-t}\big{(}x_{12}+(\theta,\mathbf{v})\big{)},x_{2}+(\theta_{2},\mathbf{v}_{2})\right)$
$=$ $\imath\,\psi^{\tau}\left(x_{2}+\Big{(}\theta+\tau\,t+R_{3}(t,\theta,\mathbf{v}),B\mathbf{v}+\mathbf{R}_{2}(t,\theta,\mathbf{v})\Big{)},x_{2}+(\theta_{2},\mathbf{v}_{2})\right)$
$=$ $\imath\,(\theta+\tau\,t-\theta_{2})-\frac{1}{4\,\tau^{2}}\,(\theta+\tau\,t-\theta_{2})^{2}+\psi_{2}\big{(}B\mathbf{v},\mathbf{v}_{2}\big{)}+R_{3}(\theta,\mathbf{v},\theta_{2},\mathbf{v}_{2}).$
3 Proof of Theorem 7 (case of $\Pi^{\tau}_{\chi,\nu,\lambda}$)
3.1 Preamble
In this section, we shall prove Theorem 7 for $\Pi^{\tau}_{\chi,\nu,\lambda}$; the formal changes to the argument needed for $P^{\tau}_{\chi,\nu,\lambda}$ will be described in §5. Before giving the proof, we shall lay down some preliminaries and pieces of notation.
  1. 1.
    With $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$ as in (10), we shall denote by
    $U^{\tau}_{\sqrt{\rho}}(t):=e^{\imath\,t\,\mathfrak{D}^{\tau}_{\sqrt{\rho}}}\qquad(t\in\mathbb{R})$
    the $1$-parameter group of unitary Toeplitz operators generated by $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$.
  2. 2.
    The distributional kernel of $\Pi^{\tau}_{\chi,\lambda}$ in (13) is related to that of $U^{\tau}_{\sqrt{\rho}}(t)$ by
    $\Pi^{\tau}_{\chi,\lambda}(x,y)=\frac{1}{\sqrt{2\,\pi}}\,\int_{-\infty}^{+\infty}e^{-\imath\,\lambda\,t}\,\chi(t)\,U^{\tau}_{\sqrt{\rho}}(t;x,y)\,\mathrm{d}t.$ (55)
  3. 3.
    For $t\in\mathbb{R}$, let us denote by $\Pi^{\tau}_{-t}:L^{2}(X^{\tau})\rightarrow L^{2}(X_{\tau})$ the operator having distributional kernel
    $\Pi^{\tau}_{-t}(x,y):=\Pi^{\tau}\left(\Gamma^{\tau}_{-t}(x),y\right).$
  4. 4.
    Following Zelditch (see e.g. [Z-1997] and [Z-2020]), there exists a zeroth order pseudodifferential operator $P^{\tau}_{t}$ on $X^{\tau}$ (depending smoothly on $t$), such that
    $U^{\tau}_{\sqrt{\rho}}(t)\sim\Pi^{\tau}\circ P^{\tau}_{t}\circ\Pi^{\tau}_{-t},$ (56)
    where $\sim$ stands for ‘equal up to smoothing operators’. More precisely, there exists a classical polyhomogeneous symbol of the form
    $\sigma^{\tau}_{t}(x,r)\sim\sum_{j=0}^{+\infty}\sigma^{\tau}_{t,j}(x)\,r^{-j},$ (57)
    such that
    $P^{\tau}_{t}\sim\sigma^{\tau}_{t}\left(x,D^{\tau}_{\sqrt{\rho}}\right).$
    The leading order term in (57) (equivalently, the principal symbol of $P^{\tau}_{t}$) can be described, up to a unitary factor, as follows.
  5. 5.
    Let us set $J^{\tau}_{t}:=\mathrm{d}\Gamma^{\tau}_{t}\circ J\circ\mathrm{d}\Gamma^{\tau}_{-t}$. Then $J^{\tau}_{t}$ is a new CR structure on $X^{\tau}$, with corresponding Hardy space
    $H(X^{\tau})_{t}:={\Gamma^{\tau}_{-t}}^{*}\big{(}H(X^{\tau})\big{)},$
    and corresponding Szegő kernel
    $\tilde{\Pi}^{\tau}_{t}:={\Gamma^{\tau}_{-t}}^{*}\circ\Pi^{\tau}\circ{\Gamma^{\tau}_{t}}^{*}.$
    Thus, $\tilde{\Pi}^{\tau}_{0}=\Pi^{\tau}$; furthermore, the distributional kernel of $\tilde{\Pi}^{\tau}_{t}$ is
    $\tilde{\Pi}^{\tau}_{t}(x,y):=\left(\Gamma^{\tau}_{-t}\times\Gamma^{\tau}_{-t}\right)^{*}\left(\Pi^{\tau}\right)(x,y)=\Pi^{\tau}\left(\Gamma^{\tau}_{-t}(x),\Gamma^{\tau}_{-t}(y)\right).$
  6. 6.
    For every $x\in X^{\tau}$, the vacuum states at $x$ associated to $J$ and $J_{t}$, denoted $\sigma_{J}^{(x)}$ and $\sigma_{J_{t}}^{(x)}$, are Gaussian functions on the horizontal tangent space at $x$. Their $L^{2}$-pairing $\langle\sigma_{J}^{(x)},\sigma_{J_{t}}^{(x)}\rangle$ varies smoothly with $x$ and $t$ and is nowhere vanishing. Furthermore, there is a smooth function $\theta^{\tau}_{t}(x)$ such that
    $\sigma^{\tau}_{t,0}(x)=e^{\imath\,\theta^{\tau}_{t}(x)}\cdot\langle\sigma_{J}^{(x)},\sigma_{J_{t}}^{(x)}\rangle^{-1}.$ (58)
  7. 7.
    Given the usual description of $\Pi^{\tau}$ as an FIO with complex phase of positive type recalled in section 2.2, we have
    $\Pi^{\tau}_{-t}(x,y)\sim\int_{0}^{+\infty}e^{\imath\,v\,\psi^{\tau}\left(\Gamma^{\tau}_{-t}(x),y\right)}\,s^{\tau}\left(\Gamma^{\tau}_{-t}(x),y,v\right)\,\mathrm{d}v.$
    It follows that the Schwartz kernel of $\mathcal{P}^{\tau}_{t}:=P^{\tau}_{t}\circ\Pi^{\tau}_{-t}$ in (56) is given by
    $\mathcal{P}^{\tau}_{t}(x,y)\sim\int_{0}^{+\infty}e^{\imath\,v\,\psi^{\tau}\left(\Gamma^{\tau}_{-t}(x),y\right)}\,r^{\tau}_{t}\left(x,y,v\right)\,\mathrm{d}v$ (59)
    where
    $r^{\tau}_{t}\left(x,y,v\right)\sim\sum_{j\geq 0}v^{d-1-j}\,r_{t,j}^{\tau}(x,y),$ (60)
    and
    $r_{t,0}^{\tau}(x,y)=\sigma_{t,0}^{\tau}(x)\cdot s^{\tau}_{0}\left(\Gamma^{\tau}_{-t}(x),y\right).$ (61)
3.2 The proof
We can now attack the proof of Theorem 7. Some of the arguments are an equivariant version of others in [P-2024], but we try to make the exposition reasonably self-contained. We shall divide the statement of the Theorem in two parts, and prove each part separately:
  1. 1.
    $\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1},x_{2})=O\left(\lambda^{-\infty}\right)$ uniformly for $\mathrm{dist}_{X^{\tau}}(x_{1},x_{2}^{G\times\mathrm{supp}(\chi)})\geq C\,\lambda^{\epsilon^{\prime}-\frac{1}{2}}$.
  2. 2.
    $\Pi_{\chi,\nu,\lambda}^{\tau}(x_{1},x_{2})=O\left(\lambda^{-\infty}\right)$ uniformly for
    $\max\left\{\mathrm{dist}_{X^{\tau}}\left(x_{1},Z^{\tau}\right),\mathrm{dist}_{X^{\tau}}\left(x_{2},Z^{\tau}\right)\right\}\geq C\,\lambda^{\epsilon^{\prime}-\frac{1}{2}}.$
Proof of Theorem 7, Part 1..
By (18), we have for $x_{1},\,x_{2}\in X^{\tau}$:
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})$
$=$ $\dim(\nu)\,\int_{G}\,\Xi_{\nu}\left(g^{-1}\right)\,\Pi_{\chi,\lambda}\left(\mu^{\tau}_{g^{-1}}(x_{1}),x_{2}\right)\,\mathrm{d}V_{G}(g).$
In view of (55), we may reformulate (3.2) as
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})$
$=$ $\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{G}\,\mathrm{d}V_{G}(g)\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\left[\Xi_{\nu}\left(g^{-1}\right)\,e^{-\imath\,\lambda\,t}\,\chi(t)\,U^{\tau}_{\sqrt{\rho}}\left(t;\mu^{\tau}_{g^{-1}}(x_{1}),x_{2}\right)\right].$
Therefore, given (56), for $\lambda\rightarrow+\infty$ we have
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})$
$\sim$ $\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{G}\,\mathrm{d}V_{G}(g)\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\left[\Xi_{\nu}\left(g^{-1}\right)\,e^{-\imath\,\lambda\,t}\,\chi(t)\,\left(\Pi^{\tau}\circ P^{\tau}_{t}\circ\Pi^{\tau}_{-t}\right)\left(\mu^{\tau}_{g^{-1}}(x_{1}),x_{2}\right)\right];$
here, $\sim$ means ‘has the same asymptotics as’.
The singular support of $\left(\Pi^{\tau}\circ P^{\tau}_{t}\circ\Pi^{\tau}_{-t}\right)$ is the set of pairs $(x,y)\in X^{\tau}\times X^{\tau}$ such that $x=\Gamma^{\tau}_{t}(y)$. Hence, $\left(\Pi^{\tau}\circ P^{\tau}_{t}\circ\Pi^{\tau}_{-t}\right)$ is smooth at $\left(\mu^{\tau}_{g^{-1}}(x_{1}),x_{2}\right)$ unless $x_{1}=\mu^{\tau}_{g}\circ\Gamma^{\tau}_{t}(x_{2})$.
Suppose then that $x_{1}\not\in x_{2}^{G\times\mathrm{supp}(\chi)}$. Then the function
$t\mapsto\chi(t)\cdot\int_{G}\,\mathrm{d}V_{G}(g)\,\left[\Xi_{\nu}\left(g^{-1}\right)\,\left(\Pi^{\tau}\circ P^{\tau}_{t}\circ\Pi^{\tau}_{-t}\right)\left(\mu^{\tau}_{g^{-1}}(x_{1}),x_{2}\right)\right]$
is smooth and compactly supported. Hence its Fourier transform is of rapid decrease.
We conclude the following. Let us define
$\mathcal{K}^{G\times\mathrm{supp}(\chi)}:=\left\{(x_{1},x_{2})\in X^{\tau}\times X^{\tau}\,:\,x_{1}\in x_{2}^{G\times\mathrm{supp}(\chi)}\right\},$
a compact subset of $X^{\tau}\times X^{\tau}$.
Lemma 71.
Let $X^{\prime}\Subset X^{\tau}\times X^{\tau}$ be any open neighbourhood of $\mathcal{K}^{G\times\mathrm{supp}(\chi)}$. Then
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})=O\left(\lambda^{-\infty}\right)\quad\text{as}\quad\lambda\rightarrow+\infty$
uniformly for $(x_{1},x_{2})\not\in X^{\prime}$.
We may thus assume that $x_{1}$ belongs to an arbitrarily small neighbourhood of $x_{2}^{G\times\mathrm{supp}(\chi)}$. We may rewrite (3.2) as follows:
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})$
$\sim$ $\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{G}\,\mathrm{d}V_{G}(g)\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{X^{\tau}}\,\mathrm{d}V_{X^{\tau}}(y)$
$\left[\Xi_{\nu}\left(g^{-1}\right)\,e^{-\imath\,\lambda\,t}\,\chi(t)\,\Pi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)\mathcal{P}^{\tau}_{t}\left(y,x_{2}\right)\right],$
where $\mathcal{P}^{\tau}_{t}$ is as in (59).
The singular support of $\Pi^{\tau}$ is the diagonal in $X^{\tau}\times X^{\tau}$ [BdM-S]. Hence, only a negligible contribution to the asymptotics of (3.2) is lost, if integration in $y$ is restricted to a small neighbourhood of $x_{1}^{G}$. More precisely, let $\varrho_{1}(g,\cdot)$ be a cut-off function, smoothly varying with $g$, identically equal to $1$ sufficiently close to $\mu^{\tau}_{g^{-1}}(x_{1})$, but vanishing identically outside a small open neighbourhood of the point. Then only a rapidly decreasing contribution to the asymptotics is lost, if the integrand in (3.2) is multiplied by $\varrho_{1}(g,y)$. Similarly, the singular support of $\mathcal{P}^{\tau}_{t}$ in (59) is the set of pairs $(x^{\prime},x^{\prime\prime})$ with $x^{\prime}=\Gamma^{\tau}_{t}(x^{\prime\prime})$. Again, we conclude that the asymptotics of (3.2) will be unaltered, if the integrand is further multiplied by $\varrho_{2}(t,y)$, where $\varrho_{2}(t,\cdot)$ varies smoothly with $t$, is supported on a small neighbourhood of $\Gamma^{\tau}_{t}(x_{2})$, and identically equal to $1$ sufficiently close to it.
Given this, the pairs $\big{(}\mu^{\tau}_{g^{-1}}(x_{1}),y\big{)}$ and $\big{(}\Gamma^{\tau}_{-t}(y),x_{2}\big{)}$ belong to small neighbourhoods of the diagonal. Therefore, on the domain of integration we may replace $\Pi^{\tau}$ and $\mathcal{P}^{\tau}_{t}$ by the representations as FIO’s with complex phase [BdM-S], perhaps at the cost of losing a negligible contribution to the asymptotics.
Thus, as $\lambda\rightarrow+\infty$,
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})$
$\sim$ $\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{G}\,\mathrm{d}V_{G}(g)\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{X^{\tau}}\,\mathrm{d}V_{X^{\tau}}(y)\,\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{0}^{+\infty}\,\mathrm{d}v$
$\left[\Xi_{\nu}\left(g^{-1}\right)\,e^{-\imath\,\lambda\,t}\,\chi(t)\,e^{\imath\,u\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)+\imath\,v\,\psi^{\tau}\left(\Gamma^{\tau}_{-t}(y),x_{2}\right)}\right$.
$\left.\varrho_{1}(g,y)\,\varrho_{2}(t,y)\,s^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y,u\right)\,r^{\tau}_{t}\left(y,x_{2},v\right)\right],$
where $r^{\tau}_{t}$ is as in (60). Let us now operate the rescaling $u\mapsto\lambda\,u,\,v\mapsto\lambda\,v$:
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})$
$\sim$ $\lambda^{2}\,\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{G}\,\mathrm{d}V_{G}(g)\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{X^{\tau}}\,\mathrm{d}V_{X^{\tau}}(y)\,\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{0}^{+\infty}\,\mathrm{d}v$
$\left[\Xi_{\nu}\left(g^{-1}\right)\,\chi(t)\,e^{\imath\,\lambda\,\left[u\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)+v\,\psi^{\tau}\left(\Gamma^{\tau}_{-t}(y),x_{2}\right)-t\right]}\right$.
$\left.\varrho_{1}(g,y)\,\varrho_{2}(t,y)\,s^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y,\lambda\,u\right)\,r^{\tau}_{t}\left(y,x_{2},\lambda\,v\right)\right].$
Let us set
$\Psi(x_{1},x_{2};g,t,y,u,v):=u\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)+v\,\psi^{\tau}\left(\Gamma^{\tau}_{-t}(y),x_{2}\right)-t.$ (68)
The next step will be to argue that integration in $u$ and $v$ may be restricted to certain compact neighbourhoods of $1/\tau$ in $\mathbb{R}_{+}$.
Proposition 72.
There exist cut-off functions $f_{1},\,f_{2}\in\mathcal{C}^{\infty}_{c}(\mathbb{R})$, identically equal to $1$ near $1/\tau$, such that the asymptotics of (3.2) are unaltered, if the integral is multiplied by $f_{1}(v)\cdot f_{2}(u)$.
For notational simplicity, the cut-off $f_{1}(v)\cdot f_{2}(u)$ will be implicitly absorbed in the amplitude.
Proof.
As remarked, by the previous reductions, on the domain of integration $\big{(}\mu^{\tau}_{g^{-1}}(x_{1}),y\big{)}$ and $\left(\Gamma^{\tau}_{-t}(y,)x_{2}\right)$ now belong to a small neighbourhood of the diagonal of, say, tubular radius $\delta>0$. Hence, in local coordinates we have
$\mathrm{d}_{\big{(}\mu^{\tau}_{g^{-1}}(x_{1}),\,y\big{)}}\psi^{\tau}$ $=$ $\left(\alpha^{\tau}_{\mu^{\tau}_{g^{-1}}(x_{1})},-\alpha^{\tau}_{\mu^{\tau}_{g^{-1}}(x_{1})}\right)+O(\delta),$ (69)
$\mathrm{d}_{\left(\Gamma^{\tau}_{-t}(y),\,x_{2}\right)}\psi^{\tau}$ $=$ $\left(\alpha^{\tau}_{x_{2}},-\alpha^{\tau}_{x_{2}}\right)+O(\delta).$
[…]
$g:=\tilde{g}\,e^{\xi},\quad y=\tilde{x}+(\theta,\mathbf{v}),\quad t=\tilde{t}+a.$
In local coordinates we get:
$\Gamma^{\tau}_{-t}(y)=\Gamma^{\tau}_{-a}\circ\Gamma_{-\tilde{t}}(y)=\Gamma^{\tau}_{-a}(x_{2})+O(\delta).$
Therefore, we conclude from Lemma 60 that
$\partial_{t}\Psi=v\,\tau-1+O(\delta).$
Since the variable $t$ is compactly supported, it is legitimate to integrate by parts in $t$, and we conclude that the contribution to the asymptotics of (3.2) of the locus where $0<v\ll 1/\tau$ or $v\gg 1/\tau$ is rapidly decreasing.
More precisely, we conclude the following.
Lemma 73.
The asymptotics of (3.2) are unchanged, if the integrand is multiplied by $f_{1}(v)$, where $f_{1}\in\mathcal{C}^{\infty}_{c}(\mathbb{R}_{+})$ is identically equal to $1$ on a suitable neighbourhood of $1/\tau$.
In the following, to simplify notation the cut-off $f_{1}(v)$ will be absorbed in the amplitude of (3.2).
Let us adopts NHLC’s centered at $\tilde{x}:=\mu^{\tau}_{\tilde{g}^{-1}}(x_{1})$, and write
$y=\tilde{x}+(\theta,\mathbf{v}),$
where $\lVert(\theta,\mathbf{v})\rVert=O(\delta)$. Then
$\imath\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)=\imath\,\psi^{\tau}\left(\mu^{\tau}_{e^{-\xi}}\left(\tilde{x}\right),\tilde{x}+(\theta,\mathbf{v})\right)$
$=$ $\imath\,\psi^{\tau}\left(\tilde{x}+\Big{(}\langle\Phi(\tilde{x}),\xi\rangle+R_{3}(\xi),-\xi^{\sharp}_{X^{\tau}}(\tilde{x})+\mathbf{R}_{2}(\xi)\Big{)},\tilde{x}+(\theta,\mathbf{v})\right)$
$=$ $\imath\,\big{(}\langle\Phi(\tilde{x}),\xi\rangle-\theta\big{)}-\frac{1}{4\,\tau^{2}}\,\big{(}\langle\Phi(\tilde{x}),\xi\rangle-\theta\big{)}^{2}-\frac{1}{2}\,\left\|\xi^{\sharp}_{X^{\tau}}(\tilde{x})\right\|^{2}+R_{3}(\xi,\theta).$
[…]
$\Gamma^{\tau}_{-a-\tilde{t}}(y)$ $=$ $\Gamma^{\tau}_{-a}\circ\Gamma^{\tau}_{-\tilde{t}}\Big{(}\tilde{x}+(\theta,\mathbf{v}\big{)}\Big{)}$
$=$ $\Gamma^{\tau}_{-a}\left(\Gamma^{\tau}_{-\tilde{t}}(\tilde{x})+\Big{(}\theta+R_{3}(\theta),M_{\tilde{t}}\,\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v})\Big{)}\right)$
$=$ $\Gamma^{\tau}_{-\tilde{t}}(\tilde{x})+\Big{(}\theta+\tau\,a+R_{3}(\theta,a),M_{\tilde{t}}\,\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v},a)\Big{)}$
By assumption $\Gamma^{\tau}_{-\tilde{t}}(\tilde{x})=x_{2}+\big{(}R_{1}(\delta),\mathbf{R}_{1}(\delta)\big{)}$. Hence by Corollary 62
$\Gamma^{\tau}_{-a-\tilde{t}}(y)=\Gamma^{\tau}_{-\tilde{t}}(\tilde{x})+\Big{(}\theta+\tau\,a+R_{3}(\theta,a),M_{\tilde{t}}\,\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v},a)\Big{)}$
$=$ $\Big{(}x_{2}+\big{(}R_{1}(\delta),\mathbf{R}_{1}(\delta)\big{)}\Big{)}+\Big{(}\theta+\tau\,a+R_{3}(\theta,a),M_{\tilde{t}}\,\mathbf{v}+\mathbf{R}_{2}(\theta,\mathbf{v},a)\Big{)}$
$=$ $x_{2}+\Big{(}\theta+\tau\,a+R_{3}(\theta,a)+R_{1}(\delta),M_{\tilde{t}}\,\mathbf{v}+\mathbf{R}_{1}(\delta)+\mathbf{R}_{2}(\theta,\mathbf{v},a)\Big{)}.$
Therefore,
$\imath\,\psi^{\tau}\left(\Gamma_{-a-\tilde{t}}(y),x_{2}\right)$ (73)
$=$ $\imath\,(\theta+\tau\,a)-\frac{1}{4\,\tau^{2}}\,(\theta+\tau\,a)^{2}-\frac{1}{2}\,\|M_{\tilde{t}}\,\mathbf{v}\|^{2}+R_{2}(\delta,a).$ (74)
Given that $|a|,\,|\theta|\leq C\,(\epsilon+\delta)$, by (68) we conclude that
$\partial_{\theta}\Psi$ $=$ $-u+v+O(\epsilon+\delta).$ (75)
The variable $\theta$ is also compactly supported, and $v$ ranges in a certain neighbourhood of $1/\tau$; therefore, upon choosing $\epsilon$ and $\delta$ suitably small, we conclude that the contribution to the asymptotics of the locus where $0<u\ll 1/\tau$ or $u\gg 1/\tau$ is also negligible. In other words, the asymptotics of (3.2) are unaltered, if the integrand is multiplied by a cut-off function of the form $f_{2}(u)$, where $f_{2}\in\mathcal{C}^{\infty}_{c}(\mathbb{R}_{+})$ is identically equal to $1$ on an appropriate neighbourhood of $1/\tau$.
The proof of Proposition 72 is complete.
By Corollary 1.3 in [BdM-S], there exists a constant $C^{\tau}>0$ such that for any $x^{\prime},x^{\prime\prime}\in X^{\tau}$
$\left|\psi^{\tau}(x^{\prime},x^{\prime\prime})\right|\geq\Im\left(\psi^{\tau}(x^{\prime},x^{\prime\prime})\right)\geq C^{\tau}\,\mathrm{dist}_{X^{\tau}}(x^{\prime},x^{\prime\prime})^{2}.$
Let $\rho_{\lambda}^{\prime}=\rho_{\lambda}^{\prime}(x_{1};g,y)$ be a smooth function identically equal to $1$ where $\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)<a^{\tau}\,\lambda^{\epsilon^{\prime}-1/2}$ and vanishes where $\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)>A^{\tau}\,\lambda^{\epsilon^{\prime}-1/2}$, for certain constants $A^{\tau}>a^{\tau}>0$. Similarly, let $\rho_{\lambda}^{\prime\prime}=\rho_{\lambda}^{\prime\prime}(x_{2};t,y)$ be a smooth function identically equal to $1$ where $\mathrm{dist}_{X^{\tau}}\left(\Gamma^{\tau}_{-t}(y,)x_{2}\right)<b^{\tau}\,\lambda^{\epsilon^{\prime}-1/2}$ and vanishes where $\mathrm{dist}_{X^{\tau}}\left(\Gamma^{\tau}_{-t}(y,)x_{2}\right)>B^{\tau}\,\lambda^{\epsilon^{\prime}-1/2}$, for certain constants $B^{\tau}>b^{\tau}>0$.
Lemma 74.
The asymptotics of (3.2) are unaltered, if the integrand is further multiplied by $\rho_{\lambda}^{\prime}(x_{1};g,y)\,\rho_{\lambda}^{\prime\prime}(x_{2};t,y)$.
Proof.
On the support of $1-\rho_{\lambda}^{\prime}$, we can write
$e^{\imath\,\lambda\,u\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)}=-\frac{\imath}{\lambda}\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)^{-1}\,\partial_{u}\left(e^{\imath\,\lambda\,u\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)}\right).$
Furthermore, for some constant $D^{\tau}>0$
$\left|\frac{\imath}{\lambda}\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)^{-1}\right|\leq D^{\tau}\lambda^{-2\,\epsilon^{\prime}}.$
Hence, by iteratively integrating by parts in the compactly supported variable $u$ we conclude that the contribution to the asymptotics of the locus where $\rho_{\lambda}^{\prime}\neq 1$ is $O\left(\lambda^{-\infty}\right)$.
The argument for $\rho^{\prime\prime}_{\lambda}$ is similar. ∎
Let us choose $r>1$ such that
$\mathrm{dist}_{X^{\tau}}(x^{\prime},x^{\prime\prime})\leq r\,\mathrm{dist}_{X^{\tau}}\left(\Gamma^{\tau}_{t}(x^{\prime}),\Gamma^{\tau}_{t}(x^{\prime\prime})\right)\quad\forall\,x^{\prime},\,x^{\prime\prime}\in X^{\tau},\quad-t\in\mathrm{supp}(\chi).$
On the support of the product $\chi(t)\cdot\rho_{\lambda}^{\prime}(x_{1};g,y)\,\rho_{\lambda}^{\prime\prime}(x_{2};t,y)$, we have
$\mathrm{dist}_{X^{\tau}}\left(x_{1},\Gamma^{\tau}_{t}\circ\mu^{\tau}_{g}(x_{2})\right)$ $=$ $\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1}),\Gamma^{\tau}_{t}(x_{2})\right)$
$\leq$ $\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)+\mathrm{dist}_{X^{\tau}}\left(y,\Gamma^{\tau}_{t}(x_{2})\right)$
$\leq$ $r\,\left[\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)+\mathrm{dist}_{X^{\tau}}\left(\Gamma^{\tau}_{-t}(y),x_{2}\right)\right]$
$=$ $\lambda^{\epsilon^{\prime}-1/2}\cdot r\,\left(A^{\tau}+B^{\tau}\right).$
Given any $C>0$ and $\epsilon^{\prime}>0$, on the other hand, for $0<\epsilon^{\prime\prime}<\epsilon^{\prime}$ and $\lambda\gg 0$ one has
$\lambda^{\epsilon^{\prime\prime}-1/2}\cdot r\,\left(A^{\tau}+B^{\tau}\right)<C\,\lambda^{\epsilon^{\prime}-1/2}.$
This completes the proof of Part 1. of Theorem 7.
Proof of Theorem 7, Part 2..
By the arguments in the proof of Part 1., integration in (3.2) may be restricted to the locus where $\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)\leq A\,\lambda^{\epsilon^{\prime}-1/2}$. Hence we may refine (69) to
$\mathrm{d}_{\big{(}\mu^{\tau}_{g^{-1}}(x_{1}),y\big{)}}\psi^{\tau}$ $=$ $\left(\alpha^{\tau}_{\mu^{\tau}_{g^{-1}}(x_{1})},-\alpha^{\tau}_{\mu^{\tau}_{g^{-1}}(x_{1})}\right)+O\left(\lambda^{\epsilon^{\prime}-\frac{1}{2}}\right).$ (77)
Working locally near a given $g_{0}\in G$, we can write $g=g_{0}\,e^{\xi}$, where $\xi\in\mathfrak{g}$. In view of Corollary 48 and (68), we conclude that on the domain of integration
$\partial_{\xi}\Psi=u\,\langle\Phi(x_{1}),\xi\rangle+O\left(\lambda^{\epsilon^{\prime}-\frac{1}{2}}\right)$
(recall Notation 49). In other words,
$\partial_{g}\Psi=u\,\Phi(x_{1})+O\left(\lambda^{\epsilon^{\prime}-\frac{1}{2}}\right).$
Since $u$ is now bounded away from zero, we conclude by iteratively integrating by parts in $g$ that the contribution to the asymptotics of the locus where $\|\Phi(x_{1})\|\geq C\,\lambda^{\epsilon^{\prime}-\frac{1}{2}}$ is negligible. Given that by assumption 0 is a regular value of $\Phi$, we conclude that
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})=O\left(\lambda^{-\infty}\right),$
uniformly for $\mathrm{dist}_{X^{\tau}}(x_{1},Z^{\tau})\geq C\,\lambda^{\epsilon^{\prime}-1/2}$. The similar statement for $x_{2}$ in the following way: set $\chi_{-}(t):=\chi(-t)$, so that $\overline{\hat{\chi}}=\widehat{\chi_{-}}$, then by definition we have:
$\overline{\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})}$ $=$ $\sum_{j}\overline{\hat{\chi}(\lambda-\lambda_{j})}\,\sum_{k}\rho^{(\nu)}_{j,k}(x_{2})\cdot\overline{\rho^{(\nu)}_{j,k}(x_{1})}$ (78)
$=$ $\sum_{j}\widehat{\chi_{-}}(\lambda-\lambda_{j})\,\sum_{k}\rho^{(\nu)}_{j,k}(x_{2})\cdot\overline{\rho^{(\nu)}_{j,k}(x_{1})}$
$=$ $\Pi^{\tau}_{\chi_{-},\nu,\lambda}(x_{2},x_{1}).$
Applying a similar argument to $\Pi^{\tau}_{\chi_{-},\nu,\lambda}(x_{2},x_{1})$ yields the same conclusion for $x_{2}$. ∎
4 Scaling asymptotics for $\Pi^{\tau}_{\chi,\nu,\lambda}$
In this section, we shall prove Theorems 15 (for $\Pi^{\tau}_{\chi,\nu,\lambda}$), 25, and 28; the adaptations required for $P_{\chi,\nu,\lambda}$ will be dealt with in §5. Before discussing the proofs, let us interject some recalls and remarks.
If $x_{1}\in Z^{\tau}$ and $x_{1}\in x_{2}^{G\times\mathrm{supp}(\chi)}$, then $x_{2}\in Z^{\tau}$ and the action of $G\times\mathbb{R}$ is locally free at both $x_{1}$ and $x_{2}$; furthermore, $x_{1}^{G}\cap x_{2}^{\mathrm{supp}(\chi)}=\{x_{12}\}$ for a unique point $x_{12}$ (Corollary 58), and $\Sigma_{\chi}(x_{1},x_{2})$ is as in (39). Let us fix NHLC’s at $x_{1}$ and $x_{2}$. Given $h_{1}$ as in (39), we obtain from the NHLC’s at $x_{1}$ a system of NHLC’s at $x_{12}=\mu^{\tau}_{h_{1}^{-1}}(x_{1})$, simply by composing with $\mu^{\tau}_{h_{1}}$.
We shall first set the common stage for the proofs, and then specialize the argument for each Theorem separately.
Remark 75.
Having chosen (normal) Heisenberg local coordinates at $x_{1}$ and $x_{2}$, in the following computations we may replace $\omega_{x_{j}}$ (restricted to $\mathcal{H}^{\tau}_{x_{j}}$) by the standard symplectic form $\omega_{0}$ on $\mathbb{R}^{2d-2}$ (see (63) of [P-2024]).
4.1 General arguments for the scaling asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}$
With the aim to study the asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$, with $x_{j,\lambda}$ as in (21), we start from (3.2), with $x_{j,\lambda}$ in place of $x_{j}$. The following arguments will eventually depend on an application of the Stationary Phase Lemma, and all the previously introduced cut-offs are identically equal to $1$ near the stationary point. With abuse of notation, we shall occasionally implicitly absorb the cut-offs in the amplitude, unless it is relevant for the argument to do otherwise. Thus we may write
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$
$\sim$ $\lambda^{2}\,\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{G}\,\mathrm{d}V_{G}(g)\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{X^{\tau}}\,\mathrm{d}V_{X^{\tau}}(y)\,\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{0}^{+\infty}\,\mathrm{d}v$
$\left[e^{\imath\,\lambda\,\Psi(x_{1,\lambda},x_{2,\lambda};g,t,y,u,v)}\,\Xi_{\nu}\left(g^{-1}\right)\,\chi(t)\,s^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1,\lambda}),y,\lambda\,u\right)\,r^{\tau}_{t}\left(y,x_{2,\lambda},\lambda\,v\right)\right].$
where
$\Psi(x_{1,\lambda},x_{2,\lambda};g,t,y,u,v)=u\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1,\lambda}),y\right)+v\,\psi^{\tau}\left(\Gamma^{\tau}_{-t}(y),x_{2,\lambda}\right)-t.$ (80)
By the previous reductions, integration has been reduced to a locus where
$\max\left\{\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1,\lambda}),y\right),\mathrm{dist}_{X^{\tau}}\left(\Gamma^{\tau}_{-t}(y),x_{2,\lambda}\right)\right\}=O\left(\lambda^{\epsilon^{\prime}-\frac{1}{2}}\right);$
hence, we also have
$\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1}),\Gamma^{\tau}_{t}(x_{2})\right)=O\left(\lambda^{\epsilon^{\prime}-\frac{1}{2}}\right).$
Since the action of $G\times\mathbb{R}$ on $Z^{\tau}$ is locally free, this entails that (once the previous cut-offs have been taken into account) $(g,t)$ ranges in a neighbourhood of radius $O\left(\lambda^{\epsilon^{\prime}-1/2}\right)$ of the set $\Sigma_{\chi}(x_{1},x_{2})$ in (38) and (39). We can reformulate this slightly more explicitly as follows.
Let $\gamma^{\mathfrak{g}}:\mathfrak{g}\rightarrow\mathbb{R}$ denote a bump function supported in a small neighbourhood of $0\in\mathfrak{g}$, and identically equal to $1$ sufficiently close to 0. Similarly, let $\gamma^{\mathbb{R}}:\mathbb{R}\rightarrow\mathbb{R}$ denote a compactly supported bump function compactly supported on a small neighbourhood of $0\in\mathbb{R}$, and identically equal to $1$ sufficiently close to 0. We conclude that the asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ are unaltered, if the integrand in (4.1) is further multiplied by the expression
$\gamma^{\mathbb{R}}\left(\lambda^{-\epsilon^{\prime}+1/2}\,(t-t_{1})\right)\cdot\sum_{l=1}^{r}\gamma^{\mathfrak{g}}\left(\lambda^{-\epsilon^{\prime}+1/2}\,\log_{G}\left(g\,h_{1}^{-1}\,\kappa_{l}^{-1}\right)\right),$ (81)
where $\log_{G}=\exp_{G}^{-1}$ is defined on some neighbourhood of $e_{G}$. The cut-off (81) plays the same role as the product $\chi(t)\cdot\rho_{\lambda}^{\prime}(x_{1};g,y)\,\rho_{\lambda}^{\prime\prime}(x_{2};t,y)$ preceding (3.2), and may replace it in the integrand.
If we multiply the integrand in (4.1) by the cut-off in (81), we obtain
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})\sim\sum_{l=1}^{r}\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}$ (82)
where $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}$ is given by right hand side of (4.1), with the integrand multiplied by the $l$-th summand in (81).
We are thus reduced to computing the asymptotics of each summand in (82). To this end, let us make the change of variables
$t\mapsto t_{1}+t,\,g=e^{\xi}\,\kappa_{l}\,h_{1},$
where $(t,\xi)$ now ranges in a neighbourhood of the origin in $\mathbb{R}\times\mathfrak{g}$. The $l$-th summand in (81) now takes the form
$\gamma^{\mathbb{R}}\left(\lambda^{-\epsilon^{\prime}+1/2}\,t\right)\cdot\gamma^{\mathfrak{g}}\left(\lambda^{-\epsilon^{\prime}+1/2}\,\xi\right).$ (83)
On the current domain of integration we have
$\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1}),x_{12}\right)$ $=$ $\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{h_{1}^{-1}}\circ\mu^{\tau}_{\kappa_{l}^{-1}}\circ\mu^{\tau}_{e^{-\xi}}(x_{1}),\mu^{\tau}_{h_{1}^{-1}}(x_{1})\right)$
$=$ $O\left(\lambda^{\epsilon^{\prime}-\frac{1}{2}}\right),$
and $\mathrm{dist}_{X^{\tau}}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)=O\left(\lambda^{\epsilon^{\prime}-\frac{1}{2}}\right)$, whence also $\mathrm{dist}_{X^{\tau}}\left(y,x_{12}\right)=O\left(\lambda^{\epsilon^{\prime}-\frac{1}{2}}\right)$. We can then express $y$ in NHLC’s at $x_{12}$ as
$y=x_{12}+\left(\theta,\mathbf{u}\right),$
where $\|(\theta,\mathbf{u})\|\leq C^{\prime}\,\lambda^{\epsilon^{\prime}-\frac{1}{2}}$ for some constant $C^{\prime}>0$. By (67) and (117) of [P-2024],
$\mathrm{d}V_{X^{\tau}}(y)=\mathcal{V}(\theta,\mathbf{u})\,\mathrm{d}\theta\,\mathrm{d}\mathbf{u},\quad\mathcal{V}(0,\mathbf{0})=\frac{2^{d-1}}{\tau}.$ (84)
Let us operate the rescaling
$(\theta,\,\mathbf{u},\,\xi,\,t)\mapsto\left(\frac{\theta}{\sqrt{\lambda}},\frac{\mathbf{u}}{\sqrt{\lambda}},\,\frac{\xi}{\sqrt{\lambda}},\,\frac{t}{\sqrt{\lambda}}\right);$ (85)
and accordingly rewrite the parametrization of $y$ as
$y_{\lambda}(\theta,\mathbf{u}):=x_{12}+\left(\frac{\theta}{\sqrt{\lambda}},\frac{\mathbf{u}}{\sqrt{\lambda}}\right).$ (86)
In the rescaled variables, the cut-off (83) takes the form
$\gamma^{\mathbb{R}}\left(\lambda^{-\epsilon^{\prime}}\,t\right)\cdot\gamma^{\mathfrak{g}}\left(\lambda^{-\epsilon^{\prime}}\,\xi\right),$ (87)
so that integration in $(\xi,t)$ is now over a ball centered at the origin and radius $O\left(\lambda^{\epsilon^{\prime}}\right)$ in $\mathfrak{g}\times\mathbb{R}$. Similarly, since prior to rescaling $(\theta,\mathbf{v})$ range in a shrinking ball centered at the origin of radius $O\left(\lambda^{\epsilon^{\prime}-1/2}\right)$, the rescaled variables will range in an expanding ball of radius $O\left(\lambda^{\epsilon^{\prime}}\right)$ in $\mathbb{R}\times\mathbb{R}^{2d-2}$.
Let $\mathrm{d}\xi$ denote the Lebesgue measure on $\mathfrak{g}$ associated to a bi-invariant metric on $G$ whose Riemannian density is the Haar measure. We obtain
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}$
$=$ $\lambda^{2-d-d_{G}/2}\,\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{-\infty}^{+\infty}\,\mathrm{d}\theta\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{0}^{+\infty}\,\mathrm{d}v$
$\left[e^{\imath\,\lambda\,\Psi_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,\xi,\theta,\mathbf{v},u,v)_{l}}\,\mathcal{A}_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,\xi,\theta,\mathbf{u},u,v)_{l}\right],$
where, recalling (68),
$\Psi_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,\xi,\theta,\mathbf{u},u,v)_{l}$
$:=$ $\Psi\left(x_{1,\lambda},x_{2,\lambda};e^{\xi/\sqrt{\lambda}}\,\kappa_{l}\,h_{1},t_{1}+\frac{t}{\sqrt{\lambda}},y_{\lambda}(\theta,\mathbf{u}),u,v\right),$
$\mathcal{A}_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,\xi,\theta,\mathbf{u},u,v)_{l}$
$:=$ $\Xi_{\nu}\left(h_{1}^{-1}\,\kappa_{l}^{-1}\,e^{-\xi/\sqrt{\lambda}}\right)\,\chi\left(t_{1}+\frac{t}{\sqrt{\lambda}}\right)\cdot\mathcal{V}\left(\frac{\theta}{\sqrt{\lambda}},\frac{\mathbf{u}}{\sqrt{\lambda}}\right)$
$\cdot s^{\tau}\left(\mu^{\tau}_{(e^{\xi/\sqrt{\lambda}}\kappa_{l}h_{1})^{-1}}(x_{1,\lambda}),y_{\lambda}(\theta,\mathbf{u}),\lambda\,u\right)\,r^{\tau}_{t}\big{(}y_{\lambda}(\theta,\mathbf{u}),x_{2,\lambda},\lambda\,v\big{)}$
$\cdot\tilde{\gamma}_{\lambda}(t,\theta,\xi,\mathbf{u})\cdot f_{1}(v)\cdot f_{2}(u);$
here $\mathcal{V}$ is as in (84), and we have collected in $\tilde{\gamma}$ the cut-offs in the rescaled variables (which may be assumed to be all of the form (87)); finally, $f_{1}$ and $f_{2}$ are as in Proposition 72.
The next step will be to expand (4.1) in descending powers of $\lambda$. Recalling (80), we shall first expand the individual summands. In view of (2.6.1),
$\imath\,u\,\psi^{\tau}\left(\mu^{\tau}_{(e^{\xi/\sqrt{\lambda}}\kappa_{l}h_{1})^{-1}}(x_{1,\lambda}),y_{\lambda}(\theta,\mathbf{u})\right)$
$=$ $\frac{\imath}{\sqrt{\lambda}}\,u\,(\theta_{1}-\theta)$
$+\frac{1}{\lambda}\,u\,\left[\imath\,\omega_{x_{1}}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}\big{)}-\frac{1}{4\,\tau^{2}}\,(\theta_{1}-\theta)^{2}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)\right]$
$+u\,R_{3}\left(\frac{\theta_{1}}{\sqrt{\lambda}},\frac{\mathbf{v}_{1}}{\sqrt{\lambda}},\frac{\theta}{\sqrt{\lambda}},\frac{\mathbf{u}}{\sqrt{\lambda}},\frac{t}{\sqrt{\lambda}},\frac{\xi}{\sqrt{\lambda}}\right);$
in the following, we shall replace $\omega_{x_{j}}$ by $\omega_{0}$ (Remark 75), and abridge $k$-th order remainder terms such as the one on the latter line of (4.1) by the short-hand $R_{k}(\bullet/\sqrt{\lambda})$.
Regarding the second summand in (80), in view of (2.6.2) we have
$\imath\,v\,\psi\left(\Gamma^{\tau}_{-t_{1}-t/\sqrt{\lambda}}\left(x_{12}+\left(\frac{\theta}{\sqrt{\lambda}},\frac{\mathbf{u}}{\sqrt{\lambda}}\right)\right),x_{2}+\left(\frac{\theta_{2}}{\sqrt{\lambda}},\frac{\mathbf{v}_{2}}{\sqrt{\lambda}}\right)\right)$
$=$ $\frac{\imath}{\sqrt{\lambda}}\,v\,(\theta+\tau\,t-\theta_{2})$
$+\frac{v}{\lambda}\,\left[-\frac{1}{4\,\tau^{2}}\,(\theta+\tau\,t-\theta_{2})^{2}+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}\big{)}\right]+R_{3}\left(\frac{\bullet}{\sqrt{\lambda}}\right).$
The upshot is an expansion for (4.1) of the form
$\imath\,\lambda\,\Psi_{\lambda}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,\xi,\theta,\mathbf{u},u,v)_{l}$
$=$ $-\imath\,\lambda\,t_{1}+\imath\,\sqrt{\lambda}\,\Psi^{\tau}_{\theta_{1},\theta_{2}}(t,v,\theta,u)+\mathcal{S}(x,\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,v,\theta,u,\mathbf{u},\xi)_{l}$
$+\lambda\,R_{3}\left(\frac{\bullet}{\sqrt{\lambda}}\right),$
where:
$\Psi^{\tau}_{\theta_{1},\theta_{2}}(t,v,\theta,u):=u\,(\theta_{1}-\theta)+v\,(\theta+\tau\,t-\theta_{2})-t,$ (94)
$\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{1},\mathbf{v}_{1},t,v,\theta,u,\mathbf{u},\xi)_{l}$
$:=$ $u\,\left[\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}\big{)}-\frac{1}{4\,\tau^{2}}\,(\theta_{1}-\theta)^{2}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)\right]$
$+v\,\left[-\frac{1}{4\,\tau^{2}}\,(\theta+\tau\,t-\theta_{2})^{2}+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}\big{)}\right].$
Let us set
$\mathcal{B}_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,v,\theta,u,\mathbf{u},\xi)_{l}$
$:=$ $e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};t,v,\theta,u,\mathbf{u},\xi)_{l}}\,e^{\lambda\,R_{3}\left(\frac{\bullet}{\sqrt{\lambda}}\right)}\cdot\mathcal{A}_{\lambda}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,\xi,\theta,\mathbf{u},u,v)_{l}.$
Then (4.1) may be rewritten as
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}$
$=$ $e^{-\imath\,\lambda\,t_{1}}\cdot\lambda^{2-d-d_{G}/2}\,\frac{\dim(\nu)}{\sqrt{2\,\pi}}\cdot\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\big{[}I_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}\big{]},$
where
$I_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}$
$:=$ $\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{-\infty}^{+\infty}\,\mathrm{d}\theta\,\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{0}^{+\infty}\,\mathrm{d}v$
$\left[e^{\imath\,\sqrt{\lambda}\,\Psi^{\tau}_{\theta_{1},\theta_{2}}(t,v,\theta,u)}\cdot\mathcal{B}_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,v,\theta,u,\mathbf{u},\xi)_{l}\right].$
We can now pair Taylor expansion in the rescaled variables in (4.1) and in the factor $e^{\lambda\,R_{3}\left(\bullet/\sqrt{\lambda}\right)}$ with the asymptotic expansions of the classical symbols $s^{\tau}$ and $r^{\tau}_{t}$. The same arguments leading to the asymptotic expansion (142) of [P-2024] in the action-free case yields a similar expansion in the present setting. Before stating it, let us make the following remarks regarding the leading order terms in $s^{\tau}$ and $r^{\tau}_{t}$ in NHLC’s.
First, by Theorem 51, in NHLC’s at $x_{12}$ we have
$s^{\tau}_{0}\left(\mu^{\tau}_{(\kappa_{l}h_{1})^{-1}}(x_{1}),x_{12}\right)=s^{\tau}_{0}\left(x_{12},x_{12}\right)=\frac{\tau}{(2\,\pi)^{d}}.$ (99)
Second, recalling (61),
$r_{t_{1},0}^{\tau}(x_{12},x_{2}\big{)}$ $=$ $\sigma_{t_{1},0}^{\tau}(x_{12})\cdot s^{\tau}_{0}\left(\Gamma^{\tau}_{-t_{1}}(x_{12}),x_{2}\right)$ (100)
$=$ $\sigma_{t_{1},0}^{\tau}(x_{1})\cdot s^{\tau}_{0}\left(x_{2},x_{2}\right)=\sigma_{t_{1},0}^{\tau}(x_{1})\cdot\frac{\tau}{(2\,\pi)^{d}},$
where the front factor is as in (58). We then have the following.
Lemma 76.
As $\lambda\rightarrow+\infty$, there is an asymptotic expansion
$\mathcal{B}_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,v,\theta,u,\mathbf{u},\xi)_{l}$
$\sim$ $e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,v,\theta,u,\mathbf{u},\xi)_{l}}\cdot\chi(t_{1})\cdot\overline{\Xi_{\nu}\left(\kappa_{l}\,h_{1}\right)}\cdot\frac{2^{d-1}}{\tau}\cdot\lambda^{2\,(d-1)}\cdot(u\,v)^{d-1}\cdot\frac{\tau^{2}}{(2\,\pi)^{2d}}$
$\cdot\sigma_{t_{1},0}^{\tau}(x_{1})\cdot\beta\left(\lambda^{-\epsilon^{\prime}}\,(t,\xi,\theta,\mathbf{u})\right)\cdot f_{1}(u)\cdot f_{2}(v)$
$\cdot\left[1+\sum_{k\geq 1}\lambda^{-k/2}\,B_{k}(u,v,\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,v,\theta,u,\mathbf{u},\xi)_{l}\right],$
where $B_{k}(u,v,\cdot)$ is a polynomial in the rescaled variables, of degree $\leq 3\,k$ and parity $k$.
The latter is indeed an asymptotics expansion for $\epsilon^{\prime}\in(0,1/6)$.
Integration in $u$ and $v$ is compactly supported. Furthermore, in view of (94)
$\partial_{u}\Psi^{\tau}_{\theta_{1},\theta_{2}}(t,v,\theta,u)=\theta_{1}-\theta,\quad\partial_{v}\Psi^{\tau}_{\theta_{1},\theta_{2}}(t,v,\theta,u)=\theta+\tau\,t-\theta_{2}.$
Hence the partial differential $\partial_{u,v}\Psi^{\tau}_{\theta_{1},\theta_{2}}$ satisfies
$\left\|\partial_{u,v}\Psi^{\tau}_{\theta_{1},\theta_{2}}(t,v,\theta,u)\right\|$ $\geq$ $C_{\tau}\,\left\|\left(\theta-\theta_{1},t-\frac{\theta_{2}-\theta_{1}}{\tau}\right)\right\|.$
For any given $\delta>0$, by iteratively integrating by parts in $(u,v)$ on the locus where $\left\|\big{(}\theta-\theta_{1},(\theta_{2}-\theta_{1})/\tau\big{)}\right\|\geq\delta$, one introduces at each step a factor $O\left(\lambda^{-1/2}\right)$. On the other hand, the radius of domain of integration in (4.1) grows like $\lambda^{\epsilon^{\prime}}$; furthermore, once divided by $\lambda^{2\,(d-1)}$ the amplitude $\mathcal{B}_{\lambda}$ remains bounded on the latter domain by Lemma 76 and (4.1). One then has the following.
Lemma 77.
Only a rapidly decreasing contribution to the asymptotics of (4.1) - and (4.1) - is lost, if integration in $(\theta,t)$ is restricted to a fixed and arbitrarily small neighbourhood of $\big{(}\theta_{1},(\theta_{2}-\theta_{1})/\tau\big{)}$.
Leaving a corresponding cut-off function in $(\theta,t)$ implicit, we may now study the asymptotics of $I_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}$ in (4.1) using the stationary phase Lemma. The phase (4.1) has already been considered in [P-2024] (where it is denoted $\Upsilon^{\tau}$). By Lemma 64 of the same paper, (94) has a unique stationary point $P_{s}=(t_{s},v_{s},\theta_{s},u_{s})$, given by
$P_{s}=\left(\frac{\theta_{2}-\theta_{1}}{\tau},\frac{1}{\tau},\theta_{1},\frac{1}{\tau}\right);$ (101)
furthermore, the Hessian determinant and signature at the critical point are, respectively, $\tau^{2}$ and 0. Arguing as in §4 of [P-2024] (in particular, as in the derivation of (147) in loc. cit.) we then obtain an asymptotic expansion for $I_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}$ in (4.1) of the following form:
$I_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}$
$\sim$ $\left(\frac{2\,\pi}{\sqrt{\lambda}}\right)^{2}\cdot\frac{1}{\tau}\cdot e^{\imath\,\sqrt{\lambda}\,\frac{\theta_{1}-\theta_{2}}{\tau}}\cdot e^{\frac{1}{\tau}\,\left[\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}\big{)}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}\big{)}\right]}$
$\cdot\sigma_{t_{1},0}^{\tau}(x_{1})\cdot\chi(t_{1})\cdot\overline{\Xi_{\nu}\left(\kappa_{l}\,h_{1}\right)}\cdot\frac{2^{d-1}}{\tau}\cdot\left(\frac{\lambda}{\tau}\right)^{2\,(d-1)}\cdot\frac{\tau^{2}}{(2\,\pi)^{2d}}$
$\cdot\beta^{\prime}\left(\lambda^{-\epsilon^{\prime}}\,(\xi,\mathbf{u})\right)\cdot\left[1+\sum_{k\geq 1}\lambda^{-k/2}\,F_{k}^{\tau}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}\right]$
$=$ $e^{\imath\,\sqrt{\lambda}\,\frac{\theta_{1}-\theta_{2}}{\tau}}\cdot\frac{\lambda^{2d-3}}{(2\,\pi^{2}\,\tau^{2})^{d-1}}\cdot\sigma_{t_{1},0}^{\tau}(x_{1})\cdot\chi(t_{1})\cdot\overline{\Xi_{\nu}\left(\kappa_{l}\,h_{1}\right)}\cdot e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}}$
$\cdot\beta^{\prime}\left(\lambda^{-\epsilon^{\prime}}\,(\xi,\mathbf{u})\right)\cdot\left[1+\sum_{k\geq 1}\lambda^{-k/2}\,F_{k}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}\right]$
where
$\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}$
$:=$ $\frac{1}{\tau}\,\left[\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}\big{)}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}\big{)}\right],$
$\beta^{\prime}$ is a cut-off identically equal to $1$ near the origin, and $F_{k}(\cdot)_{l}$ is a polynomial in the indicated variables, of degree $\leq 3\,k$ and parity $k$ (and implicitly depending on $x_{1},\,x_{2}$).
More precisely, let us set
$L:=\frac{1}{\tau}\,\left(\frac{\partial^{2}}{\partial t\,\partial\,u}+\frac{\partial^{2}}{\partial t\,\partial\,v}\right)-\frac{\partial^{2}}{\partial\theta\,\partial\,u}.$
Then, when applying the Stationary Phase Lemma to (4.1), the $k$-th summand in the asymptotic expansion of the amplitude in Lemma 76 will yield an asymptotic expansion whose $j$-th term (where $j\geq 0$) is a multiple of
$\frac{4\,\pi^{2}}{\tau}\,\sigma_{t_{1},0}^{\tau}(x_{1})\cdot\chi(t_{1})\cdot\overline{\Xi_{\nu}\left(\kappa_{l}\,h_{1}\right)}\cdot\frac{2^{d-1}}{\tau}\cdot\lambda^{2d-3-\frac{k+j}{2}}\cdot\frac{\tau^{2}}{(2\,\pi)^{2d}}$
$\left.L^{j}\left(e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,v,\theta,u,\mathbf{u},\xi)_{l}}\cdot(u\,v)^{d-1}\cdot B_{k}(u,v,\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,v,\theta,u,\mathbf{u},\xi)_{l}\right)\right|_{P_{s}}.$
Using that $\mathcal{S}$ is homogeneous of degree two in the rescaled variables, one can check inductively that the resulting expansion has the stated form.
Again in view of the cut-offs and the exponential, the asymptotic expansion (4.1) may be integrated term by term. Thus
$\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\big{[}I_{\lambda}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}\big{]}$
$\sim$ $e^{\imath\,\sqrt{\lambda}\,\frac{\theta_{1}-\theta_{2}}{\tau}}\cdot\frac{\lambda^{2d-3}}{(2\,\pi^{2}\,\tau^{2})^{d-1}}\cdot\sigma_{t_{1},0}^{\tau}(x_{1})\cdot\chi(t_{1})\cdot\overline{\Xi_{\nu}\left(\kappa_{l}\,h_{1}\right)}$
$\cdot\sum_{k=0}^{+\infty}\lambda^{-k/2}\mathcal{I}_{k}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})_{l},$
where
$\mathcal{I}_{k}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})_{l}$
$:=$ $\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\left[e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}}\cdot F_{k}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}\right];$
we have set $F_{0}=1$.
Let us first consider the leading order terms in (4.1) and in (4.1), bearing in mind the direct sum decompositions of $T_{x}X^{\tau}$ induced by $\alpha^{\tau}$ at any $x\in X^{\tau}$, and by $\mu^{\tau}$ at any $x\in Z^{\tau}$ (see (8) and Definition 18).
Since in NHLC’s we are unitarily identifying $\mathcal{H}^{\tau}_{x_{1}}\cong\mathcal{H}^{\tau}_{x_{12}}\cong\mathbb{R}^{2d-2}$ (recall Notation 20 above), we have
$\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}$ $=$ $\int_{\mathbb{R}^{d_{G}}_{v}}\,\mathrm{d}\mathbf{u}^{v}\,\int_{\mathbb{R}^{d_{G}}_{t}}\,\mathrm{d}\mathbf{u}^{t}\,\int_{\mathbb{R}^{2d-2-2d_{G}}_{h}}\,\mathrm{d}\mathbf{u}^{h}.$
Furthermore, let us make the following remarks on the geodesic flow $\Gamma^{\tau}_{t}:X^{\tau}\rightarrow X^{\tau}$:
  1. 1.
    $\Gamma^{\tau}_{t}$ preserves $\alpha$ and commutes with $\mu^{\tau}$, hence it preserves the vector bundle decompositions (8) on $X^{\tau}$ and in Definition 18 on $Z^{\tau}$;
  2. 2.
    for any $\xi\in\mathfrak{g}$ and $t\in\mathbb{R}$, the induced vector field $\xi_{X^{\tau}}$ is self-correlated under $\Gamma^{\tau}_{t}$.
Let us consider the case $k=0$ in (4.1). We have (Remark 75):
$\mathcal{I}_{0}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})_{l}$
$=$ $\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\left[e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}}\right]$
$=$ $\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\left[e^{\frac{1}{\tau}\,\left[\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}\big{)}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}\big{)}\right]}\right].$
Let us set
$\mathbf{v}_{j}(\tau):=\frac{1}{\sqrt{\tau}}\,\mathbf{v}_{j},$ (107)
and perform the change of coordinates
$\xi\rightarrow\sqrt{\tau}\,\xi,\quad\mathbf{u}\rightarrow\sqrt{\tau}\,\mathbf{u}.$
We obtain
$\mathcal{I}_{0}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})_{l}$
$=$ $\tau^{d-1+d_{G}/2}\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\left[e^{\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}(\tau)\big{)}+\psi_{2}\left(\mathbf{v}_{1}(\tau)^{(l)}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}(\tau)\big{)}}\right].$
We have
$\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}(\tau)\big{)}+\psi_{2}\left(\mathbf{v}_{1}(\tau)^{(l)}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}(\tau)\big{)}$ (109)
$=$ $\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}(\tau)\big{)}-\imath\,\omega_{x}\left(\mathbf{v}_{1}(\tau)^{(l)},\mathbf{u}\right)$
$-\frac{1}{2}\,\left\|\mathbf{v}^{t}_{1}(\tau)^{(l)}-\mathbf{u}^{t}\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{1}(\tau)^{(l)}-\mathbf{u}^{h}\right\|^{2}$
$+\imath\,\omega_{0}\left(\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{v}_{1}^{t}(\tau)^{(l)}+\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|\xi_{X^{\tau}}(x_{1})^{(l)}+\mathbf{u}^{v}\right\|^{2}.$
We first compute the $\xi$-integral. We can transfer the integral over $\mathfrak{g}$ to $T_{x_{1}}^{v}X^{\tau}\cong\mathbb{R}^{d_{G}}_{v}$, the tangent space to $x_{1}^{G}$. If $r=|G_{x_{1}}|$, the action map $G\rightarrow G\cdot x_{1}$ is $r_{x_{1}}:1$ ($r_{x_{1}}=|G_{x_{1}}|$). We then have the replacement
$\int_{\mathfrak{g}}\,\mathrm{d}\xi\quad\text{by}\quad\frac{1}{r_{x_{1}}\cdot V_{eff}(x_{1})}\cdot\int_{T_{x_{1}}^{v}X^{\tau}}\mathrm{d}\mathbf{w}^{v}=\frac{1}{r_{x_{1}}\cdot V_{eff}(x_{1})}\cdot\int_{\mathbb{R}^{d_{G}}_{v}}\mathrm{d}\mathbf{a}.$
Then
$\int_{\mathfrak{g}}e^{\imath\,\omega_{0}\left(\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{v}_{1}^{t}(\tau)^{(l)}+\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|\xi_{X^{\tau}}(x_{1})^{(l)}+\mathbf{u}^{v}\right\|^{2}}\,\mathrm{d}\xi$
$=$ $\frac{1}{r_{x_{1}}\cdot V_{eff}(x_{1})}\cdot\int_{\mathbb{R}^{d_{G}}}e^{\imath\,\omega_{0}\left(\mathbf{a}^{(l)},\mathbf{v}_{1}^{t}(\tau)^{(l)}+\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|\mathbf{a}^{(l)}+\mathbf{u}^{v}\right\|^{2}}\mathrm{d}\mathbf{a}$
$=$ $\frac{(2\,\pi)^{d_{G}/2}}{r_{x_{1}}\cdot V_{eff}(x_{1})}\cdot e^{-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{v}_{1}^{t}(\tau)^{(l)}+\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|\mathbf{v}_{1}^{t}(\tau)^{(l)}+\mathbf{u}^{t}\right\|^{2}}.$
We can insert (4.1) in (4.1) and obtain
$\mathcal{I}_{0}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})_{l}$
$=$ $\frac{(2\,\pi)^{d_{G}/2}}{r\cdot V_{eff}(x_{1})}\cdot\tau^{d-1+d_{G}/2}\,\int_{\mathbb{R}^{2d-2}}\,e^{A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})_{l}}\,\mathrm{d}\mathbf{u},$
where
$A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})_{l}$ $=$ $\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}(\tau)\big{)}-\imath\,\omega_{0}\left(\mathbf{v}_{1}(\tau)^{(l)},\mathbf{u}\right)$
$-\frac{1}{2}\,\left\|\mathbf{v}^{t}_{1}(\tau)^{(l)}-\mathbf{u}^{t}\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{1}(\tau)^{(l)}-\mathbf{u}^{h}\right\|^{2}$
$-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{v}_{1}^{t}(\tau)^{(l)}+\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|\mathbf{v}_{1}^{t}(\tau)^{(l)}+\mathbf{u}^{t}\right\|^{2}$
$=$ $-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}-\left\|\mathbf{u}^{t}\right\|^{2}-\imath\,\omega_{0}\left(\mathbf{v}^{h}_{1}(\tau)^{(l)},\mathbf{u}^{h}\right)-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{1}(\tau)^{(l)}-\mathbf{u}^{h}\right\|^{2}$
$-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{u}^{t}\right)+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}(\tau)\big{)}.$
4.2 The leading order term in the action free case
In this section, we shall determine the leading order term in in Theorem 15. To avoid repetitions, we shall later give a general argument for the lower order terms covering the general equivariant case.
Determination of the leading order term in Theorem 15.
In the action-free case, we have $\mathbf{v}_{j}=\mathbf{v}_{j}^{h}$, and the suffix $l$ may be omitted. Thus
$\mathcal{I}_{0}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})=\tau^{d-1}\,\int_{\mathbb{R}^{2d-2}}\,e^{F(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})}\,\mathrm{d}\mathbf{u},$ (113)
where
$F\big{(}\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u}\big{)}=\psi_{2}\big{(}\mathbf{v}_{1}(\tau),\mathbf{u}\big{)}+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}(\tau)\big{)}.$
As is Definition 8, following [F], let us write
$B^{-1}_{c}=\begin{pmatrix}P&Q\\ \overline{Q}&\overline{P}\end{pmatrix}$, (114)
where $P$ is invertible and $\|P\mathbf{v}\|\geq\|\mathbf{v}\|$ for every $\mathbf{v}\in\mathbb{C}^{d}$. Furthermore, let us choose a metaplectic lift $\tilde{B}^{-1}$ of $B^{-1}$, and denote by $K_{1,\tilde{B}^{-1}}$ the integral kernel of the metaplectic representation of $\tilde{B}^{-1}$. With the present normalizations (here $\omega_{0}=(\imath/2)\,\sum_{j}\mathrm{d}z_{j}\wedge\mathrm{d}\overline{z}_{j}$), it follows from the discussion in §4 of [ZZ-2018] (and Ch. 4 of [F]) that, with the notation of Definition 9, $K_{1,\tilde{B}^{-1}}:\mathbb{C}^{d-1}\times\mathbb{C}^{d-1}\rightarrow\mathbb{C}$ is given by
$K_{1,\tilde{B}^{-1}}(Z,W):=\pi^{-(d-1)}\cdot\det(P)^{-1/2}\cdot\exp\left[\Psi_{B_{c}^{-1}}(Z,W)\right],$ (115)
the square root being well-defined at the metaplectic level. We may alternatively view $K_{1,\tilde{B}^{-1}}$ as defined on $K_{1,\tilde{B}^{-1}}:\mathbb{R}^{2d-2}\times\mathbb{R}^{2d-2}\rightarrow\mathbb{C}$.
On the other hand, let $\Pi_{1}$ be the integral kernel of the level-$1$ Szegő kernel on $(\mathbb{C}^{d-1},\omega_{0})$ with the standard polarization. Explicitly, this is given by
$\Pi_{1}(Z,W):=\pi^{-(d-1)}\,e^{\psi_{2}(Z,W)}.$
where $\psi_{2}$ is as in Definition 11 and Notation 12 (there is some abuse of language here, since the Szegő kernel is really defined on $X\times X$, where $X\cong\mathbb{C}^{d-1}\times S^{1}$ is an appropriate unit circle bundle on $\mathbb{C}^{d-1}$, and $\Pi_{1}$ above is in fact its pull-back under the map $Z\mapsto(Z,1)$). Again, we can equivalently view $\Pi_{1}$ as defined on $\mathbb{R}^{2d-2}\times\mathbb{R}^{2d-2}$.
The relation between $K_{1,\tilde{B}^{-1}}$ and $\Pi_{1}$ is described in Proposition 4.4 of [ZZ-2018] (building on the theory of [D]). Namely, if $Z,\,W\in\mathbb{C}^{d-1}$ correspond to $\mathbf{v},\,\mathbf{w}\in\mathbb{R}^{2d-2}$, then
$K_{1,\tilde{B}^{-1}}(Z,W)$ $=$ $K_{1,\tilde{B}^{-1}}(\mathbf{v},\mathbf{w})$
$=$ $\det(P^{*})^{1/2}\,\int_{\mathbb{R}^{2d-2}}\,\Pi_{1}(\mathbf{v},B^{-1}\mathbf{u})\,\Pi_{1}(\mathbf{u},\mathbf{w})\,\mathrm{d}\mathbf{u}.$
Given (115) and (4.2), we may rewrite (113) as
$\mathcal{I}_{0}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})$ $=$ $\tau^{d-1}\,\int_{\mathbb{R}^{2(d-1)}}\,e^{\psi_{2}(\mathbf{v}_{1}(\tau),\mathbf{u})+\psi_{2}(B\mathbf{u},\mathbf{v}_{2}(\tau))}\,\mathrm{d}\mathbf{u}$ (117)
$=$ $\pi^{2(d-1)}\,\tau^{d-1}\,\int_{\mathbb{C}^{d-1}}\,\Pi_{1}\big{(}\mathbf{v}_{1}(\tau),\mathbf{u}\big{)}\,\Pi_{1}\big{(}B\mathbf{u},\mathbf{v}_{2}(\tau)\big{)}\,\mathrm{d}\mathbf{u}$
$=$ $\pi^{2(d-1)}\,\tau^{d-1}\,\int_{\mathbb{C}^{d-1}}\,\Pi_{1}\big{(}\mathbf{v}_{1}(\tau),B^{-1}\mathbf{u}\big{)}\,\Pi_{1}\big{(}\,\mathbf{u},\mathbf{v}_{2}(\tau)\big{)}\,\mathrm{d}\mathbf{u}$
$=$ $\pi^{2(d-1)}\,\tau^{d-1}\,\det(P^{*})^{-1/2}\,K_{1,\tilde{B}_{c}^{-1}}\big{(}\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau)\big{)}$
$=$ $(\pi\,\tau)^{d-1}\,|\det(P)|^{-1}\,\cdot\exp\left[\Psi_{B^{-1}}\big{(}\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau)\big{)}\right].$
Using (117), we obtain that to leading order (4.1) is
$\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\big{[}I_{\lambda}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}\big{]}$
$\asymp$ $e^{\imath\,\sqrt{\lambda}\,\frac{\theta_{1}-\theta_{2}}{\tau}}\cdot\frac{\lambda^{2d-3}}{(2\,\pi\,\tau)^{d-1}}\cdot\sigma_{t_{1},0}^{\tau}(x_{1})\cdot\chi(t_{1})$
$\cdot|\det(P)|^{-1}\,\cdot\exp\left[\frac{1}{\tau}\,\Psi_{B^{-1}}\big{(}\mathbf{v}_{1},\mathbf{v}_{2}\big{)}\right]$
Lemma 78.
Let the unitary factor $e^{\imath\,\theta^{\tau}_{t}(x)}$ be as in (58). Then
$\sigma_{t_{1},0}^{\tau}(x_{1})\cdot\left|\det(P)\right|^{-1}=e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}.$
Proof of Lemma 78.
Let $B$ be as in (25). In view of (114) and the computations in Ch. 4 of [F] (especially Proposition 4.17) we have
$B_{c}=\begin{pmatrix}P^{*}&-Q^{\dagger}\\ -Q^{*}&P^{\dagger}\end{pmatrix}$.
Let us write $B$ in square block form:
$B=\begin{pmatrix}A^{\prime}&B^{\prime}\\ C^{\prime}&D^{\prime}\end{pmatrix}$.
It follows from the computation in the proof of Lemma 3.3 of [Z-2020] (based on [D]) that, with the notation in (58),
$\langle\sigma_{J}^{(x)},\sigma_{J_{t_{1}}}^{(x_{1})}\rangle$ $=$ $\frac{2^{d-1}}{\left|\det\big{(}A^{\prime}+D^{\prime}+\imath\,(B^{\prime}-C^{\prime})\big{)}\right|}$ (119)
$=$ $\frac{2^{d-1}}{\left|\det\big{(}2\,P\big{)}\right|}=\frac{1}{|\det(P)|},$
where on the last line we have made use of (54) of [Z-2020]. The claim follows. ∎
Thus, in the action-free case, going back to (4.1) we obtain that to leading order
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ (121)
$\asymp$ $e^{-\imath\,\lambda\,t_{1}}\cdot\lambda^{2-d}\,\frac{1}{\sqrt{2\,\pi}}\cdot e^{\imath\,\sqrt{\lambda}\,\frac{\theta_{1}-\theta_{2}}{\tau}}\cdot\frac{\lambda^{2d-3}}{(2\,\pi\,\tau)^{d-1}}\cdot\sigma_{t_{1},0}^{\tau}(x_{1})\cdot\chi(t_{1})$
$\cdot|\det(P)|^{-1}\,\cdot\exp\left[\frac{1}{\tau}\,\Psi_{B^{-1}}\big{(}\mathbf{v}_{1},\mathbf{v}_{2}\big{)}\right]$
$=$ $e^{-\imath\,\lambda\,t_{1}}\cdot\frac{\chi(t_{1})}{\sqrt{2\,\pi}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}$
$\cdot\exp\left\{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})+\Psi_{B^{-1}}\big{(}\mathbf{v}_{1},\mathbf{v}_{2}\big{)}\right]\right\},$
as claimed (with $\theta^{\tau}(x_{1},x_{2})=\theta^{\tau}_{t_{1}}(x_{1})$).
4.3 Proof of Theorem 25
Let us return to the general equivariant setting.
Proof of statement 1 of Theorem 25.
We are assuming $\mathbf{v}_{j}=\mathbf{v}_{j}^{t}$. As in Definition 18, let us decompose $\mathbf{u}=\mathbf{u}^{h}+\mathbf{u}^{t}+\mathbf{u}^{v}$. Then (4.1) reduces to
$A\big{(}\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u}\big{)}_{l}$ $=$ $-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}-\left\|\mathbf{u}^{t}\right\|^{2}$
$-\frac{1}{2}\,\left\|\mathbf{u}^{h}\right\|^{2}-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{u}^{t}\right)+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}^{t}(\tau)\big{)};$
in particular, in this case $A(\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{u})_{l}$ is independent of $l$. Hence (4.1) may be rewritten
$\mathcal{I}_{0}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})_{l}$
$=$ $\frac{(2\,\pi)^{d_{G}/2}}{r\cdot V_{eff}(x_{1})}\cdot\tau^{d-1+d_{G}/2}\,e^{-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}}\cdot A_{x_{1},x_{2}}\left(\mathbf{v}_{2}^{t}(\tau)\right),$
where $A_{x_{1},x_{2}}\left(\mathbf{v}_{2}^{t}(\tau)\right)$ is independent of $\mathbf{v}_{1}(\tau)$ (and $l$).
Let $A_{\chi}:\mathfrak{X}_{\chi}^{\tau}\rightarrow\mathbb{C}$ be as in Definition 22.
Lemma 79.
Let $\chi_{-}(t):=\chi(-t)$. With the previous hypothesis and notation,
$\overline{e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot A_{\chi}(x_{1},x_{2})}$ $=$ $e^{-\imath\,\theta^{\tau}_{-t_{1}}(x_{1})}\cdot A_{\chi_{-}}(x_{2},x_{1}),$
$e^{-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}}\cdot A_{x_{1},x_{2}}(\mathbf{v}_{2}^{t}(\tau))$ $=$ $e^{-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}-\left\|\mathbf{v}_{2}^{t}(\tau)\right\|^{2}}\cdot A_{\chi}(x_{1},x_{2}).$
Proof.
Let $\chi_{-}(t)=\chi(-t)$ then $\overline{\hat{\chi}}=\widehat{\chi_{-}}$ and $\chi_{-}\in\mathcal{C}^{\infty}_{c}\big{(}[-t_{0}-\epsilon,-t_{0}+\epsilon]\big{)}$. Referring to (38),
$\Sigma_{\chi_{-}}(x_{2},x_{1})=\Sigma_{\chi}(x_{1},x_{2})^{-1},$ (124)
where inversion is meant in the group $G\times\mathbb{R}$.
The computation in (78) yields:
$\overline{\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})}=\Pi^{\tau}_{\chi_{-},\nu,\lambda}(x_{2,\lambda},x_{1,\lambda}).$ (125)
By assumption, $G_{x_{2}}=h_{1}^{-1}\,G_{x_{1}}\,h_{1}$. Suppose $G_{x_{1}}=\left\{\kappa_{1},\ldots,\kappa_{r}\right\}$; then
$G_{x_{2}}=\left\{h_{1}^{-1}\,\kappa_{1}\,h_{1},\ldots,h_{1}^{-1}\,\kappa_{r}\,h_{1}\right\}=\left\{\tilde{\kappa}_{1},\ldots,\tilde{\kappa}_{r}\right\},$
where $\tilde{\kappa}_{j}:=h_{1}^{-1}\,\kappa_{j}^{-1}\,h_{1}$. For every $l=1,\ldots,r$,
$\overline{\Xi_{\nu}\left(\tilde{\kappa}_{l}\,h_{1}^{-1}\right)}=\overline{\Xi_{\nu}\left(h_{1}^{-1}\,\kappa_{l}^{-1}\,h_{1}\,h_{1}^{-1}\right)}=\overline{\Xi_{\nu}\left(h_{1}^{-1}\,\kappa_{l}^{-1}\right)}=\Xi_{\nu}\left(\kappa_{l}\,h_{1}\right).$
Let us apply (4.1), (4.1), and (4.1) with $x_{1,\lambda}$ and $x_{2,\lambda}$ swapped and $\chi$ replaced by $\chi_{-}$. Then $t_{1}$ is replaced with $t_{1}$, and $B$ with $B^{-1}$; hence $P$ gets replaced with $P^{*}$ in (114) (see the proof of Proposition 4.17 of [F]). Therefore, $\left|\sigma_{-t_{1},0}^{\tau}(x_{2})\right|=\left|\sigma_{t_{1},0}^{\tau}(x_{1})\right|$. More precisely, by Lemma 78,
$\sigma_{t_{1},0}^{\tau}(x_{1})=e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\left|\det(P)\right|,\quad\sigma_{-t_{1},0}^{\tau}(x_{2})=e^{\imath\,\theta^{\tau}_{-t_{1}}(x_{2})}\cdot\left|\det(P)\right|.$ (126)
Then on the one hand we have that to leading order
$\Pi^{\tau}_{\chi_{-},\nu,\lambda}(x_{2,\lambda},x_{1,\lambda})_{l}$
$\asymp$ $e^{\imath\,\lambda\,t_{1}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\,\frac{\dim(\nu)}{\sqrt{2\,\pi}}\cdot e^{\imath\,\sqrt{\lambda}\,\frac{\theta_{2}-\theta_{1}}{\tau}}\cdot e^{\imath\,\theta^{\tau}_{-t_{1}}(x_{2})}\cdot\left|\det(P)\right|$
$\cdot\chi(t_{1})\cdot\Xi_{\nu}\left(\kappa_{l}\,h_{1}\right)\cdot\frac{1}{r\cdot V_{eff}(x_{1})}\cdot e^{-\left\|\mathbf{v}_{2}^{t}(\tau)\right\|^{2}}\cdot\frac{A_{x_{2},x_{1}}\left(\mathbf{v}_{1}^{t}(\tau)\right)}{\pi^{d-1}}.$
On the other hand, given (125), taking the complex conjugate of the expansion for $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}$ we also have
$\overline{\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}}$
$\asymp$ $e^{\imath\,\lambda\,t_{1}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\,\frac{\dim(\nu)}{\sqrt{2\,\pi}}\cdot e^{\imath\,\sqrt{\lambda}\,\frac{\theta_{2}-\theta_{1}}{\tau}}\cdot e^{-\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\left|\det(P)\right|$
$\cdot\chi(t_{1})\cdot\Xi_{\nu}\left(\kappa_{l}\,h_{1}\right)\cdot\frac{1}{r\cdot V_{eff}(x_{1})}\cdot e^{-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}}\cdot\frac{\overline{A_{x_{1},x_{2}}\left(\mathbf{v}_{2}^{t}(\tau)\right)}}{\pi^{d-1}}.$
Therefore,
$e^{-\left\|\mathbf{v}_{2}^{t}(\tau)\right\|^{2}}\cdot A_{x_{2},x_{1}}(\mathbf{v}_{1}^{t}(\tau))\cdot e^{\imath\,\theta^{\tau}_{-t_{1}}(x_{2})}=e^{-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}}\cdot\overline{e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot A_{x_{1},x_{2}}(\mathbf{v}_{2}^{t}(\tau))}.$
If
$A_{\chi}^{\prime}(x_{1},x_{2}):=e^{\left\|\mathbf{v}^{t}\right\|^{2}}\cdot A_{x_{1},x_{2}}(\mathbf{v}^{t}),$
we conclude that $A_{\chi}(x_{1},x_{2})$ is independent of $\mathbf{v}^{t}$, and this implies the stated equalities. Setting $\mathbf{v}_{j}=\mathbf{0}$, one obtains $A^{\prime}_{\chi}(x_{1},x_{2})=A_{\chi}(x_{1},x_{2})$.
Returning to (4.1), again in view of (126) we obtain that to leading order
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}$
$\asymp$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$
$\cdot e^{\imath\,\sqrt{\lambda}\,\frac{1}{\tau}\,(\theta_{1}-\theta_{2})-\frac{1}{\tau}\left(\|\mathbf{v}_{1}^{t}\|^{2}+\|\mathbf{v}_{2}^{t}\|^{2}\right)},$
where $\mathcal{F}_{\chi}(x_{1},x_{2})$ and $\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$ are as in Definition 23 (here $g_{l}=\kappa_{l}\,h_{1}$).
Let us now consider the lower order terms in the asymptotic expansion of (4.1), hence in (4.1). For each $k$, arguing as in (4.1) and (109) one verifies that $\mathcal{I}_{k}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})_{l}$ in (4.1) is a linear combination of integrals of the form
$\theta_{1}^{a}\,\mathbf{v}_{1}^{A}\,\theta_{2}^{a^{\prime}}\,\mathbf{v}_{2}^{A^{\prime}}$ (130)
$\cdot\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\left[\mathbf{u}^{C}\,e^{\frac{1}{\tau}\,\left[\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}\big{)}-\imath\,\omega_{0}\left(\mathbf{v}_{1}^{(l)},\mathbf{u}\right)-\frac{1}{2}\,\left\|(\mathbf{v}^{t}_{1})^{(l)}-\mathbf{u}^{t}\right\|^{2}-\frac{1}{2}\,\left\|(\mathbf{v}^{h}_{1})^{(l)}-\mathbf{u}^{h}\right\|^{2}\right]}\right$.
$\left.\cdot\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\left[e^{\imath\,\omega_{0}\left(\xi_{X^{\tau}}(x_{1})^{(l)},(\mathbf{v}_{1}^{t})^{(l)}+\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|\xi_{X^{\tau}}(x_{1})^{(l)}+\mathbf{u}^{v}\right\|^{2}}\cdot\xi^{D}\right]\right],$
where $a,\,a^{\prime}\geq 0$, $A,A^{\prime},C,D\geq 0$ are multi-indexes and $a+a^{\prime}+|A|+|A^{\prime}|+|C|+|D|$ has the same parity as $k$ and is $\leq 3\,k$.
Let us consider the inner $\xi$-integral in (130). Arguing as for (4.1), we can transfer integration to $T^{v}X^{\tau}\cong\mathbb{R}^{d_{G}}$, and conclude that, up to some multiplicative constant factor, the $\xi$-integral is given by
$\int_{\mathbb{R}^{d_{G}}}\mathbf{a}^{D}\,e^{\frac{1}{\tau}\left[\imath\,\omega_{0}\left(\mathbf{a}^{(l)},(\mathbf{v}_{1}^{t})^{(l)}+\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|\mathbf{a}^{(l)}+\mathbf{u}^{v}\right\|^{2}\right]}\mathrm{d}\mathbf{a}$
$=$ $\sum_{D^{\prime}+D^{\prime\prime}=D}\beta_{D^{\prime},D^{\prime\prime}}\cdot\mathbf{u}^{D^{\prime\prime}}\,e^{-\imath\,\omega_{0}\left(\mathbf{u}^{v},(\mathbf{v}_{1}^{t})^{(l)}+\mathbf{u}^{t}\right)}\int_{\mathbb{R}^{d_{G}}}\mathbf{b}^{D^{\prime}}\,e^{\frac{1}{\tau}\left[\imath\,\omega_{0}\left(\mathbf{b},(\mathbf{v}_{1}^{t})^{(l)}+\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|\mathbf{b}\right\|^{2}\right]}\mathrm{d}\mathbf{a},$
for certain $\beta_{D^{\prime},D^{\prime\prime}}\in\mathbb{C}$.
On the other hand, for any $\mathbf{r}\in\mathbb{R}^{d_{G}}$ for certain constants we have
$\int_{\mathbb{R}^{d_{G}}}\,\mathbf{b}^{D^{\prime}}\,e^{\imath\,\mathbf{b}\cdot\mathbf{r}-\frac{1}{2}\,\|\mathbf{b}\|^{2}}\,\mathbf{d}\mathbf{b}=\sum_{L\geq 0}c_{L}\mathbf{r}^{L}\,e^{-\frac{1}{2}\,\|\mathbf{r}\|^{2}},$ (132)
where $c_{L}\neq 0$ only if $L\leq D^{\prime}$, and $|L|$ and $|D^{\prime}|$ have the same parity. We conclude that (4.3) can be rewritten in the form
$\sum_{D^{\prime}+L}\tilde{\beta}_{D^{\prime},L}\cdot\mathbf{u}^{D^{\prime\prime}}\,\mathbf{v}_{1}^{L}e^{-\imath\,\omega_{0}\left(\mathbf{u}^{v},(\mathbf{v}_{1}^{t})^{(l)}+\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|(\mathbf{v}_{1}^{t})^{(l)}+\mathbf{u}^{t}\right\|^{2}},$ (133)
where $\tilde{\beta}_{D^{\prime},L}\neq 0$ only if $D^{\prime}+L\leq D$ and $|D^{\prime}|+|L|$ has the same parity as $|D|$.
If we insert (133) in (130), assuming that $\mathbf{v}_{j}=\mathbf{v}_{j}^{t}$, we conclude that the latter is in turn a linear combination of integrals of the form
$\theta_{1}^{a}\,\mathbf{v}_{1}^{A}\,\theta_{2}^{a^{\prime}}\,\mathbf{v}_{2}^{A^{\prime}}\,e^{-\left\|\mathbf{v}^{t}_{1}\right\|^{2}}$
$\cdot\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\left[\mathbf{u}^{C}\,e^{\frac{1}{\tau}\,\left[\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}^{t}\big{)}-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{u}^{t}\right)-\left\|\mathbf{u}^{t}\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{u}^{h}\right\|^{2}\right]}\right],$
where $a+a^{\prime}+|A|+|A^{\prime}|+|C|$ has the same parity as $k$, and is $\leq 3\,k$.
Let us define $\mathcal{L}:\mathbb{R}^{2d-2}\rightarrow\mathbb{C}$ by setting
$\mathcal{L}(\mathbf{v}):=\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\left[\mathbf{u}^{C}\,e^{\frac{1}{\tau}\,\left[\psi_{2}\big{(}B\mathbf{u},\mathbf{v}\big{)}-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{u}^{t}\right)-\left\|\mathbf{u}^{t}\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{u}^{h}\right\|^{2}\right]}\right].$
The exponent in the integrand has the form
$\left\langle\mathbf{u},\frac{1}{\tau}J_{0}B^{-1}J_{0}\mathbf{v}\right\rangle+\imath\,\left\langle\mathbf{u},\frac{1}{\tau}\,J_{0}\,B^{-1}\,\mathbf{v}\right\rangle-\frac{1}{2}\,\langle\mathbf{u},R\,\mathbf{u}\rangle,$
where $R$ is a certain complex symmetric matrix with positive definite real part. It follows that $\mathcal{L}(\mathbf{v})$ is a linear combination of terms of the form
$\mathbf{v}^{D}\,e^{-\frac{1}{2}\,\left\langle\mathbf{v},\tilde{R}^{-1}\,\mathbf{v}\right\rangle},$
where $\tilde{R}$ is another complex symmetric matrix with $\Re(\tilde{R})\gg 0$, $|D|\leq|C|$ and $|D|$, $|C|$ have the same parity. If we finally insert $\mathbf{v}=\mathbf{v}_{2}^{t}$, we obtain the claimed statement.
Remark 80.
The same arguments for the lower order term can be applied to the action-free case (with minor modifications), thus completing the proof of Theorem 15.
Proof of Statement 2 of Theorem 25.
In this case, (4.1) is
$A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})_{l}$
$=$ $-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{1}(\tau)\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}_{2}^{t}(\tau)\right\|^{2}$
$-\imath\,\omega_{0}\left(\mathbf{v}^{h}_{1}(\tau)^{(l)},\mathbf{u}^{h}\right)+g_{0}\left(\mathbf{v}^{h}_{1}(\tau)^{(l)},\mathbf{u}^{h}\right)-\imath\,\omega_{0}\left(B\mathbf{u},\mathbf{v}_{2}^{t}(\tau)\right)+g_{0}(B\mathbf{u},\mathbf{v}_{2}^{t}(\tau)\big{)}$
$-\left\|\mathbf{u}^{t}\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{u}^{h}\right\|^{2}-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|B\mathbf{u}\right\|^{2}.$
$=$ $-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{1}(\tau)\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}_{2}^{t}(\tau)\right\|^{2}-\imath\,\langle\mathbf{V}(\tau),\,A_{l}\,\mathbf{u}\rangle-\frac{1}{2}\,\mathbf{u}^{t}\,R\,\mathbf{u},$
where $\mathbf{V}(\tau)^{\dagger}=\begin{pmatrix}\mathbf{v}_{1}^{t}(\tau)^{\dagger}&\mathbf{v}_{1}^{h}(\tau)^{\dagger}&\mathbf{v}_{2}^{t}(\tau)^{\dagger}\end{pmatrix}$, and $A_{l}$ and $R$ are complex matrices, with $R=R^{\dagger}$ and $\Re(R)\gg 0$. Thus we may rewrite (4.1) as
$\mathcal{I}_{0}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})_{l}$ $=$ $\frac{(2\,\pi)^{d_{G}/2}}{r\cdot V_{eff}(x_{1})}\cdot\tau^{d-1+d_{G}/2}\,e^{-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{1}(\tau)\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}_{2}^{t}(\tau)\right\|^{2}}$ (136)
$\cdot\int_{\mathbb{R}^{2d-2}}\,e^{-\imath\,\langle\mathbf{V}(\tau),\,A_{l}\,\mathbf{u}\rangle-\frac{1}{2}\,\mathbf{u}^{t}\,R\,\mathbf{u}}\,\mathrm{d}\mathbf{u}$
$=$ $\frac{(2\,\pi)^{d_{G}/2}}{r\cdot V_{eff}(x_{1})}\cdot\tau^{d-1+d_{G}/2}\,\frac{(2\,\pi)^{d-1}}{\mathrm{det}(R)^{\frac{1}{2}}}\,e^{-\frac{1}{2}\,\langle\mathbf{V}(\tau),D_{l}\,\mathbf{V}(\tau)\rangle},$
for a certain matrix $D=D^{\dagger}$. The ambiguity in the choice of the square root of the determinant of the matrix $R$ is resolved using analytic continuation arguments, as in Appendix A of [D]. Therefore, after performing the integral, in place of (4.3) we obtain
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}$ (138)
$\asymp$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot\mathrm{det}(R)^{-\frac{1}{2}}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$
$\cdot e^{\imath\,\sqrt{\lambda}\,\frac{1}{\tau}\,(\theta_{1}-\theta_{2})-\frac{1}{2}\,\langle\mathbf{V}(\tau),D_{l}\,\mathbf{V}(\tau)\rangle}.$
Furthermore, for $\theta_{j}=0$ and $\mathbf{V}=\mathbf{0}$ we obtain an asymptotic expansion for $\Pi_{k,\nu,\lambda}(x_{1},x_{2})$ which must agree with the one obtained setting $\theta_{j}=0$, $\mathbf{v}_{j}=\mathbf{v}_{j}^{t}=\mathbf{0}$ in the previous case. Hence $\det(R)=1$.
In order to verify that $\Re(D_{l})\gg 0$, let us consider the special case where $\nu$ is the trivial representation and $\theta_{j}=0$. Then $\mathcal{B}_{\nu}(x_{1},x_{2})_{l}=1$ and so
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ (139)
$\sim$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\sum_{l=1}^{r_{x_{1}}}e^{-\frac{1}{2}\,\langle\mathbf{V}(\tau),D_{l}\,\mathbf{V}(\tau)\rangle}$
$\cdot\left[1+O\left(\lambda^{-1/2}\right)\right]$
uniformly for $\|\mathbf{V}\|\leq C\,\lambda^{\epsilon^{\prime}}$.
On the other hand, in view of Definition 18, we have
$\delta^{\prime}_{\lambda}$ $:=$ $\max\left\{\mathrm{dist}_{X^{\tau}}\left(x_{1,\lambda},Z^{\tau}\right),\mathrm{dist}_{X^{\tau}}\left(x_{2,\lambda},Z^{\tau}\right)\right\}$
$\geq$ $\frac{1}{2\,\sqrt{\lambda}}\,\max\left\{\left\|\mathbf{v}_{1}^{t}\right\|,\left\|\mathbf{v}_{2}^{t}\right\|\right\},$
and
$\delta^{\prime\prime}_{\lambda}$ $:=$ $\mathrm{dist}_{X^{\tau}}\left(x_{1,\lambda},x_{2,\lambda}^{G\times\mathrm{supp}(\chi)}\right)$
$\geq$ $\mathrm{dist}_{X^{\tau}}\left(x_{1,\lambda},x_{2,\lambda}^{G}\right)\geq\frac{1}{2\,\sqrt{\lambda}}\,\left\|\mathbf{v}_{1}^{h}\right\|.$
Let us set $\delta_{\lambda}:=\max\{\delta^{\prime}_{\lambda},\,\delta^{\prime\prime}_{\lambda}\}$. If $\|\mathbf{V}\|=\lambda^{\epsilon^{\prime}}$ for some $\epsilon^{\prime}\in(0,1/6)$, we conclude that $\delta_{\lambda}\geq C\,\lambda^{\epsilon-1/2}$, and therefore $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})=O\left(\lambda^{-\infty}\right)$ by Theorem 7. However, this will be true only if every summand in (139) is rapidly decreasing, i.e. if $\Re(D_{l})\gg 0$ for every $l$.
The remaining arguments are a repetition of previous ones.
4.4 Proof of Theorem 28
We now consider near-diagonal scaling asymptotics, thus with $x_{1}=x_{2}=x$, and assume $\chi\in\mathcal{C}^{\infty}\big{(}(-\epsilon,\epsilon)\big{)}$ for some $\epsilon>0$ suitably small; then $t_{1}=0$, $x_{12}=x$, and $B=I_{2d-2}$ (the identity matrix) in (25). Furthermore, we may take $h_{1}=e$ in (39). We adopt the notation (27).
Proof of Theorem 28.
The exponent (4.1) becomes
$A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})_{l}$
$=$ $-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}-\left\|\mathbf{u}^{t}\right\|^{2}-\imath\,\omega_{0}\left(\mathbf{v}^{h}_{1}(\tau)^{(l)},\mathbf{u}^{h}\right)-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{1}(\tau)^{(l)}-\mathbf{u}^{h}\right\|^{2}$
$-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{u}^{t}\right)+\psi_{2}\big{(}\mathbf{u},\mathbf{v}_{2}(\tau)\big{)}$
$=$ $A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})^{h}_{l}+A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})^{vt}_{l},$
where
$A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})^{h}_{l}:=-\imath\,\omega_{0}\left(\mathbf{v}^{h}_{1}(\tau)^{(l)}-\mathbf{v}^{h}_{2}(\tau),\mathbf{u}^{h}\right)$
$-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{2}(\tau)\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{1}(\tau)^{(l)}\right\|^{2}-\left\|\mathbf{u}^{h}\right\|^{2}+g_{0}\left(\mathbf{v}^{h}_{1}(\tau)^{(l)}+\mathbf{v}^{h}_{2}(\tau),\mathbf{u}^{h}\right),$
$A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})^{vt}_{l}:=-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}-\left\|\mathbf{u}^{t}\right\|^{2}$
$-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{u}^{t}\right)-\frac{1}{2}\,\left\|\mathbf{u}^{v}\right\|^{2}-\omega_{0}\left(\mathbf{u}^{v},\mathbf{v}_{2}^{t}(\tau)\right)-\frac{1}{2}\,\left\|\mathbf{u}^{t}-\mathbf{v}_{2}^{t}(\tau)\right\|^{2},$
By standard Gaussian integrations, we obtain
$\int_{\mathbb{R}^{2\,(d-1-d_{G})}}e^{A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})^{h}_{l}}\,\mathrm{d}\mathbf{u}^{h}$ (143)
$=$ $e^{-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{2}(\tau)\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{v}^{h}_{1}(\tau)^{(l)}\right\|^{2}-\imath\,\omega_{0}\left(\mathbf{v}^{h}_{1}(\tau)^{(l)},\mathbf{v}^{h}_{2}(\tau)\right)+\frac{1}{4}\,\left\|\mathbf{v}^{h}_{1}(\tau)^{(l)}+\mathbf{v}^{h}_{2}(\tau)\right\|^{2}}$
$\cdot\int_{\mathbb{R}^{2\,(d-1-d_{G})}}e^{-\|\mathbf{a}\|^{2}-\imath\,\omega_{0}\big{(}\mathbf{v}_{1}^{h}(\tau)^{(l)}-\mathbf{v}_{2}(\tau),\mathbf{a}\big{)}}\,\mathrm{d}\mathbf{a}$
$=$ $\pi^{d-1-d_{G}}\,e^{\psi_{2}\left(\mathbf{v}^{h}_{1}(\tau)^{(l)},\mathbf{v}^{h}_{2}(\tau)\right)},$
$\int_{\mathbb{R}^{2\,d_{G}}}e^{A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})^{vt}_{l}}\,\mathrm{d}\mathbf{u}^{v}\,\mathrm{d}\mathbf{u}^{t}=e^{-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}}$
$\cdot\int_{\mathbb{R}^{d_{G}}}\,\mathrm{d}\mathbf{u}^{t}\,\left[e^{-\left\|\mathbf{u}^{t}\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{u}^{t}-\mathbf{v}_{2}^{t}(\tau)\right\|^{2}}\int_{\mathbb{R}^{d_{G}}}e^{-\imath\,\omega_{0}\left(\mathbf{u}^{v},\mathbf{u}^{t}+\mathbf{v}_{2}^{t}(\tau)\right)-\frac{1}{2}\,\left\|\mathbf{u}^{v}\right\|^{2}}\,\mathrm{d}\mathbf{u}^{v}\right]$
$=$ $(2\,\pi)^{d_{G}/2}\,e^{-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}}\cdot\int_{\mathbb{R}^{d_{G}}}\,e^{-\left\|\mathbf{u}^{t}\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{u}^{t}-\mathbf{v}_{2}^{t}(\tau)\right\|^{2}-\frac{1}{2}\,\left\|\mathbf{u}^{t}+\mathbf{v}_{2}^{t}(\tau)\right\|^{2}}\,\mathrm{d}\mathbf{u}^{t}$
$=$ $\pi^{d_{G}}\,e^{-\left\|\mathbf{v}_{1}^{t}(\tau)\right\|^{2}-\left\|\mathbf{v}_{2}^{t}(\tau)\right\|^{2}}.$
Hence, recalling (107), (4.1) reduces to
$\mathcal{I}_{0}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2})_{l}$
$=$ $\frac{(2\,\pi)^{d_{G}/2}}{r\cdot V_{eff}(x_{1})}\cdot\tau^{d-1+d_{G}/2}\,\int_{\mathbb{R}^{2d-2}}\,e^{A(\mathbf{v}_{1}(\tau),\mathbf{v}_{2}(\tau),\mathbf{u})_{l}}\,\mathrm{d}\mathbf{u}$
$=$ $\frac{(2\,\pi)^{d_{G}/2}}{r\cdot V_{eff}(x_{1})}\cdot\tau^{d-1+d_{G}/2}\,\pi^{d-1}\,e^{\frac{1}{\tau}\,\left[-\left\|\mathbf{v}_{1}^{t}\right\|^{2}-\left\|\mathbf{v}_{2}^{t}\right\|^{2}+\psi_{2}\left((\mathbf{v}^{h}_{1})^{(l)},\mathbf{v}^{h}_{2}\right)\right]}.$
Since $\sigma^{\tau}_{0,0}=1$ by (126), the leading order term in (4.1) becomes
$\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\big{[}I_{\lambda}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}\big{]}$
$\sim$ $e^{\imath\,\sqrt{\lambda}\,\frac{\theta_{1}-\theta_{2}}{\tau}}\cdot\frac{\lambda^{2d-3}}{(2\,\pi^{2}\,\tau^{2})^{d-1}}\cdot\sigma_{0,0}^{\tau}(x_{1})\cdot\chi(0)\cdot\overline{\Xi_{\nu}\left(\kappa_{l}\right)}$
$\cdot\frac{(2\,\pi)^{d_{G}/2}}{r\cdot V_{eff}(x_{1})}\cdot\tau^{d-1+d_{G}/2}\,\pi^{d-1}\,e^{\frac{1}{\tau}\,\left[-\left\|\mathbf{v}_{1}^{t}\right\|^{2}-\left\|\mathbf{v}_{2}^{t}\right\|^{2}+\psi_{2}\left((\mathbf{v}^{h}_{1})^{(l)},\mathbf{v}^{h}_{2}\right)\right]}$
$=$ $e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})-\left\|\mathbf{v}_{1}^{t}\right\|^{2}-\left\|\mathbf{v}_{2}^{t}\right\|^{2}+\psi_{2}\left((\mathbf{v}^{h}_{1})^{(l)},\mathbf{v}^{h}_{2}\right)\right]}\cdot\frac{\lambda^{2d-3}}{(2\,\pi\,\tau)^{d-1-d_{G}/2}}\cdot\frac{\chi(0)\cdot\overline{\Xi_{\nu}\left(\kappa_{l}\right)}}{r\cdot V_{eff}(x_{1})}.$
Inserting (4.4) in (4.1), we finally obtain that to leading order
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})_{l}$ $\asymp$ $\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\,\frac{\dim(\nu)}{\sqrt{2\,\pi}}\cdot\frac{\chi(0)\cdot\overline{\Xi_{\nu}\left(\kappa_{l}\right)}}{r\cdot V_{eff}(x_{1})}$
$\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})-\left\|\mathbf{v}_{1}^{t}\right\|^{2}-\left\|\mathbf{v}_{2}^{t}\right\|^{2}+\psi_{2}\left((\mathbf{v}^{h}_{1})^{(l)},\mathbf{v}^{h}_{2}\right)\right]},$
The argument for the lower order terms is the same as for Theorem 25.
5 Scaling asymptotics for $P_{k,\nu,\lambda}$
The previous techniques may be applied to the asymptotics of the complexifications of the equivariant eigenfunctions of the positive Laplacian operator $\Delta$ on $(M,\kappa)$ (recall Theorem 1). In this section we describe the necessary adaptations to the previous arguments.
As in (3), let the $\mu_{j}$’s be the distinct eigenvalues of $\sqrt{\Delta}$. Let us choose for every $j$ a real orthonormal basis $(\varphi_{j,k})_{k=1}^{\ell_{j}^{\prime}}$ of the eigenspace $W_{j}$ of $\mu_{j}$. Similarly, for every $j$ such that the equivariant component $W_{j,\nu}$ of $W_{j}$ is non-zero let $(\varphi_{j,\nu,k})_{k=1}^{\ell_{j,\nu}^{\prime}}$ be a real orthonormal basis of $W_{j,\nu}$.
Let us recall the following basic facts; a detailed discussion may be found in [Z-2012], [Z-2013], [Z-2020].
  1. 1.
    The wave operator $U(t)=e^{\imath\,t\,\sqrt{\Delta}}:L^{2}(M)\rightarrow L^{2}(M)$ at time $t\in\mathbb{R}$ is the unitary operator with distributional kernel
    $U(t)(m,n)=\sum_{j=1}^{+\infty}e^{\imath\,\mu_{j}\,t}\,\sum_{k=1}^{\ell^{\prime}_{j}}\,\varphi_{j,k}(m)\cdot\varphi_{j,k}(n).$
  2. 2.
    For every $\tau>0$, the distributional kernel
    $U(\imath\,\tau)(m,n):=\sum_{j=1}^{+\infty}e^{-\mu_{j}\,\tau}\,\sum_{k=1}^{\ell^{\prime}_{j}}\,\varphi_{j,k}(m)\cdot\varphi_{j,k}(n)$
    is globally real-analytic on $M\times M$; there exists $\tau_{1}\in(0,\tau_{0}]$ such that if $\tau\in(0,\tau_{1}]$ then $U(\imath\,\tau)$ admits a holomorphic extension on $M^{\tau}\times M$ in the first variable,
    $\widetilde{U}(\imath\,\tau)(x,n)=\sum_{j=1}^{+\infty}e^{-\mu_{j}\,\tau}\,\sum_{k=1}^{\ell^{\prime}_{j}}\,\widetilde{\varphi}_{j,k}(x)\cdot\varphi_{j,k}(n).$
  3. 3.
    For $\tau\in(0,\tau_{1})$, the restriction of $\widetilde{U}(\imath\,\tau)$ to $X^{\tau}\times M$ is the distributional kernel of an operator
    $\mathfrak{P}^{\tau}:\mathcal{C}^{\infty}(M)\rightarrow\mathcal{O}(X^{\tau}),$ (148)
    which is a Fourier integral operator with complex phase of degree (d1)/4-(d-1)/4;
  4. 4.
    $\mathfrak{P}^{\tau}$ controls the complexification of the eigenfunctions $\widetilde{\varphi}_{j,k}$: leaving restriction to $X^{\tau}$ implicit,
    $\widetilde{\varphi}_{j,k}=e^{\tau\,\mu_{j}}\,\mathfrak{P}^{\tau}(\varphi_{j,k}).$
  5. 5.
    For every $s\in\mathbb{R}$, $\mathfrak{P}^{\tau}$ determines a continuous isomorphism of Sobolev spaces
    $\mathfrak{P}^{\tau}:W^{s}(M)\rightarrow\mathcal{O}^{s+\frac{d-1}{4}}(X^{\tau}),$
    where $W^{s}(M)$ is the $s$-th Sobolev space of $M$, and $\mathcal{O}^{k}(X^{\tau})$ is the intersection of the $k$-th Sobolev space of $X^{\tau}$ with the space of CR (generalized) functions on $X^{\tau}$.
  6. 6.
    The composition $U_{\mathbb{C}}(t+2\,\imath\,\tau):=\mathfrak{P}^{\tau}\circ U(t)\circ{\mathfrak{P}^{\tau}}^{*}$ is a Fourier integral operator with complex phase of degree (d1)/2-(d-1)/2 on $X^{\tau}$, whose distributional kernel admits the spectral description
    $U_{\mathbb{C}}(t+2\,\imath\,\tau)(x,y)$ $=$ $\sum_{j}e^{(-2\,\tau+\imath\,t)\,\mu_{j}}\,\sum_{k=1}^{\ell^{\prime}_{j}}\,\widetilde{\varphi}_{j,k}(x)\,\overline{\widetilde{\varphi}_{j,k}(y)}$
    $=$ $\sum_{j}e^{\imath\,t\,\mu_{j}}\,\sum_{k=1}^{\ell^{\prime}_{j}}\,\mathfrak{P}^{\tau}\left(\widetilde{\varphi}_{j,k}\right)(x)\,\overline{\mathfrak{P}^{\tau}(\varphi)_{j,k}(y)};$
    we shall also use the notation $U_{\mathbb{C}}(t+2\,\imath\,\tau,x,y)=U_{\mathbb{C}}(t+2\,\imath\,\tau)(x,y)$.
  7. 7.
    For every $s\in\mathbb{R}$, $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ determines a continuous isomorphism of CR Sobolev spaces
    $U_{\mathbb{C}}(t+2\,\imath\,\tau):\mathcal{O}^{s}(X^{\tau})\rightarrow\mathcal{O}^{s+\frac{d-1}{2}}(X^{\tau}).$
For the following computations, we need the description (due to Zelditch) of of the complexified Poisson-wave operators $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ as a ‘dynamical Toeplitz operator’. The latter relies in turn on the study of the composition
$\mathfrak{Q}^{\tau}:={\mathfrak{P}^{\tau}}^{*}\circ\mathfrak{P}^{\tau}:\mathcal{C}^{\infty}(M)\rightarrow\mathcal{C}^{\infty}(M),$
where $\mathfrak{P}^{\tau}$ is as in (148). As explained in §3 of [Z-2007], $\mathfrak{Q}^{\tau}$ is an elliptic pseudodifferential operator on $M$, of degree (d1)/2-(d-1)/2, and its principal symbol $\sigma(\mathfrak{Q}^{\tau})$ appears in the description of $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ in terms of dynamical Toeplitz operators (see [Z-2012], [Z-2014], [Z-2020]).
Since $\mathfrak{P}^{\tau}$$^{*}$ depends on the choice of volume form on $X^{\tau}$, so does $\mathfrak{Q}^{\tau}$. In this section, we review the computation of $\sigma(\mathfrak{Q}^{\tau})$, which was carried out by Zelditch in §3 of [Z-2007], in light of the choices that we have adopted (see also the discussion in [P-2024]). We follow the general heuristic strategy in [Z-2007]: first consider the Euclidean case, and then reduce the general case to the latter.
Thus we first assume that $M=\mathbb{R}^{d}$ with the standard metric, and $X^{\tau}\cong\mathbb{R}^{d}\times S^{d-1}_{\tau}$ (where $S^{d-1}_{\tau}=\tau\,S^{d-1}$ is the sphere of radius $\tau$ centered at the origin).
In the real domain, the wave kernel at time $t$ for $\mathbb{R}^{d}$ with the standard metric is
$E_{t}(x,y):=\frac{1}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{\imath\,(t\,\|\xi\|+\langle\xi,x-y\rangle)}\,\mathrm{d}\xi.$ (150)
This may be analytically continued to the complex domain in $t$ and $x$ by replacing $t$ with $t+\imath\,\tau$ for $\tau>0$ and $x$ with $\zeta=x+\imath\,p\in\mathbb{C}^{d}$. In particular, for $t=0$ we obtain the kernel
$\mathfrak{E}^{\tau}(\zeta,y)=\frac{1}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{-\tau\,\|\xi\|-\langle\xi,p\rangle+\imath\,\langle\xi,x-y\rangle}\,\mathrm{d}\xi,$ (151)
which is absolutely convergent and holomorphic in $\zeta$ on the locus where $\|p\|<\tau$ (which plays the role of $M^{\tau}$). The distributional kernel of $\mathfrak{P^{\tau}}$ is given by the restriction of (151) to the locus where $\|p\|=\tau$ (which plays the role of $X^{\tau}$).
With respect to the given volume forms, $(\mathfrak{P}^{\tau})^{*}$ is represented by the integral kernel
$(\mathfrak{P}^{\tau})^{*}(y,\zeta)=\overline{\mathfrak{P}^{\tau}(\zeta,y)}=\frac{1}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{-\tau\,\|\xi\|-\langle\xi,p\rangle-\imath\,\langle\xi,x-y\rangle}\,\mathrm{d}\xi,$
where $\zeta=x+\imath\,p$.
Therefore $\mathfrak{Q}^{\tau}$ is represented by the operator kernel
$\mathfrak{Q}^{\tau}(x,y)=\int_{X^{\tau}}(\mathfrak{P}^{\tau})^{*}(x,\zeta)\,\,\mathfrak{P}^{\tau}(\zeta,y)\,\mathrm{d}V_{X^{\tau}}(\zeta).$ (152)
Let us write $\zeta=x^{\prime}+\imath\,\tau\,\omega$, where $x^{\prime}\in\mathbb{R}^{d}$ and $\omega\in S^{d-1}$. Then $\mathrm{d}V_{X^{\tau}}(x^{\prime})=\tau^{d-1}\,\mathrm{d}x^{\prime}\,\mathrm{d}\omega$. We obtain
$\mathfrak{Q}^{\tau}(x,y)$
$=$ $\frac{\tau^{d-1}}{(2\,\pi)^{2\,d}}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}x^{\prime}\,\int_{S^{d-1}}\mathrm{d}\omega\left[(\mathfrak{P}^{\tau})^{*}(x,\zeta)\,\,\mathfrak{P}^{\tau}(\zeta,y)\,\right]$
$=$ $\frac{\tau^{d-1}}{(2\,\pi)^{2\,d}}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}x^{\prime}\,\int_{S^{d-1}}\mathrm{d}\omega\,\int_{\mathbb{R}^{d}}\,\mathrm{d}\xi_{1}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}\xi_{2}$
$\left[e^{-\tau\,(\|\xi_{1}\|+\|\xi_{2}\|)-\tau\,\langle\xi_{1}+\xi_{2},\omega\rangle+\imath\,\langle\xi_{2},x^{\prime}-y\rangle-\imath\,\langle\xi_{1},x^{\prime}-x\rangle}\,\right]$
$=$ $\frac{\tau^{d-1}}{(2\,\pi)^{2\,d}}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}x^{\prime}\,\int_{S^{d-1}}\mathrm{d}\omega\,\int_{\mathbb{R}^{d}}\,\mathrm{d}\xi_{1}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}\xi_{2}$
$\left[e^{-\tau\,(\|\xi_{1}\|+\|\xi_{2}\|)-\tau\,\langle\xi_{1}+\xi_{2},\omega\rangle+\imath\,\langle\xi_{2}-\xi_{1},x^{\prime}\rangle+\imath\,\langle\xi_{1},x\rangle-\imath\,\langle\xi_{2},y\rangle}\,\right].$
Using the distributional identity
$\frac{1}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{\imath\,\langle\xi_{2}-\xi_{1},x^{\prime}\rangle}\,\mathrm{d}x^{\prime}=\delta(\xi_{2}-\xi_{1}),$
we can rewrite (5) as
$\mathfrak{Q}^{\tau}(x,y)=\frac{\tau^{d-1}}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{\imath\,\langle\xi,x-y\rangle}\left[\int_{S^{d-1}}e^{-2\,\tau\,\|\xi\|-2\,\tau\,\langle\xi,\omega\rangle}\,\mathrm{d}\omega\right]\mathrm{d}\xi.$ (154)
We aim to evaluate the inner integral asymptotically for $\xi\rightarrow\infty$. To this end, we set $\xi=\lambda\,\eta$, where $\lambda>0$ and $\eta\in S^{d-1}$, and let $\lambda\rightarrow+\infty$. We obtain
$\int_{S^{d-1}}e^{-2\,\tau\,\|\xi\|-2\,\tau\,\langle\xi,\omega\rangle}\,\mathrm{d}\omega$ $=$ $\int_{S^{d-1}}e^{\imath\lambda\cdot\Psi_{\eta}^{\tau}(\omega)}\,\mathrm{d}\omega,$ (155)
where
$\Psi_{\eta}^{\tau}(\omega):=\imath\,2\,\tau\,\big{(}1+\langle\eta,\omega\rangle\big{)}.$
Thus $\Psi_{\eta}^{\tau}$ is purely imaginary, and $\Im\left(\Psi_{\eta}^{\tau}\right)\geq 0$. Furthermore, $\Im\left(\Psi_{\eta}^{\tau}\right)$ vanishes only at $\omega=-\eta$. Hence, without altering the asymptotics for $\lambda\rightarrow+\infty$, we may replace integration over $S^{d-1}$ by integration over an arbitrarily small open neighbourhood $S_{\eta}\subset S^{d-1}$ of $\eta$.
Furthermore, given the form of $\Psi_{\eta}^{\tau}:S^{d-1}\rightarrow\mathbb{R}$, there is no loss of generality in assuming that $\eta$ is the last vector of the standard basis of $\mathbb{R}^{d}$, that is, $\eta=\begin{pmatrix}\mathbf{0}&1\end{pmatrix}^{\dagger}$. Thus any $\omega\in S_{\eta}$ may be written $\omega=\begin{pmatrix}\mathbf{u}&-\sqrt{1-\|\mathbf{u}\|^{2}}\end{pmatrix}^{\dagger}$, where $\mathbf{u}\in\mathbb{R}^{d-1}$ ranges in a small neighbourhood of the origin. Then $\mathrm{d}\omega=\mathcal{V}(\mathbf{u})\,\mathrm{d}\mathbf{u}$ on $S_{\eta}$ where $\mathcal{V}(\mathbf{0})=1$. For $\mathbf{u}\sim\mathbf{0}$,
$\Psi_{\eta}^{\tau}(\omega)=\imath\,2\,\tau\,\left(1-\sqrt{1-\|\mathbf{u}\|^{2}}\right)=\imath\,\tau\,\left(\|\mathbf{u}\|^{2}+R_{3}(\mathbf{u})\right).$ (156)
Thus there is a unique critical point $\mathbf{u}=\mathbf{0}$ (that is, $\omega=-\eta$), with Hessian matrix $H(\Psi_{\eta}^{\tau})=2\,\imath\,\tau\,I_{d-1}$.
Thus
$\sqrt{\det\left(\frac{\lambda\,H(\Psi_{\eta}^{\tau})}{2\pi\,\imath}\right)}=\left(\frac{\lambda\tau}{\pi}\right)^{\frac{d-1}{2}}.$
Applying the Lemma of stationary phase, we obtain for (155) an aymptotic expansion in descending powers of $\lambda=\|\xi\|$, with leading order term $(\pi/\lambda\tau)^{(d-1)/2}$. In view of (154) the principal symbol of $\mathfrak{Q}^{\tau}$ is therefore
$\sigma(\mathfrak{Q}^{\tau})=\tau^{d-1}\,\left(\frac{\pi}{\lambda\,\tau}\right)^{\frac{d-1}{2}}=\left(\frac{\pi\,\tau}{\|\xi\|}\right)^{\frac{d-1}{2}}.$ (157)
Before considering the general case, let us premise a remark considering scalar rescalings of a metric. Suppose $X=X^{\tau}$ and consider the operator $\mathfrak{P}^{\tau}:\mathcal{C}^{\infty}(M)\rightarrow\mathcal{O}(X)$. To emphasize the role of the metric, let us write $\mathfrak{P}^{\tau}=\mathfrak{P}^{\tau}_{\kappa}$, $X=X^{\tau}_{\kappa}$.
When $\kappa$ is replaced by $\kappa^{\lambda}:=\lambda^{2}\,\kappa$ for some $\lambda\in\mathbb{R}_{+}$, we have $X=X^{\lambda\,\tau}_{\kappa^{\lambda}}$. If $\Delta_{\kappa}$ and $\Delta_{\kappa^{\lambda}}$ are the Laplacians for $\kappa$ and $\kappa^{\lambda}$, respectively, then $\Delta_{\kappa^{\lambda}}=\lambda^{-2}\,\Delta_{\kappa}$. Hence,
$e^{-\tau\,\sqrt{\Delta_{\kappa}}}=e^{-\lambda\,\tau\,\sqrt{\Delta_{\kappa^{\lambda}}}}.$
Thus it makes sense to denote $\mathfrak{P}^{\tau}_{\kappa}$ as $\mathfrak{P}^{X}$, without reference to a specific rescaling of $\kappa$.
On the other hand, the rescaling affects the adjoint operator, since it modifies the volume form on $X$. Let $\mathrm{vol}_{X,\kappa}$ and let $\mathrm{vol}_{X,\kappa^{\lambda}}$ be the volume forms on $X$ viewed as $X^{\tau}_{\kappa}$ and $X^{\lambda\,\tau}_{\kappa^{\lambda}}$. Similarly, let $\mathrm{vol}_{M,\kappa}$ and $\mathrm{vol}_{M,\kappa^{\lambda}}$ be the Riemannian volume forms on $M$ associated to $\kappa$ and $\kappa^{\lambda}$, respectively. Then
$\mathrm{vol}_{M,\kappa^{\lambda}}=\lambda^{d}\,\mathrm{vol}_{M,\kappa},\quad\mathrm{vol}_{X,\kappa^{\lambda}}=\lambda^{2d-1}\,\mathrm{vol}_{X,\kappa}.$ (158)
Let $({\mathfrak{P}}^{X})^{*}_{\kappa}$ and $({\mathfrak{P}}^{X})^{*}_{\kappa^{\lambda}}$ be the adoints of $\mathfrak{P}^{X}$ with respect to $\kappa$ and $\kappa^{\lambda}$ (that is, using the pairs of volume forms $(\mathrm{vol}_{M,\kappa},\mathrm{vol}_{X,\kappa})$, $(\mathrm{vol}_{M,\kappa^{\lambda}},\mathrm{vol}_{X,\kappa^{\lambda}})$, respectively). One concludes from (158) that
$({\mathfrak{P}}^{X})^{*}_{\kappa^{\lambda}}=\lambda^{d-1}\,({\mathfrak{P}}^{X})^{*}_{\kappa}.$
Since $X=X^{\tau}_{\kappa}=X^{\lambda\,\tau}_{\kappa^{\lambda}}$, we obtain
$\mathfrak{Q}^{\lambda\,\tau}_{\kappa^{\lambda}}=\lambda^{d-1}\,\mathfrak{Q}^{\tau}_{\kappa};$ (159)
therefore the same relation holds between the respective principal symbols. Since $\mathfrak{Q}^{\tau}_{\kappa}$ has degree (d1)/2-(d-1)/2, we conclude that its principal symbol has the form
$\sigma(\mathfrak{Q}^{\tau}_{\kappa})(m,\xi)=c_{m}(\xi)\,\left(\frac{\tau\,\pi}{\|\xi\|}\right)^{\frac{d-1}{2}},$ (160)
where $c_{m}(\xi)$ is homogeneous of degree 0 in $\xi$.
Let us consider a general real-analytic $(M,\kappa)$ and fix $m\in M$. Let us choose a real-analytic coordinate chart $\varphi:B\rightarrow U$ centered at $m$ that is isometric at the origin. Thus, $B\subseteq\mathbb{R}^{d}$ is an open neighborhood of the origin (say, an open ball centered at $\mathbf{0}$), $U$ an open neighbourhood of $m$ in $M$, and $\varphi$ a real-analytic diffeomorphism such that $\mathrm{d}_{\mathbf{0}}\varphi$ is a linear isometry between $(\mathbb{R}^{d},g_{st})$ (where $g_{st}$ is the standard Euclidean product) and $(T_{m}M,\kappa_{x})$. The pull-back metric $\varphi^{*}(\kappa)$ admits a convergent power series expansion
$\varphi^{*}(\kappa)_{\mathbf{x}}=g_{st}+\sum_{|I|\geq 1}x^{I}\,g_{I}\qquad(\mathbf{x}\in B),$
where the $g_{I}$’s are fixed symmetric 2-tensors. Here $x^{I}=x_{1}^{i_{1}}\cdots x_{d}^{i_{d}}$ if $\mathbf{x}=\begin{pmatrix}x_{1}&\ldots&x_{d}\end{pmatrix}^{\dagger}$.
For some sufficiently small $\epsilon>0$ let $U_{\epsilon}:=\varphi\big{(}B_{d}(\mathbf{0},\epsilon)\big{)}$ denote be the image of the open ball of radius $\epsilon$. Since the singular support of $\mathfrak{Q}^{\tau}$ is the diagonal, the computation of the principal symbol of $\mathfrak{Q}^{\tau}$ at $(x,\xi)$ may be localized to $U_{\epsilon}$, meaning that in local coordinates it is given by the leading order term of the asymptotic expansion for $\xi\rightarrow\infty$ of the integral
$\int_{B_{d}(\mathbf{0},\epsilon)}e^{\imath\,\langle\xi,\mathbf{y}\rangle}\,\mathfrak{Q}^{\tau}(\mathbf{0},\mathbf{y})\,\rho(\epsilon^{-1}\,\mathbf{y})\,\mathcal{V}(\mathbf{y})\,\mathrm{d}\mathbf{y},$
where $\rho$ is a suitable fixed cut-off function identically equal to $1$ near the origin, and $\mathcal{V}(\mathbf{y})\,\mathrm{d}\mathbf{y}$ is the pull-back by $\varphi$ of the Riemannian density on $M$ (here we occasionally blend intrinsic and local coordinate notation). If we pull this back by the dilation $\rho_{\epsilon}:\mathbf{x}\in B_{d}(\mathbf{0},1)\mapsto\mathbf{y}:=\epsilon\,\mathbf{x}\in B_{d}(\mathbf{0},\epsilon)$, we obtain the symbol at $(0,\epsilon\,\xi)$ of the corresponding $\mathfrak{Q}^{\tau}$ referred to the metric
$(\varphi\circ\rho_{\epsilon})^{*}(\kappa)=\epsilon^{2}\,g_{\epsilon},\quad\text{where}\quad g_{\epsilon}:=g_{st}+\sum_{|I|\geq 1}\epsilon^{|I|}\,x^{I}\,g_{I}.$ (161)
Now $g_{\epsilon}$ is a real-analytic Riemannian metric on $B_{d}(\mathbf{0},1)$, it is defined for sufficiently small $\epsilon$ and depends real analytically on $\epsilon$; hence the same holds of the corresponding operators $\mathfrak{Q}_{g_{\epsilon}}^{\tau}$ and their principal symbols. By (157), (160), and (161) we conclude that
$\sigma\left(\mathfrak{Q}_{g_{\epsilon}}^{\tau}\right)=\left(\frac{\pi\,\tau}{\|\xi\|}\right)^{\frac{d-1}{2}}\cdot\big{(}1+\epsilon\,F_{m}(\xi)\big{)},$
where $F_{m}$ is homogeneous of degree 0 in $\xi$.
On the other hand, in view of (159) on $U_{\epsilon}$ we have
$\sigma\left(\mathfrak{Q}^{\tau}_{(\varphi\circ\rho_{\epsilon})^{*}(\kappa)}\right)(m,\epsilon\,\xi)=\epsilon^{d-1}\,\sigma\left(\mathfrak{Q}^{\tau/\epsilon}_{g_{\epsilon}}\right)(x,\epsilon\,\xi)=\left(\frac{\pi\,\tau}{\|\xi\|}\right)^{\frac{d-1}{2}}\cdot\big{(}1+\epsilon\,F_{m}(\xi)\big{)}.$
Since the result must be independent of $\epsilon$, we conclude that $F_{m}=0$.
Summing up, we conclude the following:
Lemma 81.
The principal symbol of $\mathfrak{Q}^{\tau}$ is
$\sigma(\mathfrak{Q}^{\tau})(m,\xi)=\left(\frac{\pi\,\tau}{\|\xi\|}\right)^{\frac{d-1}{2}}.$
Going over the arguments in §2, §3 and especially §4 of [Z-2014] or §5, §6 and especially §7 of [Z-2020] in light of Lemma 81, one obtains the following description of $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ as a dynamical Toeplitz operator.
  1. 1.
    Let us denote by $\Pi^{\tau}_{-t}$ the zeroth order Fourier integral operator on $X^{\tau}$ having distributional kernel
    $\Pi^{\tau}_{-t}(x,y):=\Pi^{\tau}\left(\Gamma^{\tau}_{-t}(x),y\right);$
    then on $X^{\tau}$ there exist a smoothly varying pseudodifferential operator $Q^{\tau}_{t}$ of degree (d1)/2-(d-1)/2 and a smoothly varying operator $R^{\tau}_{t}$ with $\mathcal{C}^{\infty}$ kernel such that
    $U_{\mathbb{C}}(t+2\,\imath\,\tau)=\Pi^{\tau}\circ Q^{\tau}_{t}\circ\Pi^{\tau}_{-t}+R^{\tau}_{t}.$ (162)
  2. 2.
    In a conic neighoburhood of the symplectic cone $\Sigma^{\tau}$, $Q^{\tau}_{t}$ admits the following microlocal description. Let us set $D^{\tau}_{\sqrt{\rho}}:=\imath\,\upsilon_{\sqrt{\rho}}$. Then there exists a polyhomogeneous classical symbol on $X^{\tau}\times\mathbb{R}_{+}$, of the form
    $\gamma^{\tau}_{t}(x,r)\sim\sum_{j\geq 0}\gamma^{\tau}_{t,j}(x)\,r^{-\frac{d-1}{2}-j},$ (163)
    such that $Q^{\tau}_{t}\sim\gamma^{\tau}_{t}\big{(}x,D^{\tau}_{\sqrt{\rho}}\big{)}$.
  3. 3.
    Similarly to (58), the leading coefficient in (163) is
    $\gamma^{\tau}_{t,0}(x)=(\pi\,\tau)^{\frac{d-1}{2}}\cdot e^{\imath\,\tilde{\theta}^{\tau}_{t}(x)}\cdot\langle\sigma_{J}^{(x)},\sigma_{J_{t}}^{(x)}\rangle^{-1}$ (164)
    for a certain smooth function $\tilde{\theta}^{\tau}_{t}:X^{\tau}\rightarrow\mathbb{R}$. The additional factor $\tau^{\frac{d-1}{2}}$ with respect to the computation of Zelditch (see e.g. [Z-2012]) is due to our choice of volume form, which affects the construction of $P^{\tau}$$^{*}$ (see the discussion in [P-2024]).
Proof of Theorem 29.
We are interested in the asymptotics of the kernel
$P^{\tau}_{\chi,\lambda}(x,y):=\sum_{j}\hat{\chi}(\lambda-\mu_{j})\,e^{-2\,\tau\,\mu_{j}}\,\sum_{k}\widetilde{\varphi}^{\tau}_{j,k}(x)\,\overline{\widetilde{\varphi}^{\tau}_{j,k}(y)}\qquad(x,y\in X^{\tau}),$
and more generally of its equivariant version
$P^{\tau}_{\chi,\nu,\lambda}(x,y):=\sum_{j}\hat{\chi}(\lambda-\mu_{j})\,e^{-2\,\tau\,\mu_{j}}\,\sum_{k}\widetilde{\varphi}^{\tau}_{j,\nu,k}(x)\,\overline{\widetilde{\varphi}^{\tau}_{j,\nu,k}(y)}.$ (165)
The analysis run parallel to the one conducted for $\Pi^{\tau}_{\chi,\nu,\lambda}$.
More precisely, the following analogue of (55) describes the relation between $P^{\tau}_{\chi,\lambda}$ and $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ in (6) is given by :
$P^{\tau}_{\chi,\lambda}(x,y)=\frac{1}{\sqrt{2\,\pi}}\,\int_{-\infty}^{+\infty}e^{-\imath\,\lambda\,t}\,\chi(t)\,U_{\mathbb{C}}(t+2\,\imath\,\tau,x,y)\,\mathrm{d}t;$ (166)
furthermore, operatorially we have the analogue of (18):
$P^{\tau}_{\chi,\nu,\lambda}=P_{\nu}\circ P^{\tau}_{\chi,\lambda}.$ (167)
Arrguing as in the proof of Theorem 7, with $U_{\mathbb{C}}(t+2\,\imath\,\tau,x,y)$ in place of $U^{\tau}_{\sqrt{\rho}}(t;x,y)$ and (162) in place of (56), we obtain in place of (3.2):
$P^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})$
$\sim$ $\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{G}\,\mathrm{d}V_{G}(g)\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\left[\Xi_{\nu}\left(g^{-1}\right)\,e^{-\imath\,\lambda\,t}\,\chi(t)\,\left(\Pi^{\tau}\circ Q^{\tau}_{t}\circ\Pi^{\tau}_{-t}\right)\left(\mu^{\tau}_{g^{-1}}(x_{1}),x_{2}\right)\right].$
The arguments in the proof of Theorem 7 apply, except that the leading term of the amplitude has been multiplied by a factor $(\pi\,\tau)^{(d-1)/2}\,(u\,\tau)^{-(d-1)/2}$ and the unitary factor $e^{\imath\,\tilde{\theta}^{\tau}_{t}(x)}$ in (164) replaces $e^{\imath\,\theta^{\tau}_{t}(x)}$. We have used that the principal symbol of $D^{\tau}_{\sqrt{\rho}}$ along $\Sigma^{\tau}$ (or, equivalently, of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$) is
$\sigma(D^{\tau}_{\sqrt{\rho}})\left(x,v\,\alpha_{x}^{\tau}\right)=v\,\tau\qquad(v>0).$
In view of the rescaling $u\mapsto\lambda\,u$, this entails a further factor $\lambda^{-(d-1)/2}$. Furthermore, at the critical point (101) of the phase the product $u\,\tau=1$. Thus to leading order we obtain an extra overall factor $(\lambda/\pi\,\tau)^{-(d-1)/2}$. ∎
6 Near-graph uniform asymptotic expansions
Given $(x_{1},x_{2})\in Z^{\tau}\times Z^{\tau}$, we have defined $\Sigma_{\chi}(x_{1},x_{2})\subseteq G\times\mathrm{supp}(\chi)$ in Remark 59.
Lemma 82.
$\Sigma_{\chi}(x_{1},x_{2})$ has the following properties:
  1. 1.
    $\Sigma_{\chi}(x_{1},x_{2})\neq\emptyset$ if and only if $(x_{1},x_{2})\in\mathfrak{X}^{\tau}_{\chi}$;
  2. 2.
    for any $(x_{1},x_{2})\in Z^{\tau}\times Z^{\tau}$ and any neighbourhood $S$ of $\Sigma_{\chi}(x_{1},x_{2})$ in $G\times\mathrm{supp}(\chi)$, there exists a neighbourhood $Z^{\prime}$ of $(x_{1},x_{2})$ in $Z^{\tau}\times Z^{\tau}$ such that
    $(x_{1}^{\prime},x_{2}^{\prime})\in Z^{\prime}\quad\Rightarrow\quad\Sigma_{\chi}(x_{1}^{\prime},x_{2}^{\prime})\subseteq S.$
Proof.
The first statement is obvious by definition. If the second was false, for any $j=1,2,\ldots$ there would exist $(y_{j}^{\prime},y_{j}^{\prime\prime})\in Z^{\tau}\times Z^{\tau}$ having distance <1/j<1/j from $(x_{1},x_{2})$ and $(g_{j},t_{j})\in\Sigma_{\chi}(y_{j}^{\prime},y_{j}^{\prime\prime})$ having distance from $\Sigma_{\chi}(x_{1},x_{2})$ no less than $\epsilon_{0}$, for some fixed $\epsilon_{0}>0$. By compactness, we may assume without loss that $g_{j}\rightarrow g_{\infty}\in G$ and $t_{j}\rightarrow t_{\infty}$ in $\mathrm{supp}(\chi)$. By continuity,
$y_{j}^{\prime}=\mu^{\tau}_{g_{j}}\circ\Gamma^{\tau}_{t_{j}}(y_{j}^{\prime\prime})\quad\Rightarrow\quad x_{1}=\mu^{\tau}_{g_{\infty}}\circ\Gamma^{\tau}_{t_{\infty}}(x_{2})\quad\Rightarrow\quad(g_{\infty},t_{\infty})\in\Sigma_{\chi}(x_{1},x_{2}).$
hence $(g_{j},t_{j})\rightarrow(g_{\infty},t_{\infty})\in\Sigma_{\chi}(x_{1},x_{2})$, absurd. ∎
Proof of Theorem 30.
To begin with, we consider the asymptotics at fixed points (that is, with no rescaling). Let us choose $(x_{1},x_{2})\in\mathfrak{X}^{\tau}_{\chi}$, hence satisfying (38): there exist $(g,t)\in G\times\mathrm{supp}(\chi)$ such that $x_{1}=\mu^{\tau}_{g}\circ\Gamma^{\tau}_{t}(x_{2})$. We may assume that the possible pairs $(g,t)$ can be listed as in (39). We have
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})$
$\sim$ $\lambda^{2}\,\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{G}\,\mathrm{d}V_{G}(g)\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{X^{\tau}}\,\mathrm{d}V_{X^{\tau}}(y)\,\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{0}^{+\infty}\,\mathrm{d}v$
$\left[\Xi_{\nu}\left(g^{-1}\right)\,\chi(t)\,e^{\imath\,\lambda\,\left[u\,\psi^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y\right)+v\,\psi^{\tau}\left(\Gamma^{\tau}_{-t}(y),x_{2}\right)-t\right]}\right$.
$\left.\varrho_{1}(g,y)\,\varrho_{2}(t,y)\,s^{\tau}\left(\mu^{\tau}_{g^{-1}}(x_{1}),y,\lambda\,u\right)\,r^{\tau}_{t}\left(y,x_{2},\lambda\,v\right)\right].$
Let us multiply the integrand in (6) by the unrescaled cut-off
$\gamma^{\mathbb{R}}\left(t-t_{1}\right)\cdot\sum_{l=1}^{r_{x_{1}}}\gamma^{\mathfrak{g}}\left(\log_{G}\left(g\,h_{1}^{-1}\,\kappa_{l}^{-1}\right)\right),$ (170)
where $G_{x_{1}}=\{\kappa_{1},\ldots,\kappa_{r_{x_{1}}}\}$. Thus integration in $G\times\mathbb{R}$ has been restricted to a small but fixed neighbourhood of $\Sigma_{\chi}(x_{1},x_{2})$, and only a rapidly decreasing contribution to the asymptotics of (6) is lost. The same will then be true for the asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1}^{\prime},x_{2}^{\prime})$ for any $(x^{\prime}_{1},x_{2}^{\prime})$ in a fixed small neighbourhood of $(x_{1},x_{2})$, as $\Sigma_{\chi}(x_{1}^{\prime},x_{2}^{\prime})$ is then contained in a small neighbourhood of $\Sigma_{\chi}(x_{1},x_{2})$ by Lemma 82.
Thus, uniformly on an open neighbourhood of $(x_{1},x_{2})$ we have
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1}^{\prime},x_{2}^{\prime})\sim\sum_{l=1}^{r_{x_{1}}}\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1}^{\prime},x_{2}^{\prime})_{l},$
where for each $l$ (with the change of variable $t\mapsto t_{1}+t$) we have
$\Pi^{\tau}_{\chi,\nu,\lambda}(x^{\prime}_{1},x^{\prime}_{2})_{l}$
$=$ $e^{-\imath\,\lambda\,t_{1}}\,\lambda^{2}\,\frac{\dim(\nu)}{\sqrt{2\,\pi}}\,\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{-\infty}^{+\infty}\,\mathrm{d}\theta\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{0}^{+\infty}\,\mathrm{d}v$
$\left[e^{\imath\,\lambda\,\Psi_{x^{\prime}_{1},x^{\prime}_{2}}(t,v,\theta,u,\mathbf{v},\xi)_{l}}\,\mathcal{H}_{\lambda}(x^{\prime}_{1},x^{\prime}_{2},t,v,\theta,u,\mathbf{v},\xi)_{l}\right].$
At $(x_{1},x_{2})$ we have
$\Psi_{x_{1},x_{2}}(t,v,\theta,u,\mathbf{v},\xi)_{l}:=-u\,\theta+v\,(\theta+\tau\,t)-t$ (172)
$+\imath\,u\,\left[\frac{1}{4\,\tau^{2}}\,\theta^{2}-\psi_{2}\left(-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{v}\right)\right]+\imath\,v\,\left[\frac{1}{4\,\tau^{2}}\,(\theta+\tau\,t)^{2}+\frac{1}{2}\big{\|}B\mathbf{v}\big{\|}^{2}\right]$
$+R_{3}(\mathbf{v},\xi,\theta,t)$
and
$\mathcal{H}_{\lambda}(x_{1},x_{2},t,v,\theta,u,\mathbf{v},\xi)_{l}$ (173)
$:=$ $\Xi_{\nu}\left(h_{1}^{-1}\,\kappa_{l}^{-1}\,e^{-\xi}\right)\,\chi\left(t+t_{1}\right)\cdot\mathcal{V}\left(\theta,\mathbf{u}\right)\cdot\gamma(t,\theta,\xi,\mathbf{u})\cdot f_{1}(v)\cdot f_{2}(u)$
$\cdot s^{\tau}\left(\mu^{\tau}_{(e^{\xi}\kappa_{l})^{-1}}(x),y_{\lambda}(\theta,\mathbf{u}),\lambda\,u\right)\,r^{\tau}_{t}\big{(}y_{\lambda}(\theta,\mathbf{u}),x,\lambda\,v\big{)},$
where $\gamma(t,\theta,\xi,\mathbf{u})$ is a product of unrescaled cut-offs and $f_{1},\,f_{2}$ are as in Proposition 72. In particular, integration is restricted to a small neighbourhood of the locus where $\theta=t=0$, $\mathbf{v}=\xi=0$, and is compactly supported in $(u,v)$.
We have $\Im\Psi_{x_{1},x_{2}}(t,v,\theta,u,\mathbf{v},\xi)_{l}\geq 0$, and $\Im\Psi_{x_{1},x_{2}}(t,v,\theta,u,\mathbf{v},\xi)_{l}=0$ only if
$\theta=t=0,\quad\mathbf{v}=\mathbf{0}_{\mathbb{R}^{d_{G}}},\quad\xi=0_{\mathfrak{g}}.$
We have in addition a stationary point, if we also impose $u=v=1/\tau$. Thus the only nonnegligible contribution to the asymptotics may come from a neighourhood of
$P_{0}:=(t_{0},v_{0},\theta_{0},u_{0},\mathbf{v}_{0},\xi_{0})=\left(0,\frac{1}{\tau},0,\frac{1}{\tau},\mathbf{0}_{\mathbb{R}^{d_{G}}},0_{\mathfrak{g}}\right)$
Let $D_{x_{1}}^{l}$} denote the $(2d-2)\times d_{G}$ matrix representing the injective linear map $\xi\mapsto\xi_{X^{\tau}}(x_{1})^{(l)}$ with respect to the given orthonormal basis in $\mathfrak{g}$ and $\mathcal{H}^{\tau}_{x_{1}}$. At the critical point, the Hessian of $\Psi_{x_{1},x_{2}}(\cdot)_{l}$ is
$H_{P_{0}}\big{(}\Psi_{x_{1},x_{2}}(\cdot)_{l}\big{)}=\begin{pmatrix}\imath/2&\tau&\imath/(2\,\tau^{2})&0&\mathbf{0}^{\dagger}&\mathbf{0}^{\dagger}\\ \tau&0&1&0&\mathbf{0}^{\dagger}&\mathbf{0}^{\dagger}\\ \imath/(2\,\tau^{2})&1&\imath/\tau^{3}&-1&\mathbf{0}^{\dagger}&\mathbf{0}^{\dagger}\\ 0&0&-1&0&\mathbf{0}^{\dagger}&\mathbf{0}^{\dagger}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathcal{A}&\mathcal{B}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathcal{B}^{\dagger}&\mathcal{C}\end{pmatrix}$
where
$\mathcal{A}:=\frac{\imath}{\tau}\,\left(I_{2d-2}+B^{\dagger}\,B\right),\quad\mathcal{B}:=\frac{\imath}{\tau}\,(I_{2d-2}+\imath\,J_{0})\,{D_{x_{1}}^{l}},\quad\mathcal{C}:=\frac{\imath}{\tau}\,{D_{x_{1}}^{l}}^{\dagger}\,{D_{x_{1}}^{l}}.$
Here $J_{0}$ is the standard complex structure on $\mathbb{C}^{d-1}\cong\mathbb{R}^{d-1}\oplus\mathbb{R}^{d-1}$. Thus the Hessian matrix is in block diagonal form, and the determinants of the two blocks are as follows.
First we have
$\det\left(\frac{1}{\imath}\,\begin{pmatrix}\imath/2&\tau&\imath/(2\,\tau^{2})&0\\ \tau&0&1&0\\ \imath/(2\,\tau^{2})&1&\imath/\tau^{3}&-1\\ 0&0&-1&0\end{pmatrix}\right)=\tau^{2}.$
As to the second block, we have
$\det\left(\frac{1}{\imath}\,\begin{pmatrix}\mathcal{A}&\mathcal{B}\\ \mathcal{B}^{\dagger}&\mathcal{C}\end{pmatrix}\right)$
$=$ $\det\left(\frac{1}{\imath}\,\begin{pmatrix}\frac{\imath}{\tau}\,\left(I_{2d-2}+B^{\dagger}\,B\right)&\frac{\imath}{\tau}\,(I_{2d-2}+\imath\,J_{0})\,{D_{x_{1}}^{l}}\\ \frac{\imath}{\tau}\,{D_{x_{1}}^{l}}^{\dagger}\,(I_{2d-2}-\imath\,J_{0})&\frac{\imath}{\tau}\,{D_{x_{1}}^{l}}^{\dagger}\,{D_{x_{1}}^{l}}\end{pmatrix}\right)$
$=$ $\frac{1}{\tau^{2d-2+d_{G}}}\,\det\begin{pmatrix}\left(I_{2d-2}+B^{\dagger}\,B\right)&(I_{2d-2}+\imath\,J_{0})\,{D_{x_{1}}^{l}}\\ {D_{x_{1}}^{l}}^{\dagger}\,(I_{2d-2}-\imath\,J_{0})&{D_{x_{1}}^{l}}^{\dagger}\,{D_{x_{1}}^{l}}\end{pmatrix}$
$=$ $\frac{1}{\tau^{2d-2+d_{G}}}\,\det\begin{pmatrix}\left(I_{2d-2}+B^{\dagger}\,B\right)&(I_{2d-2}+\imath\,J_{0})\,{D_{x_{1}}^{l}}\\ 0&S\end{pmatrix}$,
where
$S$ $:=$ ${D_{x}^{l}}^{\dagger}\,{D_{x_{1}}^{l}}-{D_{x_{1}}^{l}}^{\dagger}\,(I_{2d-2}-\imath\,J_{0})\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,(I_{2d-2}+\imath\,J_{0})\,{D_{x_{1}}^{l}}.$
Thus $S=S^{\dagger}$, and the real part of $S$ is
$\Re(S)$ $:=$ ${D_{x_{1}}^{l}}^{\dagger}\,{D_{x}^{l}}-{D_{x}^{l}}^{\dagger}\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,\,{D_{x}^{l}}$
$-{D_{x_{1}}^{l}}^{\dagger}\,J_{0}\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,J_{0}\,{D_{x_{1}}^{l}}.$
We have (since $B$ is a symplectic matrix)
$\left(I_{2d-2}+B^{\dagger}\,B\right)\,J_{0}$ $=$ $J_{0}+B^{\dagger}\,B\,J_{0}$
$=$ $J_{0}+J_{0}\,\left(B^{\dagger}\,B\right)^{-1}$
$=$ $J_{0}\,\left(I_{2d-2}+\left(B^{\dagger}\,B\right)^{-1}\right).$
Hence, taking inverses we get
$J_{0}\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}=\left(I_{2d-2}+\left(B^{\dagger}\,B\right)^{-1}\right)^{-1}\,J_{0}.$
Thus
$\Re(S)$ $:=$ ${D_{x_{1}}^{l}}^{\dagger}\,{D_{x_{1}}^{l}}-{D_{x_{1}}^{l}}^{\dagger}\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,\,{D_{x_{1}}^{l}}$
$-{D_{x_{1}}^{l}}^{\dagger}\,\left(I_{2d-2}+\left(B^{\dagger}\,B\right)^{-1}\right)^{-1}\,J_{0}\,J_{0}\,{D_{x_{1}}^{l}}$
$=$ ${D_{x_{1}}^{l}}^{\dagger}\,\left[I_{2d-2}-\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\right]\,\,{D_{x_{1}}^{l}}$
$+{D_{x_{1}}^{l}}^{\dagger}\,\left(I_{2d-2}+\left(B^{\dagger}\,B\right)^{-1}\right)^{-1}\,{D_{x_{1}}^{l}},$
whence $\Re(S)\gg 0$.
Hence $P_{0}$ is a non-degenerate stationary point; by the complex stationary phase lemma of [MS], we obtain an asymptotic expansion for $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})$ which must agree with our previous derivation. Moreover, since the complex stationary phase holds with parameters, we can replace $(x_{1},x_{2})$ by a general $(x_{1}^{\prime},x_{2}^{\prime})$ varying in some small open neighbourhood of $(x_{1},x_{2})$, and obtain an asymptotic expansions
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1}^{\prime},x_{2}^{\prime})\sim\sum_{l=1}^{r_{x_{1}}}\Pi^{\tau}_{\chi,\nu,\lambda}(x^{\prime}_{1},x_{2}^{\prime})_{l},$ (174)
but for some $l$ the corresponding stationary point (accounting for the expansion of the $l$-th summand in (174)) might cease to be real; this happens when $(x_{1}^{\prime},x_{2}^{\prime})\in\mathfrak{X}_{\chi}^{\tau}$ but $r_{x_{1}^{\prime}}<r_{x_{1}}$, or when $(x_{1}^{\prime},x_{2}^{\prime})\not\in\mathfrak{X}_{\chi}^{\tau}$. In this case, $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1}^{\prime},x_{2}^{\prime})_{l}=O\left(\lambda^{-\infty}\right)$.
In particular, let us now replace $x_{j}$ by, say, $x_{j}^{\prime}:=x_{j}+(\theta_{j},\mathbf{v}_{j}^{t})$ for some nearby $x_{j}^{\prime}=x_{j}+(\theta_{j},\mathbf{v}_{j})$, with $(\theta_{j},\mathbf{v}_{j})$ suitably small. By non-degeneracy, the critical point will vary smoothly, except that it will generically move to the complex domain, which accounts for the exponential decay in the scaling asymptotics. Nonetheless, the asymptotic expansions will still hold, and if we pair this with the scaling asymptotics in the previous theorems, we obtain the statement of Theorem 30.
Remark 83.
In our situation, $R_{3}$ and $L_{\nu,l,s}\left(x_{1},x_{2};\cdot\right),\,K_{\nu,l,s}\left(x_{1},x_{2};\cdot\right)$ are complex valued real-analytic functions (meaning that their real and imaginary components are real-analytic). Hence there are expansions of the form
$e^{\lambda\,R_{3}\left(\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)}$ $=$ $\sum_{j\geq 0}\frac{\lambda^{j}}{j!}\,R_{3}\left(\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)^{j}$
$=$ $\sum_{j\geq 0}\frac{\lambda^{j}}{j!}\,\left(\sum_{a\geq 0}P_{3+a}\left(\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right)^{j}$
$=$ $\sum_{j\geq 0}\lambda^{j}\,\sum_{a\geq 0}P_{3j+a}\left(\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)$
where $P_{k}\left(\cdot\right)$ is a generic homogeneous polynomial of degree $k$ and, say,
$L_{\nu,l,s}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)=\sum_{k\geq 0}L_{\nu,l,s,k}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right),$
where each $L_{\nu,l,s,k}\left(x_{1},x_{2};\cdot\right)$ is a homogeneous polynomial of total degree $k$. Upon rescaling, the product of these expansions gives rise to terms of the form
$\lambda^{j-s-\frac{k}{2}-\frac{3j}{2}-\frac{a}{2}}\,P_{3j+a}\left(\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\,L_{\nu,l,s,k}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)$
$=$ $\lambda^{-\frac{1}{2}\,(j+2s+a+k)}\,\tilde{P}_{k+3j+a}\left(\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)$
where $\tilde{P}_{k+3j+a}$ is homogeneous of total degree $3j+a+k$. We obtain the same conclusions as before, that is general terms of the form $\lambda^{-k/2}\,P_{r}$ with $r\leq 3\,k$ and rkr-k even.
7 Applications
7.1 The equivariant Weyl law for $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$
Proof of Theorem 31.
Let $\chi$ have sufficiently small support and satisfy $\hat{\chi}>0$ and $\chi(0)>0$. Recalling (18),
$\sum_{j}\hat{\chi}(\lambda-\lambda_{j})\,\dim H^{\tau}(X)_{j,\nu}$
$=$ $\sum_{j}\hat{\chi}(\lambda-\lambda_{j})\,\int_{X^{\tau}}\,\Pi^{\tau}_{j,\nu}(x,x)\,\mathrm{d}V_{X^{\tau}}(x)$
$=$ $\int_{X^{\tau}}\,\left[\sum_{j}\hat{\chi}(\lambda-\lambda_{j})\,\Pi^{\tau}_{j,\nu}(x,x)\right]\,\mathrm{d}V_{X^{\tau}}(x)$
$=$ $\int_{X^{\tau}}\,\Pi^{\tau}_{\chi,\nu,\lambda}(x,x)\,\mathrm{d}V_{X^{\tau}}(x).$
Let us fix $C>0$ and $\epsilon^{\prime}\in(0,1/6)$. By Theorem 7, we only loose a rapidly decreasing contribution in (7.1), if integration is restricted to a shrinking tubular neighourhood of $Z^{\tau}$ of radius $C\,\lambda^{\epsilon^{\prime}-\frac{1}{2}}$. Furthermore, any such neighbourhood may be locally parametrized using smoothly varying systems of NHLC’s centered at moving points $x\in Z^{\tau}$.
More precisely, in view of Remark 19, for any $x\in Z^{\tau}$ we may find an open neighbourhood $Z^{\prime}\subseteq Z^{\tau}$ of $x$ and a smoothly varying family of normal Heisenberg coordinates centered at points $x^{\prime}\in Z^{\prime}$, such that the map
$(x^{\prime},\mathbf{v}^{t})\in Z^{\prime}\times\mathbb{R}^{d_{G}}\mapsto x^{\prime}+\mathbf{v}^{t}\in X^{\tau}$ (176)
is a diffeomorphism onto a neighbourhood of $x$ in $X^{\tau}$ (here we use normal Heisenberg coordinates at each $x^{\prime}$ to identify $\mathbb{R}^{d_{G}}\cong\mathbb{R}^{d_{G}}_{t}\cong T_{x^{\prime}}^{t}X^{\tau}$). In view of Corollary 35 of [P-2024], if $x=x^{\prime}+\mathbf{v}^{t}$ then
$\mathrm{d}V_{X^{\tau}}(x)=2^{d_{G}/2}\,\mathcal{V}(x^{\prime}+\mathbf{v}^{t})\,\mathrm{d}V_{Z^{\tau}}(x^{\prime})\,\mathrm{d}\mathbf{v}^{t},$ (177)
where $\mathrm{d}V_{Z^{\tau}}$ is the Riemannian volume form on $Z^{\tau}$ for the restricted metric, and $\mathcal{V}(x^{\prime})=1$ for any $x^{\prime}\in Z^{\prime}$.
If we pass to the rescaled local parametrization
$x_{\lambda}(\mathbf{v}^{t}):=x^{\prime}+\frac{\mathbf{v}^{t}}{\sqrt{\lambda}}\qquad(x^{\prime}\in Z^{\prime},\,\mathbf{v}^{t}\in\mathbb{R}^{d_{G}}),$ (178)
then uniformly for $\|\mathbf{v}^{t}\|\leq C\,\lambda^{\epsilon^{\prime}}$ there are asymptotic expansion
$\Pi^{\tau}_{\chi,\nu,\lambda}\left(x_{\lambda}(\mathbf{v}^{t}),x_{\lambda}(\mathbf{v}^{t})\right)$ $\sim$ $\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot\frac{\dim(\nu)^{2}}{V_{eff}(x_{1})}$
$\cdot e^{-\frac{2}{\tau}\,\left\|\mathbf{v}^{t}\right\|^{2}}\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k/2}\,F_{k,\chi,\nu}\left(x^{\prime};\mathbf{v}^{t}\right)\right],$
where $F_{k,\chi,\nu}(x^{\prime};\cdot)$ is a polynomial of degree $\leq 3\,k$ and parity $k$.
Let us choose an open cover of $Z^{\tau}$ by open sets $Z^{\prime}_{j}$ as above, and a subordinate partition of unity $\beta_{j}$ on $Z^{\tau}$; we can then express the latter integral in (7.1) as a sum of local contributions. For notational simplicity, we shall force notation and leave the partition $\{\beta_{j}\}$ implicit. We obtain:
$\sum_{j}\hat{\chi}(\lambda-\lambda_{j})\,\dim H^{\tau}(X)_{j,\nu}$
$\sim$ $2^{d_{G}/2}\int_{Z^{\tau}}\,\mathrm{d}V_{Z^{\tau}}(x)\left[\int_{\mathbb{R}^{d_{G}}}\,\Pi^{\tau}_{\chi,\nu,\lambda}(x+\mathbf{v}^{t},x+\mathbf{v}^{t})\,\,\mathcal{V}(x+\mathbf{v}^{t})\,\mathrm{d}\mathbf{v}^{t}\right]$
$=$ $\left(\frac{\lambda}{2}\right)^{-d_{G}/2}\,\int_{Z^{\tau}}\,\mathrm{d}V_{Z^{\tau}}(x)\left[\int_{\mathbb{R}^{d_{G}}}\,\Pi^{\tau}_{\chi,\nu,\lambda}\left(x_{\lambda}(\mathbf{v}^{t}),x_{\lambda}(\mathbf{v}^{t})\right)\,\,\mathcal{V}(x_{\lambda}(\mathbf{v}^{t}))\,\mathrm{d}\mathbf{v}^{t}\right].$
Integration in $\mathrm{d}\mathbf{v}^{t}$ is now over an expanding ball of radius $O\left(\lambda^{\epsilon^{\prime}}\right)$ in $\mathbb{R}^{d_{G}}$. By multiplying the asymptotic expansion of Theorem 28 with the Taylor expansion of $\mathcal{V}(x_{\lambda}(\mathbf{v}^{t}))$, we obtain an asymptotic expansion for the integrand in (7.1) of the form
$\Pi^{\tau}_{\chi,\nu,\lambda}\left(x_{\lambda}(\mathbf{v}^{t}),x_{\lambda}(\mathbf{v}^{t})\right)\,\,\mathcal{V}(x_{\lambda}(\mathbf{v}^{t}))$
$\sim$ $\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot\frac{\dim(\nu)^{2}}{V_{eff}(x)}\cdot e^{-\frac{2}{\tau}\,\left\|\mathbf{v}^{t}\right\|^{2}}$
$\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k/2}\,F_{k,\nu}\left(x;\mathbf{v}^{t}\right)\right],$
where $F_{k,\nu}(x;\cdot)$ is a polynomial of degree $\leq 3\,k$ and parity $k$. The expansion (7.1) may be integrated term by term; we obtain an asymptotic expansion for (7.1) of the form
$\sum_{j}\hat{\chi}(\lambda-\lambda_{j})\,\dim H^{\tau}(X)_{j,\nu}$
$\sim$ $\frac{2^{d_{G}/2}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\,\lambda^{-d_{G}/2}\cdot\dim(\nu)^{2}$
$\cdot\left(\int_{Z^{\tau}}\,\frac{1}{V_{eff}(x)}\,\mathrm{d}V_{Z^{\tau}}(x)\right)\cdot\int_{\mathbb{R}^{d_{G}}}\,e^{-\frac{2}{\tau}\,\left\|\mathbf{v}^{t}\right\|^{2}}\,\mathrm{d}\mathbf{v}^{t}\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k}\,\mathfrak{f}_{k,\nu}\right]$
$=$ $\frac{1}{2^{d_{G}/2}\,\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k}\,\mathfrak{f}_{k,\nu}\right].$
As in [P-2024], we shall now follow a standard argument from spectral analysis (see e.g. [GrSj]). Let $f_{\lambda}:\mathbb{R}\times\mathbb{R}\rightarrow[0,+\infty)$ be defined by
$f_{\lambda}(s,t):=\hat{\chi}(t)\cdot H(\lambda-s-t),$ (183)
where $H$ is the Heaviside function. Let $\mathcal{L}$ be the Lebsesgue measure on $\mathbb{R}$, and let $\mathcal{P}^{\tau}_{\nu}$ be the positive measure on $\mathbb{R}$ given by
$\mathcal{Q}^{\tau}_{\nu}:=\sum_{j\geq 1}\dim H^{\tau}(X)_{j,\nu}\,\delta_{\lambda_{j}};$ (184)
here $\delta_{a}$ is the delta measure at $a\in\mathbb{R}$. Let us endow $\mathbb{R}\times\mathbb{R}$ with the product meaure $\mathcal{P}^{\tau}_{\nu}\times\mathcal{L}$. By the Fubini Theorem,
$\int_{\mathbb{R}}\,\mathrm{d}\mathcal{L}(t)\,\left[\int_{\mathbb{R}}\,f_{\lambda}(s,t)\,\mathrm{d}\mathcal{Q}^{\tau}_{\nu}(s)\right]=\int_{\mathbb{R}}\,\mathrm{d}\mathcal{Q}^{\tau}_{\nu}(s)\left[\int_{\mathbb{R}}\,f_{\lambda}(s,t)\,\mathrm{d}\mathcal{L}(t)\right].$ (185)
The right hand side in (185) is
$\int_{\mathbb{R}}\,\mathrm{d}\mathcal{Q}^{\tau}_{\nu}(s)\left[\int_{\mathbb{R}}\,f_{\lambda}(s,t)\,\mathrm{d}\mathcal{L}(t)\right]$
$=$ $\int_{\mathbb{R}}\,\mathrm{d}\mathcal{Q}^{\tau}_{\nu}(s)\left[\int_{-\infty}^{\lambda-s}\,\hat{\chi}(t)\,\mathrm{d}\mathcal{L}(t)\right]=\int_{\mathbb{R}}\,\mathrm{d}\mathcal{Q}^{\tau}_{\nu}(s)\left[\int_{-\infty}^{\lambda}\,\hat{\chi}(t-s)\,\mathrm{d}\mathcal{L}(t)\right]$
$=$ $\sum_{j}\dim H^{\tau}(X)_{j,\nu}\,\int_{-\infty}^{\lambda}\,\hat{\chi}(t-\lambda_{j})\,\mathrm{d}\mathcal{L}(t)$
$=$ $\int_{-\infty}^{\lambda}\,\left[\sum_{j}\,\hat{\chi}(t-\lambda_{j})\cdot\dim H^{\tau}(X)_{j,\nu}\right]\,\mathrm{d}\mathcal{L}(t)$
In view of (7.1), we conclude that as $\lambda\rightarrow+\infty$
$\int_{\mathbb{R}}\,\mathrm{d}\mathcal{Q}^{\tau}_{\nu}(s)\left[\int_{\mathbb{R}}\,f_{\lambda}(s,t)\,\mathrm{d}\mathcal{L}(t)\right]$
$=$ $\frac{1}{2^{d_{G}/2}\,\sqrt{2\,\pi}}\,\cdot\frac{\lambda}{d-d_{G}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
$\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k}\,F_{k,\nu}\left(x\right)\right].$
On the other hand, the left hand side in (185) is
$\int_{\mathbb{R}}\,\mathrm{d}\mathcal{L}(t)\,\left[\int_{\mathbb{R}}\,f_{\lambda}(s,t)\,\mathrm{d}\mathcal{Q}^{\tau}_{\nu}(s)\right]$
$=$ $\int_{\mathbb{R}}\,\left[\sum_{j}\dim H^{\tau}(X)_{j,\nu}\,\,\hat{\chi}(t)\cdot H(\lambda-\lambda_{j}-t)\right]\,\mathrm{d}\mathcal{L}(t)$
$=$ $\int_{\mathbb{R}}\,\left[\sum_{\lambda_{j}\leq\lambda-t}\dim H^{\tau}(X)_{j,\nu}\,\right]\,\hat{\chi}(t)\,\mathrm{d}\mathcal{L}(t)$
$=$ $\int_{\mathbb{R}}\,\mathcal{W}_{\nu}^{\tau}(\lambda-t)\,\hat{\chi}(t)\,\mathrm{d}\mathcal{L}(t)$
$=$ $\sqrt{2\,\pi}\,\chi(0)\,\mathcal{W}_{\nu}^{\tau}(\lambda)+\int_{-\infty}^{+\infty}\,\left[\mathcal{W}_{\nu}^{\tau}(\lambda-t)-\mathcal{W}_{\nu}^{\tau}(\lambda)\right]\,\hat{\chi}(t)\,\mathrm{d}\mathcal{L}(t).$
An adaptation of the argument in the proof of Lemma 70 of [P-2024], paired with (7.1), yields the estimate
$\int_{-\infty}^{+\infty}\,\left[\mathcal{W}_{\nu}^{\tau}(\lambda-t)-\mathcal{W}_{\nu}^{\tau}(\lambda)\right]\,\hat{\chi}(t)\,\mathrm{d}\mathcal{L}(t)=O\left(\lambda^{d-1-d_{G}}\right)$ (189)
for $\lambda\rightarrow+\infty$. We conclude
$\mathcal{W}_{\nu}^{\tau}(\lambda)$ $=$ $\frac{2^{d_{G}/2}}{2^{d-1}\,(2\,\pi)}\,\cdot\frac{\lambda}{d-d_{G}}\,\left(\frac{\lambda}{\pi\,\tau}\right)^{d-1-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
$\cdot\left[1+O\left(\lambda^{-1}\right)\right]$
$=$ $\frac{1}{2^{d_{G}/2}}\cdot\frac{\tau}{d-d_{G}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
$\cdot\left[1+O\left(\lambda^{-1}\right)\right]$
7.2 Pointwise estimates on eigenfunctions
Proof of Proposition 32.
If $x\not\in Z^{\tau}$, there exists an open neighbourhood $Y_{x}\subseteq X^{\tau}$ of $x$ such that $\Pi^{\tau}_{\chi,\nu,\lambda}(x^{\prime},x^{\prime})=O\left(\lambda^{-\infty}\right)$ for $\lambda\rightarrow+\infty$, uniformly for $x^{\prime}\in Y_{x}$. If $x\in X^{\tau}$, there exists similarly a neighbourhood $Y_{x}$ where the asymptotic expansion of Theorem 30 implies $\Pi^{\tau}_{\chi,\nu,\lambda}(x^{\prime},x^{\prime})\leq C_{x}\,\lambda^{d-1-d_{G}/2}$ for some constant $C_{x}$, uniformly for $x^{\prime}\in Y_{x}$. The claim follows by the compactness of $X^{\tau}$.
7.3 Equivariant operator norm estimates
We can now prove the equivariant version of the operator norm estimate of Chang and Rabinowitz (29); a similar statement may be proved with minor changes for $P^{\tau}_{\chi,\nu,\lambda}$.
Proof of Theorem 34.
By the Shur-Young inequality, for some positive constant $D^{\tau}_{p}$
$\left\|\Pi^{\tau}_{\chi,\nu,\lambda}\right\|_{L^{p}\rightarrow L^{q}}\leq D_{p}^{\tau}\,\left[\sup_{y\in X^{\tau}}\int_{X^{\tau}}\left|\Pi^{\tau}_{\chi,\nu,\lambda}(y^{\prime},y)\right|^{r}\,\mathrm{d}V_{X^{\tau}}(y^{\prime})\right]^{\frac{1}{r}},$ (190)
where
$\frac{1}{r}:=1-\frac{1}{p}+\frac{1}{q}.$
Let us choose $C>0$ and $\epsilon^{\prime}\in(0,1/6)$. By Theorem 7, $\Pi^{\tau}_{\chi,\lambda}(y^{\prime},y)=O\left(\lambda^{-\infty}\right)$ uniformly for $\mathrm{dist}_{X^{\tau}}\left(y,Z^{\tau}\right)\geq C\,\lambda^{\epsilon^{\prime}-1/2}$. Thus we may assume without loss of generality that $\mathrm{dist}_{X^{\tau}}\left(y,Z^{\tau}\right)\leq C\,\lambda^{\epsilon^{\prime}-1/2}$. Any such $y$ may be written in the form
$y=x+\frac{\mathbf{v}^{t}}{\sqrt{\lambda}},\quad\text{where}\quad x\in Z^{\tau},\quad\mathbf{v}^{t}\in T^{t}_{x}X^{\tau},\quad\|\mathbf{v}^{t}\|\leq 2\,C\,\lambda^{\epsilon^{\prime}}.$ (191)
Let $0<a<A$ be such that
$a\,\mathrm{dist}_{X^{\tau}}\left(x^{\prime},x^{\prime\prime}\right)\leq\mathrm{dist}_{X^{\tau}}\left(\Gamma^{\tau}_{t}(x^{\prime}),\Gamma^{\tau}_{t}(x^{\prime\prime})\right)\leq A\,\mathrm{dist}_{X^{\tau}}\left(x^{\prime},x^{\prime\prime}\right)$
for all $x^{\prime},x^{\prime\prime}\in X^{\tau}$ and $(g,t)\in G\times\mathrm{supp}(\chi)$.
Lemma 84.
Uniformly for $y$ as in (191) and
$\mathrm{dist}_{X^{\tau}}\left(y^{\prime},x^{G\times\chi}\right)\geq 3\,A\,C\,\lambda^{\epsilon^{\prime}-1/2}.$ (192)
we have $\Pi^{\tau}_{\chi,\lambda}(y^{\prime},y)=O\left(\lambda^{-\infty}\right)$.
Proof.
If (192) holds, then for any $(g,t)\in G\times\mathrm{supp}(\chi)$
$\mathrm{dist}_{X^{\tau}}\left(y^{\prime},\mu_{g}\circ\Gamma^{\tau}_{t}(y)\right)$ (193)
$\geq$ $\mathrm{dist}_{X^{\tau}}\left(y^{\prime},\mu_{g}\circ\Gamma^{\tau}_{t}(x)\right)-\mathrm{dist}_{X^{\tau}}\left(\mu_{g}\circ\Gamma^{\tau}_{t}(y),\mu_{g}\circ\Gamma^{\tau}_{t}(x)\right)$
$=$ $\mathrm{dist}_{X^{\tau}}\left(y^{\prime},\mu_{g}\circ\Gamma^{\tau}_{t}(x)\right)-\mathrm{dist}_{X^{\tau}}\left(\Gamma^{\tau}_{t}(y),\Gamma^{\tau}_{t}(x)\right)$
$\geq$ $\mathrm{dist}_{X^{\tau}}\left(y^{\prime},\mu_{g}\circ\Gamma^{\tau}_{t}(x)\right)-2\,A\,C\,\lambda^{\epsilon^{\prime}-\frac{1}{2}}\geq A\,C\,\lambda^{\epsilon^{\prime}-\frac{1}{2}},$
whence $\mathrm{dist}_{X^{\tau}}\left(y^{\prime},x^{G\times\chi}\right)\geq A\,C\,\lambda^{\epsilon^{\prime}-\frac{1}{2}}$. The claim follows by Theorem 7. ∎
Hence integration in (190) may be restricted to the shrinking locus where (192) is satisfied. Any $y^{\prime}\in X^{\tau}$ satisfying (192) may in turn be written uniquely (for a given choice of smoothly varying normal NHLC’s) in the form
$y^{\prime}=\Gamma^{\tau}_{t}\circ\mu^{\tau}_{g}(x)+\frac{\mathbf{u}^{t}+\mathbf{u}^{h}}{\sqrt{\lambda}},$ (194)
where $t\in\mathrm{supp}(\chi)$, $g\in G$ and
$\mathbf{u}^{t}\in T^{t}_{\Gamma^{\tau}_{t}\circ\mu^{\tau}_{g}(x)}X^{\tau},\quad\mathbf{u}^{h}\in T^{h}_{\Gamma^{\tau}_{t}\circ\mu^{\tau}_{g}(x)}X^{\tau},\quad\left\|\mathbf{u}^{t}\right\|,\,\left\|\mathbf{u}^{h}\right\|=O\left(\lambda^{\epsilon^{\prime}}\right).$
In view of statement 3. of Theorem 25 and of the previous considerations on uniformity, there exist constants $a,\,C_{\nu}>0$ such that, uniformly for all such choices,
$\left|\Pi^{\tau}_{\chi,\nu,\lambda}(y^{\prime},y)\right|^{r}$ $=$ $\left|\Pi^{\tau}_{\chi,\nu,\lambda}\left(\Gamma^{\tau}_{t}\circ\mu^{\tau}_{g}(x)+\frac{\mathbf{u}^{t}+\mathbf{u}^{h}}{\sqrt{\lambda}},x+\frac{\mathbf{v}^{t}}{\sqrt{\lambda}}\right)\right|^{r}$ (195)
$\leq$ $C_{\nu}\,\left(\frac{\lambda}{\tau}\right)^{r\,(d-1-d_{G}/2)}\,e^{-a\,\frac{r}{\tau}\left(\|\mathbf{u}^{t}\|+\|\mathbf{u}^{h}\|^{2}+\|\mathbf{v}^{t}\|^{2}\right)}$
Let us set
$F_{r}(y):=\int_{X^{\tau}}\left|\Pi^{\tau}_{\chi,\nu,\lambda}(y^{\prime},y)\right|^{r}\,\mathrm{d}V_{X^{\tau}}(y^{\prime})\qquad(y\in X^{\tau}).$ (196)
Using an unparametrized version of (194), that is, $y^{\prime}=\Gamma^{\tau}_{t}\circ\mu^{\tau}_{g}(x)+\mathbf{u}^{t}+\mathbf{u}^{h}$, we have $\mathrm{d}V_{X^{\tau}}(y^{\prime})=\mathcal{V}(g,t,\mathbf{u}^{t},\mathbf{u}^{h})\,\mathrm{d}g\,\mathrm{d}t\,\mathrm{d}\mathbf{u}^{t}\,\mathrm{d}\mathbf{u}^{h}$ for some positive function $\mathcal{V}(g,t,\mathbf{u}^{t},\mathbf{u}^{h})$, and integration in $\mathrm{d}\mathbf{u}^{t}\,\mathrm{d}\mathbf{u}^{h}$ is on a shrinking ball of radius $O\left(\lambda^{\epsilon^{\prime}-1/2}\right)$. Applying the rescaling in (194) yields $\mathrm{d}V_{X^{\tau}}(y^{\prime})=\lambda^{-(d-1-d_{G})-d_{G}/2}\,\mathcal{V}(g,t,\mathbf{0}^{\dagger},\mathbf{0}^{h})\,\mathrm{d}g\,\mathrm{d}t\,\mathrm{d}\mathbf{u}^{t}\,\mathrm{d}\mathbf{u}^{h}+L.O.T.$, where $L.O.T.$ denotes lower order terms in $\lambda$, and integration in in $\mathrm{d}\mathbf{u}^{t}\,\mathrm{d}\mathbf{u}^{h}$ is now on an expanding ball of radius $O\left(\lambda^{\epsilon^{\prime}}\right)$.
Given this, we obtain from (195) that uniformly for $x\in Z^{\tau}$ and $\mathbf{v}^{t}\in T^{t}_{x}X^{\tau}$ with $\|\mathbf{v}^{t}\|\leq 2\,C\,\lambda^{\epsilon^{\prime}}$ we have (for some constant $D_{\nu}>0$)
$F_{r}\left(x+\frac{\mathbf{v}^{t}}{\sqrt{\lambda}}\right)$ $\leq$ $D_{\nu,r}\,\left(\frac{\lambda}{\tau}\right)^{r\,\left(d-1-\frac{d_{G}}{2}\right)}\,\left(\frac{\tau}{\lambda}\right)^{d-1-\frac{d_{G}}{2}}$ (197)
$=$ $D_{\nu,r}\,\left(\frac{\lambda}{\tau}\right)^{(r-1)\,(d-1-d_{G}/2)}.$
Hence, uniformly in $y\in X^{\tau}$ we have for some constant $D_{\nu,r}^{\prime}>0$ that
$\left[\sup_{y\in X^{\tau}}\int_{X^{\tau}}\left|\Pi^{\tau}_{\chi,\nu,\lambda}(y^{\prime},y)\right|^{r}\,\mathrm{d}V_{X^{\tau}}(y^{\prime})\right]^{\frac{1}{r}}$ $\leq$ $D_{\nu,r}^{\prime}\,\left(\frac{\lambda}{\tau}\right)^{\left(1-\frac{1}{r}\right)\,\left(d-1-\frac{d_{G}}{2}\right)}$
$=$ $D_{\nu,r}^{\prime}\,\left(\frac{\lambda}{\tau}\right)^{\left(\frac{1}{p}-\frac{1}{q}\right)\,\left(d-1-\frac{d_{G}}{2}\right)}.$
7.4 Proof of Proposition 38 and Theorem 40
Proof of Proposition 38.
In the setting of Theorem 31, in place of (7.1) we have
$P^{\tau}_{\chi,\nu,\lambda}\left(x_{\lambda}(\mathbf{v}^{t}),x_{\lambda}(\mathbf{v}^{t})\right)$
$\sim$ $\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{1}{2}\right)^{d-1-d_{G}/2}\left(\frac{\lambda}{\pi\,\tau}\right)^{(d-1-d_{G})/2}\cdot\frac{\dim(\nu)^{2}}{V_{eff}(x_{1})}$
$\cdot e^{-\frac{2}{\tau}\,\left\|\mathbf{v}^{t}\right\|^{2}}\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k/2}\,F_{k,l,\nu}\left(x;\mathbf{v}^{t}\right)\right],$
where is $x_{\lambda}(\mathbf{v}^{t})$ is as in (178) and $F_{k,l,\nu}(x;\cdot)$ is a polynomial of degree $\leq 3\,k$ and parity $k$. The statement of the Theorem follows by integration. ∎
Proof of Theorem 40.
In analogy with (184), let us define the positive measure
$\widetilde{\mathcal{T}}^{\tau}_{\nu}:=\sum_{j\geq 1}\Lambda_{j,\nu}\,\delta_{\mu_{j}}$ (199)
where
$\Lambda_{j,\nu}:=e^{-2\,\tau\,\mu_{j}}\,\sum_{k}\left\|\widetilde{\varphi}^{\tau}_{j,\nu,k}\right\|_{L^{2}(X^{\tau})}^{2}.$
With $f_{\lambda}$ as in (183), we have as in (7.1)
$\int_{\mathbb{R}}\,\mathrm{d}\widetilde{\mathcal{T}}^{\tau}_{\nu}(s)\left[\int_{\mathbb{R}}\,f_{\lambda}(s,t)\,\mathrm{d}\mathcal{L}(t)\right]=\int_{-\infty}^{\lambda}\,\left[\sum_{j}\,\hat{\chi}(t-\mu_{j})\cdot\Lambda_{j,\nu}\right]\,\mathrm{d}\mathcal{L}(t).$ (200)
In view of Proposition 38, we conclude that as $\lambda\rightarrow+\infty$
$\int_{\mathbb{R}}\,\mathrm{d}\widetilde{\mathcal{T}}^{\tau}_{\nu,\lambda}(s)\left[\int_{\mathbb{R}}\,f_{\lambda}(s,t)\,\mathrm{d}\mathcal{L}(t)\right]$
$=$ $\frac{1}{\sqrt{2\pi}}\,\frac{1}{2^{d-1-d_{G}/2}}\,\left(\frac{\lambda}{\pi\,\tau}\right)^{\frac{d-1}{2}-d_{G}}\,\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
$\cdot\lambda\,\left[\dfrac{\chi(0)}{\frac{d+1}{2}-d_{G}}+\sum_{k\geq 1}\,\lambda^{-k}\,\mathfrak{f}_{k}^{\prime\prime}\right].$
Similarly, arguing as for (7.1) we obtain
$\int_{\mathbb{R}}\,\mathrm{d}\mathcal{L}(t)\,\left[\int_{\mathbb{R}}\,f_{\lambda}(s,t)\,\mathrm{d}\widetilde{\mathcal{T}}^{\tau}_{\nu}(s)\right]$
$=$ $\sqrt{2\,\pi}\,\chi(0)\,\mathcal{P}_{\nu}^{\tau}(\lambda)+\int_{-\infty}^{+\infty}\,\left[\mathcal{P}_{\nu}^{\tau}(\lambda-t)-\mathcal{P}_{\nu}^{\tau}(\lambda)\right]\,\hat{\chi}(t)\,\mathrm{d}\mathcal{L}(t)$
$=$ $\sqrt{2\,\pi}\,\chi(0)\,\mathcal{P}_{\nu}^{\tau}(\lambda)+O\left(\lambda^{\frac{d-1}{2}-d_{G}}\right),$
where the last equality follows from Proposition 38 and the argument in Lemma 70 of [P-2024].
By the Fubini Theorem, (7.4) equals (7.4); the claim follows.
8 Index of Notation
  1. 1.
    $(M,\kappa)$: the given $d$-dimensional Riemannian manifold; $(\tilde{M},J)$: its complexification;
  2. 2.
    $\rho:\tilde{M}\rightarrow[0,+\infty)$: the strictly plurisubharmonic exhaustion function determined by $\kappa$;
  3. 3.
    $h_{0},\,\omega_{0},\,g_{0}$: the standard hermitian, symplectic and Riemannian structure. $\Omega,\,\hat{\kappa}$: see (1); $\omega$ and $\tilde{\kappa}$: see Remark 14;
  4. 4.
    $\psi_{2}^{\omega}$: see Definition 11;
  5. 5.
    $G$: $d_{G}$-dimensional Lie group; $\mathfrak{g},\,\mathfrak{g}_{X^{\tau}}(x)$: see Notation 16;
  6. 6.
    $\tilde{M}^{\tau}$, $X^{\tau}$: see (2); $\alpha^{\tau}$: see (4);
  7. 7.
    $\mathcal{H}^{\tau}$, $\mathcal{R}^{\tau}$, $\mathcal{T}^{\tau}$: see (5), (7), (8); $T^{v}X^{\tau},\,T^{t}X^{\tau},\,T^{h}X^{\tau}$: see Definition 18;
  8. 8.
    $\Delta,\,\mu_{j},\,W_{j},\,\varphi_{j,k}$: see (3);
  9. 9.
    $\Pi^{\tau}$, $H(X^{\tau})$: see (9); $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$: see (10); $H(X^{\tau})_{j}$, $\rho_{j,k}$: see (12);
  10. 10.
    $P^{\tau}_{\chi,\,\lambda}$: see (11); $\lambda_{j}$, $\rho_{j,k}$, $\Pi^{\tau}_{\chi,\,\lambda}$: see (12);
  11. 11.
    $\mu:G\times M\rightarrow M$: isometric Lie group action; $\hat{G}$, $\nu$, $V_{\nu}$, $\Xi_{\nu}$, $\dim(\nu)$: see Notation 3; $V_{eff}$: see Definition 21;
  12. 12.
    $\tilde{\mu}^{\tau},\,Z,\,Z^{\tau}$: see (19);
  13. 13.
    $x_{j,\,\lambda}$: see (21); $x_{12}$: see (24);
  14. 14.
    the matrix $B$: see (23), (25);
  15. 15.
    $W_{j,\nu},\,H(X^{\tau})_{j,\nu}$: see (14) and (15); $P_{\nu}$: see (18);
  16. 16.
    $P^{\tau}_{\chi,\nu,\lambda}$: see (16); $\Pi^{\tau}_{\chi,\nu,\lambda}$: see (17);
  17. 17.
    $\mathrm{d}V_{G}(g)$: the Haar measure on $G$; $\mathrm{d}V_{X^{\tau}}(x)$: the measure on $X^{\tau}$ associated to the Riemannian volume form $\mathrm{vol}^{R}_{X^{\tau}}$ (see §3.2.3 of [P-2024]);
  18. 18.
    $\Gamma^{\tau}_{t}$, $x^{G\times\chi}$, $\mathfrak{X}^{\tau}_{\chi}$, $x^{\chi}$, $x^{G}$: see Definition 6 and the discussion immediately preceding it;
  19. 19.
    the matrices $A_{c}$, $P$, $Q$ associated to a symplectic matrix AA: see Definition 8; the function $\Psi_{A}:\mathbb{R}^{2d-2}\times\mathbb{R}^{2d-2}\rightarrow\mathbb{C}$ associated to a symplectic matrix AA: see Definition 9;
  20. 20.
    $e^{\imath\theta^{\tau}_{t}(x)},\,e^{\imath\tilde{\theta}^{\tau}_{t}(x)}$: see Remark 26;
  21. 21.
    $A_{\chi}$: see Definition 22; $\mathcal{F}_{\chi}$, $\mathcal{B}_{\nu}$: see Definition 23;
  22. 22.
    $\psi^{\tau}$, $s^{\tau}$: see (34); $r^{\tau}_{t}$: see (60);
  23. 23.
    $\Sigma_{\chi}(x_{1},x_{2})$: see (38);
  24. 24.
    $\mathcal{W}^{\tau}_{\nu}(\lambda)$: see (28); $\mathcal{P}^{\tau}_{\nu}(\lambda)$: see (32);
  25. 25.
    $\psi_{\gamma},\,J_{\mathrm{ad}}$: see Notation 41;
  26. 26.
    $\Phi,\,\varphi^{\xi},\,\xi_{X^{\tau}}^{\sharp}$: see Notation 49;
  27. 27.
    $\mathbf{v}_{g}$: see (44); $\mathbf{v}^{(l)}$: see (47); $\mathbf{v}(\tau)$: see (107); $\mathbf{v}^{v},\,\mathbf{v}^{t},\,\mathbf{v}^{h}$: see Definition 18;
  28. 28.
    $U_{\sqrt{\rho}}(t),\,\Pi^{\tau}_{t},\,\tilde{\Pi}^{\tau}_{t},\,J^{\tau}_{t},\,\mathcal{P}^{\tau}_{t},\,\sigma^{\tau}_{t,\,j},\,\sigma^{(x)}_{J}$: see §3.1;
  29. 29.
    $\Psi(x_{1},x_{2};g,t,y,u,v)$: see (68); $\mathcal{I}_{k}(\theta_{1},\mathbf{v}_{!},\theta_{2},\mathbf{v}_{2})$: see (4.1); $A(\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{u})_{l}$: see (4.1);
  30. 30.
    $U(\imath\tau),\,\mathfrak{B}^{\tau},\,\tilde{\varphi}_{j,k},\,U_{\mathbb{C}}(t+2\imath\tau),\,\mathcal{Q}^{\tau}_{t},\gamma^{\tau}_{t},\gamma^{\tau}_{t,j}$: see §5;
  31. 31.
    $\Re,\,\Im$: the real and imaginary part of a complex number;
  32. 32.
    The Fourier transform of $f\in\mathcal{S}(\mathbb{R})$ is
    $\hat{f}(\lambda)=\frac{1}{\sqrt{2\,\pi}}\,\int_{-\infty}^{+\infty}\,e^{-\imath\,\lambda\,t}\,f(t)\,\mathrm{d}t.$
  33. 33.
    $\mathrm{dist}_{X^{\tau}}:X^{\tau}\times X^{\tau}\rightarrow\mathbb{R}$: the Riemannian distance function on $X^{\tau}$.
  34. 34.
    $\bullet^{\dagger}$ indicates the transpose of a matrix or a vector, $\bullet^{*}$ indicates the conjugate transpose of a matrix.
Acknowledgments. The authors are members of GNSAGA (Gruppo Nazionale per le Strutture Algebriche, Geometriche e le loro Applicazioni) of INdAM (Istituto Nazionale di Alta Matematica ”Francesco Severi”) and thank the group for its support.